### v1

# Optimal Distributed Similarity Estimation for Unitary Channels

Abstract
We study distributed similarity estimation for unitary channels (DSEU), the task of estimating the similarity between unitary channels implemented on different quantum devices. We completely address DSEU by showing that, for $n$-qubit unitary channels, the query complexity of DSEU is $\Theta(\sqrt{d})$, where $d=2^{n}$, for both incoherent and coherent accesses. First, we propose two estimation algorithms for DSEU with these accesses utilizing the randomized measurement toolbox. The query complexities of these algorithms are both $\mathcal{O}(\sqrt{d})$. Although incoherent access is generally weaker than coherent access, our incoherent algorithm matches this complexity by leveraging additional shared randomness between devices, highlighting the power of shared randomness in distributed quantum learning. We further establish matching lower bounds, proving that $\Theta(\sqrt{d})$ queries are both necessary and sufficient for DSEU. Finally, we compare our algorithms with independent classical shadow and show that ours have a square-root advantage. Our results provide practical and theoretically optimal tools for quantum devices benchmarking and for distributed quantum learning.
I Introduction
The engineering and physical realization of quantum computers are being actively pursued across a wide range of physical platforms [1, 2, 3]. To certify their performance, numerous protocols have been developed that compare experimentally generated quantum states or channels with known theoretical targets, including direct fidelity estimation [4, 5], random benchmarking [6, 7, 8], and quantum verification [9, 10, 11, 12]. However, a central challenge is how to directly compare unknown quantum states or channels produced on different platforms, possibly at different times and locations. This task, known as cross-platform verification, becomes especially pressing as we approach regimes of quantum advantage, where classical simulation of quantum systems becomes computationally intractable. For quantum states, the core step of cross-platform verification is distributed inner product estimation. A variety of efficient algorithms have been developed for this learning task [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Notably, for $n$-qubit quantum states, it has been shown that the sample complexity of completing this task is $\Theta(\sqrt{d})$ [15], where $d=2^{n}$.
A natural next question is how to efficiently estimate the similarity between two unitary channels. We term this learning task as distributed similarity estimation for unitary channels (DSEU). Concretely, given two $n$-qubit unitaries $U$ and $V$ implemented on separate quantum devices, the goal of DSEU is to estimate their similarity using only local quantum operations and classical communication (LOCC). This task has attracted significant attention in several recent works [24, 25, 23]. DSEU has important applications in quantum information theory, including circuit equivalence checking [26, 27, 28] and quantum channel benchmarking [29, 30, 31, 32]. Despite its importance, both lower bound and the optimal algorithm for completing DSEU remain unknown.
In this work, we completely address DSEU by proving the lower bound and the corresponding optimal algorithm for both incoherent and coherent accesses. First, based on the randomized measurement toolbox [33], we provide two algorithms to accomplish DSEU with these two accesses. We show that both algorithms require only $\mathcal{O}(\sqrt{d})$ queries to the unitary channels on each device. We then establish matching lower bounds, proving that any algorithm for DSEU must make at least $\Omega(\sqrt{d})$ queries, even using multi-round LOCC, arbitrarily large ancillary systems, and adaptive state preparation and measurement (SPAM) settings. Together, these results demonstrate that our algorithms are optimal and that the query complexity of completing DSEU is $\Theta(\sqrt{d})$. We also compare our algorithms with independently performing classical shadow for unitary channels [34]. We show that independent classical shadow require $\mathcal{O}(d)$ queries to the unitary channels, whereas our protocols achieve a quadratic improvement. This square-root advantage arises from the shared randomness of SPAM settings.
II Problem Setup
Let $\mathcal{H}$ be the Hilbert space of an $n$-qubit system with dimension $d=2^{n}$. Let $\mathcal{U}(\cdot)=U(\cdot)U^{\dagger}$ and $\mathcal{V}(\cdot)=V(\cdot)V^{\dagger}$ be two unknown unitary channels implemented on different (possibly distant) quantum devices. The goal of DSEU is to estimate the similarity measure $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}\in[0,1]$ up to additive error $\varepsilon\in(0,1)$ only using LOCC. In the following, we describe the learning models relevant to this task, focusing on two aspects: learning access and shared randomness.
Figure 1: Learning models for distributed similarity estimation of unitary channels. (a) Learning access: (i) Incoherent access: the unknown channel is queried once before each measurement. (ii) Coherent access: the unknown channel is queried $T$ times before each measurement, with arbitrary intermediate quantum channels $\{\mathcal{C}_{t}\}$. (b) Shared randomness: (i) Devices share randomness used in both state preparation and measurement (SPAM) settings. (ii) Devices share randomness only for state preparation. (iii) Devices have no shared randomness.
Learning Access.
Following the classifications in [35, 36, 37, 38, 34, 39], we consider two types of learning access: incoherent and coherent accesses, as shown in Fig. 1(a). For clarity, we describe the access model for one device. (i) Incoherent access, also known as learning without quantum memory or without control, allows the learner to use the unknown channel only once before measurement. That is, in the $t$-th query, the learner prepares an arbitrary quantum state $\rho_{t}$, applies either $(\mathcal{I}\otimes\mathcal{U})(\rho_{t})$ or $(\mathcal{I}\otimes\mathcal{V})(\rho_{t})$, and then performs an arbitrary quantum measurement on the output state, which is described by a positive operator-valued measure (POVM). Here $\mathcal{I}$ is the identity channel acting on an ancillary system. (ii) Coherent access, also known as learning with quantum memory or with control, permits multiple queries of the unknown channel before measurement. In each experimental round, the learner may interleave uses of the unknown channel with arbitrary quantum channels. Therefore, the resulting output state before the measurement has the form
$(\mathcal{U}\otimes\mathcal{I})\circ\mathcal{C}_{T-1}\circ(\mathcal{U}\otimes\mathcal{I})\circ\cdots\mathcal{C}_{1}\circ(\mathcal{U}\otimes\mathcal{I})(\rho),$ (1)
with an analogous expression for $\mathcal{V}$, where $\rho$ is the initial state and $\{\mathcal{C}_{i}\}$ are arbitrary quantum channels. An arbitrary POVM is then applied to the final state.
Shared Randomness.
In addition to learning accesses, shared randomness plays a crucial role in randomized measurement-based distributed learning. As in many learning protocols, SPAM settings are essential for learning channels. In the framework of the randomized measurement toolbox, both initial states and measurement settings are sampled randomly to enhance practicality and sample efficiency. Thus, allowing devices to share part or all of randomness can significantly improve performance, as demonstrated in distributed inner product estimation [15]. In this work, we consider three levels of shared randomness, illustrated in Fig. 1(b): (i) shared randomness in both SPAM settings, (ii) shared randomness only in state preparation, and (iii) no shared randomness between devices. We will show that shared randomness provides clear advantages in the efficiency of distributed learning algorithms.
III Upper bound
Here, we provide two algorithms to complete DSEU and provide the corresponding query complexities.
III.1 Incoherent Access
We first consider learning with incoherent access and share the randomness of SPAM settings. Let $\{|{a}\rangle\!\langle{a}|\}_{a}$ be the computational basis. To complete DSEU, we run the following procedure $T$ times independently:
  1. 1.
    Randomly generate a pure state $|{\psi}\rangle\in\mathcal{H}$ from a state $4$-design ensemble on each quantum devices.
  2. 2.
    Apply the unitary channels $\mathcal{U}$ and $\mathcal{V}$ on each $|{\psi}\rangle$.
  3. 3.
    Measure two rotated states $\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)$ and $\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)$ in the basis $\{Q^{\dagger}|{a}\rangle\!\langle{a}|Q\}_{a}$ for $m$ times, where $Q$ is randomly sampled from a unitary $4$-design ensemble.
  4. 4.
    Recode the measurement results as $\{a_{i}\}_{i=1}^{m}$ and $\{b_{i}\}_{i=1}^{m}$, respectively. Compute
    $\tilde{g}=\frac{1}{m^{2}}\sum_{i,j=1}^{m}\mathbb{1}(a_{i},b_{j}),$ (2)
    where $\mathbb{1}(a,b)=1$ if $a=b$; otherwise 0.
  5. 5.
    Obtain an unbiased estimator for $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$:
    $\tilde{\omega}=\frac{(d+1)^{2}}{d}\tilde{g}-\frac{d+2}{d}.$ (3)
This incoherent algorithm is summarized in Algorithm 1. As we can see, our incoherent algorithm is ancilla-free and non-adaptive, making it immediately implementable on near-term devices. Additionally, the corresponding query complexity is shown in the following theorem.
Theorem 1.
The expectation and variance of estimator $\tilde{\omega}$ defined in Eq. (3) are given by
$\mathbb{E}\tilde{\omega}=\frac{1}{d^{2}}\operatorname{Tr}^{2}[U^{\dagger}V],\quad\mathbb{V}(\tilde{\omega})=\mathcal{O}\left(\frac{d}{m^{2}}+\frac{1}{m}+\frac{1}{d}\right).$ (4)
Consequently, the query complexity is $\mathcal{O}(\max\{1/\varepsilon^{2},\sqrt{d}/\varepsilon\})$.
The proof is provided in Appendix B.1. We remark that this incoherent algorithm naturally generalizes to estimating similarities between arbitrary channels. Concretely, we can estimate the inner product of two unknown channels’ Choi states [40], and then estimate the similarity via several inner product-based distance measures [24, 41]. Furthermore, this incoherent algorithm can also serve as a tool of unitarity estimation for an unknown quantum channel [38].
III.2 Coherent Access
We now turn to the coherent access and assume that two devices share only the randomness of state preparation. In various quantum learning scenarios, coherent access often provides substantial advantages compared to incoherent access [37, 38, 34, 39]. Here, we provide a coherent algorithm based on a symmetric collective measurement applied on quantum states in $\mathcal{H}^{\otimes T}$, labeled as $\mathcal{M}_{T}$ [42, 34]. $\mathcal{M}_{T}$ can be constructed as the following POVM:
$\mathcal{M}_{T}\mathrel{\mathop{\ordinarycolon}}=\left\{\frac{\kappa_{T}}{L}|{\phi_{j}}\rangle\!\langle{\phi_{j}}|^{\otimes T}\right\}_{j=1}^{L}\cup\{I-\Pi_{\rm sym}^{(d,T)}\},$ (5)
where $\{|{\phi_{j}}\rangle\}$ forms a state $(T+2)$-design, $\kappa_{T}\mathrel{\mathop{\ordinarycolon}}=\binom{d+T-1}{T}$, and $\Pi_{\rm sym}^{(d,T)}$ is the projector onto the symmetric subspace of $\mathcal{H}^{\otimes T}$. With this measurement, we provide the following algorithm:
  1. 1.
    Randomly generate pure states $|{\psi}\rangle^{\otimes T}$ on each quantum devices, where $|{\psi}\rangle$ from a state $4$-design ensemble;
  2. 2.
    Apply the unitary channels $\mathcal{U}^{\otimes T}$ and $\mathcal{V}^{\otimes T}$ on $|{\psi}\rangle^{\otimes T}$.
  3. 3.
    Measure $(\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes T}$ and $(\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes T}$ with the POVM $\mathcal{M}_{T}$. Recode the measurement results on two devices as $|{\phi_{A}}\rangle$ and $|{\phi_{B}}\rangle$ and compute $\tilde{f}=|\langle\phi_{A}|\phi_{B}\rangle|^{2}$.
  4. 4.
    Obtain an unbiased estimator for $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$:
    $\tilde{\chi}\mathrel{\mathop{\ordinarycolon}}=\frac{(d+1)(d+T)^{2}}{T^{2}d}\tilde{f}-\frac{(d+1)(d+2T)+T^{2}}{T^{2}d}.$ (6)
This algorithm is described in pseudocode in Algorithm 2. We summarize the query complexity of the coherent algorithm in the following theorem, proven in Appendix B.2.
Theorem 2.
The expectation and variance of estimator $\tilde{\chi}$ defined in Eq. (6) are given by
$\mathbb{E}\tilde{\chi}=\frac{1}{d^{2}}\operatorname{Tr}^{2}[U^{\dagger}V],\quad\mathbb{V}(\tilde{\chi})=\mathcal{O}\left(\frac{d}{m^{2}}+\frac{1}{m}+\frac{1}{d}\right).$ (7)
Consequently, the query complexity is $\mathcal{O}(\max\{1/\varepsilon^{2},\sqrt{d}/\varepsilon\})$.
Surprisingly, although incoherent access is generally weaker than coherent access, the two algorithms have the same query complexity of completing DSEU. Let’s elaborate the differences between these two algorithms. First, unlike the incoherent algorithm, the coherent algorithm fundamentally requires that the unknown channels be unitary and the proof relies on pure output states; check out details in Appendix B.2. Second, the coherent algorithm does not reduce to the incoherent algorithm when $T=1$, as they have different classical data post-processing, caused by the different types of applied shared randomness. Specifically, the coherent algorithm requires sharing only the randomness of state preparation, whereas the incoherent algorithm additionally relies on shared measurement settings. This shared measurement randomness effectively boosts the performance of the incoherent algorithm, enabling it to achieve the $\mathcal{O}(\sqrt{d})$ query complexity.
IV Lower Bound
In this section, we provide the lower bounds for completing DSEU, showing that both two algorithms proposed in this work achieve optimal query complexity. To identify the lower bound of DSEU, we consider a related hypothesis testing task, which serves as a restricted version of DSEU. This is a task of distinguishing between the following two cases:
Two devices perform the same unitary $U$, which is a Haar random unitary;
Two devices independently perform two unitaries $U$ and $V$, which are both independent Haar random unitaries.
If a learning algorithm can estimate values of $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$, then, we can use this learning algorithm to complete the above hypothesis testing task. Hence, a lower bound for this task directly implies a lower bound for completing DSEU. Moreover, this lower bound also applies to estimating other commonly used distance measures between quantum channels. In the following, we analyze the query complexity required to solve this hypothesis testing task with incoherent and coherent accesses.
We employ Le Cam’s two-point method [43] to upper bound the success probability of solving the hypothesis testing problem. This method relates the success probability to the total variation distance (TVD) between the probability distributions of measurement results under the two hypotheses. Let $p_{1}(\ell)$ and $p_{2}(\ell)$ be the probabilities of obtaining measurement result $\ell$ after querying two unitary channels in the first and second cases. Their TVD is defined as
$\left\lVert p_{1}-p_{2}\right\rVert_{\rm TV}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{2}\sum_{\ell}\left|p_{1}(\ell)-p_{2}(\ell)\right|,$ (8)
To reliably distinguish the two hypotheses, this TVD must be at least constantly large. Now, we are ready to prove the lower bounds for two learning accesses.
IV.1 Incoherent Access
We first consider learning with incoherent access, and follow the framework of learning tree [37, 38, 44, 45, 46]. We show that for incoherent access, completing DSEU needs exponentially many queries, even learning with multi-round LOCC, arbitrarily large ancillary systems, and adaptive SPAM settings. The result is shown in the following theorem.
Theorem 3.
For any algorithm with incoherent access that can be possible multi-round LOCC, ancilla-assisted, and adaptive, requires $\Omega(\sqrt{d})$ queries to complete DSEU.
This theorem shows that our incoherent algorithm in Section III.1 is optimal, yielding the query complexity of completing DSEU with incoherent coherent access is $\Theta(\sqrt{d})$. This scaling matches that of unitarity estimation [38], consistent with the fact that unitarity estimation is a special case of DSEU.
Proof sketch of Theorem 3.
Inspired by [38], we introduce an intermediate distribution $p_{\mathcal{D}}(\ell)$, which is the probability of obtaining measurement result $\ell$ when both channels are replaced by the completely depolarizing channel, defined as $\mathcal{D}(A)\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}[A]I/d$. Then, with the triangle inequality, the target TVD is upper bounded by two TVDs: $\left\lVert p_{1}-p_{\mathcal{D}}\right\rVert_{\rm TV}$ and $\left\lVert p_{\mathcal{D}}-p_{2}\right\rVert_{\rm TV}$. Lastly, using techniques in [38], each of these terms is upper bounded by $\mathcal{O}(T^{2}/d)$, where $T$ is the number of queries made to each device. Therefore, completing DSEU with incoherent access requires query complexity $\Omega(\sqrt{d})$. The detailed proof is provided in Appendix C.1. ∎
IV.2 Coherent Access
We now turn to coherent access, where the learner may interleave multiple queries to the unknown channel with arbitrary quantum channels before the measurement. Perhaps surprisingly, we show that even under this significantly more powerful model, completing DSEU still requires exponentially many queries. This lower bound holds even when learning with multi-way LOCC, arbitrarily large ancillary systems, and adaptive SPAM settings. This result further highlights the fundamental hardness of DSEU. Our result is stated as follows.
Theorem 4.
For any algorithm with coherent access that can be possible multi-round LOCC, ancilla-assisted, and adaptive, requires $\Omega(\sqrt{d})$ queries to complete DSEU.
Thus our coherent algorithm in Section III.2 is also optimal, and the query complexity of completing DSEU with coherent access is also $\Theta(\sqrt{d})$. Consequently, coherent access offers no advantage for completing DSEU. This phenomenon is consistent with the result of distributed inner product estimation [15], which show that symmetric collective measurements do not reduce the complexity in distributed learning tasks. Taken together, these results highlight that shared randomness, rather than coherent access or collective measurements, is the key resource that enables efficient distributed learning.
Proof sketch of Theorem 4.
Assume we query $T$ times on each devices. Generally, analyzing the complexity of coherent access is subtle, as arbitrary quantum channels may appear between queries. To handle this difficulty, inspired by [47], we show that it suffices to work with the Choi operator of $\mathcal{U}^{\otimes T}$, defined as $J_{U}^{(T)}\mathrel{\mathop{\ordinarycolon}}=(\mathcal{U}^{\otimes T}\otimes\mathcal{I})(|{\Phi}\rangle\!\langle{\Phi}|)$, where $|{\Phi}\rangle\mathrel{\mathop{\ordinarycolon}}=\sum|{ii}\rangle$ is the unnormalized maximally entangled states in $(\mathcal{H}^{\otimes T})^{\otimes 2}$. By formulating the problem in terms of Choi operators, all interleaving channels can be absorbed into an enlarged effective input system, reducing the analysis to distinguishing between the two averaged Choi operators
$\bar{J}_{1}^{(T)}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U}\;J_{U}^{(T)}\otimes\mathbb{E}_{V}\;J_{V}^{(T)},\quad\bar{J}_{2}^{(T)}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U}\;J_{U}^{(T)}\otimes J_{U}^{(T)},$ (9)
which correspond to the two hypothesis testing cases.
Subsequently, as in the proof of lower bound for incoherent access, we introduce an intermediate probability distribution $p_{a}(\ell)$, defined as the probability of obtaining measurement result $\ell$ when both channels are replaced by the approximate $T$-design channel with Choi operator $J_{a}^{(T)}$ (see definition in [47], also Appendix Eq. (25)). We again analyze the TVDs: $\left\lVert p_{1}-p_{a}\right\rVert_{\rm TV}$ and $\left\lVert p_{a}-p_{2}\right\rVert_{\rm TV}$ separately. For the first term, we directly bound it using the relative error of the approximate $T$-design channel, and the corresponding result is $\mathcal{O}(T^{2}/d)$. For the second term, we relate it to the distance between the approximate $T$-design channel and a “measure-and-prepare” map. Using tools from quantum cloning theory [48],we show that this term is also upper bounded by $\mathcal{O}(T^{2}/d)$. Therefore, completing DSEU with coherent access requires query complexity $\Omega(\sqrt{d})$. The detailed proof is provided in Appendix C.2. ∎
V Independent Classical Shadow
We now compare our algorithms with independently performing classical shadow for unitary channels [34], where no shared randomness is available. For completeness, we briefly outline how DSEU can be completed using independent classical shadow; the full procedure is given in Algorithm 3. For each device, we randomly select a state $|{\psi}\rangle$ in $\mathcal{H}$ from a state $4$-design ensemble, then apply $\mathcal{U}^{\otimes s}$ on $|{\psi}\rangle^{\otimes s}$, and finally perform the symmetric collective measurement $\mathcal{M}_{s}$. Repeating the above procedure $T$ times on each devices yields two independent sets of classical snapshots of $\mathcal{U}$ and $\mathcal{V}$, labeled as $\{X_{t}\}$ and $\{Y_{t}\}$. Lastly, the similarity can then be estimated via
$\tilde{\gamma}=\frac{1}{T^{2}d^{2}}\sum_{i,j=1}^{T}\operatorname{Tr}[X_{i}^{\dagger}Y_{j}].$ (10)
Unlike our algorithms, this method uses no shared randomness, i.e., SPAM settings are chosen independently on the two devices. We provide the corresponding query complexity of this algorithm in the following theorem.
Theorem 5.
The expectation and variance of estimator $\tilde{\gamma}$ defined in Eq. (10) are given by
$\mathbb{E}\tilde{\gamma}=\frac{1}{d^{2}}\operatorname{Tr}^{2}[U^{\dagger}V],\quad\mathbb{V}(\tilde{\gamma})=\mathcal{O}\left(\frac{d^{2}}{T^{2}s^{2}}+\frac{1}{T}\right).$ (11)
Consequently, the query complexity is $\mathcal{O}(\max\{1/\varepsilon^{2},d/\varepsilon\})$.
The proof is provided in Appendix D. This result shows that, regardless of how many queries are used to generate a single classical snapshot, completing DSEU with independent classical shadow requires $\mathcal{O}(d)$ queries. In contrast, both of our algorithms achieve query complexity $\Theta(\sqrt{d})$, revealing a square-root advantage that arises precisely from the shared randomness of SPAM settings.
VI Conclusion
In this work, we showed that the query complexity of completing DSEU is $\Theta(\sqrt{d})$, for both incoherent and coherent accesses. Firstly, we proposed two randomized measurement based estimation algorithms that effectively utilize shared randomness between the two quantum devices. Then, we proved that it is necessary and sufficient to take $\Theta(\sqrt{d})$ queries to complete DSEU. We also compared our algorithms with independent classical shadow, demonstrating a square-root advantage in query complexity. This advantage arises from the shared randomness employed in our algorithms, which allows the two devices to coordinate SPAM settings. The obtained results are summarized in Table 1.
Several questions remain open. For example, existing cross-platform verification protocols often assume that the two platforms implement the same operation, an assumption that may fail in realistic experimental settings due to device imperfections. Developing robust verification algorithms that relax this assumption is an important direction for future research.
Learning Access Upper Bound Lower Bound
Incoherent $\mathcal{O}(\sqrt{d})$ (Theorem 1) $\Omega(\sqrt{d})$ (Theorem 3)
Coherent $\mathcal{O}(\sqrt{d})$ (Theorem 2) $\Omega(\sqrt{d})$ (Theorem 4)
Classical Shadow $\mathcal{O}(d)$ (Theorem 5)
Table 1: Our results on the query complexity of completing DSEU, where $d$ is the dimension of the unitary channels.
Note added.
While completing this manuscript, we became aware of a related work by Ananth et al. [49], which also studied the distinguishing problem constructed in our lower bound with coherent access. While both works identify the $\Omega(\sqrt{d})$ barrier, the proof techniques are distinct and complementary. Our approach relies on Le Cam’s two-point method and the connection to optimal quantum cloning theory [48], whereas Ananth et al. utilized linear programming. The diversity of proof techniques enriches the understanding of the fundamental hardness of this distinguishing problem.
Acknowledgments.
This work was supported by the National Natural Science Foundation of China (Grant No. 62471126), the Jiangsu Frontier Technology Research and Development Plan (Grant No. BF2025066), the Fundamental Research Funds for the Central Universities (Grant No. 2242022k60001), the SEU Innovation Capability Enhancement Plan for Doctoral Students (Grant No. CXJH_SEU 24083), and the National Key R&D Program of China (Grant No. 2022YFF0712800).
Appendix A Preliminaries
Let $\mathcal{H}$ be the Hilbert space of an $n$-qubit system, with dimension $d=2^{n}$. The set of Hermitian operators on $\mathcal{H}$ is denoted by $\mathcal{B}(\mathcal{H})$, the set of linear operators on $\mathcal{H}$ is denoted by $\mathcal{L}(\mathcal{H})$, the set of density matrices on $\mathcal{H}$ is denoted by $\mathcal{D}(\mathcal{H})$. For a pure state $|{\phi}\rangle\!\langle{\phi}|\in\mathcal{D}(\mathcal{H})$, we use the shorthand $\phi\mathrel{\mathop{\ordinarycolon}}=|{\phi}\rangle\!\langle{\phi}|$. For a unitary $U\in\mathcal{L}(\mathcal{H})$, let $\mathcal{U}(A)\mathrel{\mathop{\ordinarycolon}}=UAU^{\dagger}$ be the corresponding unitary channel. For a quantum channel $\mathcal{E}$, we define its Choi operator as [40]
$J_{\mathcal{E}}\mathrel{\mathop{\ordinarycolon}}=(\mathcal{E}\otimes\mathcal{I})(\Phi),$ (12)
where $|{\Phi}\rangle\mathrel{\mathop{\ordinarycolon}}=\sum_{i}|{ii}\rangle$ is the unnormalized maximally entangled states. The Choi operator provides the useful identity
$\operatorname{Tr}\left[O\mathcal{E}(\rho)\right]=\operatorname{Tr}\left[\left(O\otimes\rho^{T}\right)J_{\mathcal{E}}\right],$ (13)
which we will frequently employ throughout this work.
For any operator $A\in\mathcal{B}(\mathcal{H})$, we use the notation $|A\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ to represent the corresponding vectorized operator. For example,
$||{\psi}\rangle\!\langle{\phi}|\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\mathrel{\mathop{\ordinarycolon}}=|{\psi}\rangle\otimes|{\phi^{*}}\rangle,\quad|ABC^{\dagger}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=A\otimes C^{*}|B\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}.$ (14)
The inner product between two vectorized operators is defined as $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}A|B\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=\operatorname{Tr}[A^{\dagger}B]$. For operators $A$ and $B$, we write $A\succeq B$ if ABA-B is positive semidefinite.
In the following, we introduce two central tools used in this work: the Haar measure and approximate unitary designs. Both play fundamental roles in quantum information theory [50, 47], and they are essential for establishing the query complexities of our algorithms as well as the lower bounds for completing DSEU.
A.1 Haar Measure and Permutation Operators
Let $\mu_{H}$ be the Haar measure over the unitary group. The Haar random channel is defined as
$\mathcal{E}_{H}^{(k)}(A)\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mu_{H}}U^{\otimes k}AU^{\dagger\otimes k},\quad A\in\mathcal{B}(\mathcal{H}^{\otimes k}).$ (15)
To describe the Haar random channel, we introduce permutation operators. Let $\mathcal{S}_{k}$ be the symmetric group on $k$ elements. For $\pi\in\mathcal{S}_{k}$, the corresponding permutation operator $P_{\pi}$ acts on product states as
$P_{\pi}|{\psi_{1}}\rangle\otimes\cdots\otimes|{\psi_{k}}\rangle=|{\psi_{\pi^{-1}(1)}}\rangle\otimes\cdots\otimes|{\psi_{\pi^{-1}(k)}}\rangle,\quad\forall\;\psi_{1},\cdots\psi_{k}\in\mathcal{D}(\mathcal{H}).$ (16)
With the above definition, we have the following lemma to describe the Haar random channel.
Lemma 6 (Weingarten Calculus [50, 47]).
Let $\mathcal{H}$ be a Hilbert space with dimension $d$ and $\mathcal{S}_{k}$ be the symmetric group and $A\in\mathcal{B}(\mathcal{H}^{\otimes k})$. Then, we have
$\mathcal{E}_{H}^{(k)}(A)=\sum_{\sigma,\tau\in\mathcal{S}_{k}}{\rm Wg}_{\sigma,\tau}(d)\operatorname{Tr}[AP^{\dagger}_{\sigma}]P_{\tau},\quad A\in\mathcal{B}(\mathcal{H}^{\otimes k}).$ (17)
Here, ${\rm Wg}_{\sigma,\tau}(d)$ are the elements of the $k!\times k!$ Weingarten matrix, which is defined as the inverse of the Gram matrix, $G_{\sigma,\tau}(d)=\operatorname{Tr}(P_{\sigma}P_{\tau}^{\dagger})$, given by the inner products of the permutation operators. Consequently, the Choi operator of Haar random channel can be represented as
$J_{H}^{(k)}=\sum_{\sigma,\tau\in\mathcal{S}_{k}}{\rm Wg}_{\sigma,\tau}(d)P_{\tau}\otimes P_{\sigma}.$ (18)
For the special cases $k=1$ and $k=2$, the Haar twirl admits simple closed forms.
Lemma 7 (Special Cases of Weingarten Calculus, $k=1$ and $k=2$ cases [50]).
Let $\mathcal{H}$ be a Hilbert space with dimension $d$, $A\in\mathcal{B}(\mathcal{H})$, and $B\in\mathcal{B}(\mathcal{H}^{\otimes 2})$, we have
$\mathcal{E}_{H}^{(1)}(A)=\frac{\operatorname{Tr}[A]}{d}I,\quad\mathcal{E}_{H}^{(2)}(B)=\frac{d\operatorname{Tr}[B]-\operatorname{Tr}[\mathbb{F}B]}{d(d^{2}-1)}I\otimes I+\frac{d\operatorname{Tr}[\mathbb{F}B]-\operatorname{Tr}[B]}{d(d^{2}-1)}\mathbb{F},$ (19)
where $\mathbb{F}$ is the SWAP operator.
For quantum states in $\mathcal{H}$, we define the Haar measure on quantum states as [50],
$\mathbb{E}_{|{\psi}\rangle\sim\mu_{H}}\psi^{\otimes k}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mu_{H}}U^{\otimes k}\phi^{\otimes k}U^{\dagger\otimes k}=\frac{1}{\kappa_{k}}\Pi_{\rm sym}^{(d,k)},\quad\phi\in\mathcal{D}(\mathcal{H}),$ (20)
where $\kappa_{k}\mathrel{\mathop{\ordinarycolon}}=\binom{d+k-1}{k}$ and $\Pi_{\rm sym}^{(d,k)}$ is the orthogonal projector onto the symmetric subspace of $\mathcal{H}^{\otimes k}$, defined as
$\Pi_{\rm sym}^{(d,k)}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{k!}\sum_{\pi\in\mathcal{S}_{k}}P_{\pi}.$ (21)
Additionally, a pure state ensemble $\mathcal{A}$ is said to form a state $k$-design if
$\mathbb{E}_{|{\psi}\rangle\sim\mathcal{A}}\psi^{\otimes k}=\mathbb{E}_{|{\psi}\rangle\sim\mu_{H}}\psi^{\otimes k}=\frac{\Pi_{\rm sym}^{(d,k)}}{\kappa_{k}}.$ (22)
Similarly, a unitary ensemble $\mathcal{A}$ is said to form a unitary $k$-design if and only if
$\mathcal{E}_{\mathcal{A}}^{(k)}(A)\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mathcal{A}}U^{\otimes k}AU^{\dagger\otimes k}=\mathcal{E}_{H}^{(k)}(A),\quad\forall\;A\in\mathcal{B}(\mathcal{H}^{\otimes k}).$ (23)
A.2 Approximate Unitary Designs
A unitary ensemble $\mathcal{A}$ is an $\varepsilon$-approximate unitary $k$-design if [47]
$(1-\varepsilon)\mathcal{E}_{\mathcal{A}}^{(k)}\preceq\mathcal{E}_{H}^{(k)}\preceq(1+\varepsilon)\mathcal{E}_{\mathcal{A}}^{(k)},$ (24)
where $\varepsilon$ is called relative error. Here $\mathcal{E}_{1}\preceq\mathcal{E}_{2}$ denotes that $\mathcal{E}_{2}-\mathcal{E}_{1}$ is a completely-positive map. A standard approximate $k$-design channel [47] and the corresponding Choi operator are defined as
$\mathcal{E}_{a}^{(k)}(A)\mathrel{\mathop{\ordinarycolon}}=\frac{1}{d^{k}}\sum_{\sigma\in\mathcal{S}_{k}}\operatorname{Tr}[AP_{\sigma}^{\dagger}]P_{\sigma},\quad J_{a}^{(k)}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{d^{k}}\sum_{\pi\in\mathcal{S}_{k}}P_{\pi}\otimes P_{\pi}.$ (25)
For $k^{2}\leq d$, the corresponding relative error satisfies
$\varepsilon=\frac{k^{2}/2d}{1-k^{2}/2d}.$ (26)
We also have the following inequality for $J_{a}^{(k)}$ and $J_{H}^{(k)}$.
Lemma 8 (Lemma 2 of [38]).
Suppose $\frac{d^{k}}{d(d+1)\cdots(d+T-1)}>\frac{1}{2}$, then, we have
$J^{(k)}_{H}\succeq\frac{1}{d(d+1)\cdots(d+k-1)}\sum_{\sigma\in\mathcal{S}_{k}}P_{\sigma}\otimes P_{\sigma}=\frac{d^{k}}{d(d+1)\cdots(d+k-1)}J_{a}^{(k)}.$ (27)
Appendix B Upper Bound
B.1 Incoherent Access: Proof of Theorem 1
The algorithm solving DSEU with incoherent access is summarized in Algorithm 1. In the following, we will prove Theorem 1.
Input: number of SPAM settings $T$,
    number of measurements for each SPAM setting $m$,
    TmTm queries of unknown unitary channels $\mathcal{U}$ and $\mathcal{V}$ acting on $d$-dimension Hilbert space $\mathcal{H}$.
Output: an estimation of $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$.
1 for $t=1,\cdots,T$ do
2   Randomly generate pure states $|{\psi_{t}}\rangle$ from a 44-design state ensemble on each quantum devices.
3  Apply the unitary channels $\mathcal{U}$ and $\mathcal{V}$ on $|{\psi_{t}}\rangle$.
4  Sample a random unitary $Q_{t}$ from an unitary $4$-design ensemble.
5  Measure $m$ copies of $\mathcal{U}(|{\psi_{t}}\rangle\!\langle{\psi_{t}}|)$ in the basis $\{Q_{t}^{\dagger}|{a}\rangle\!\langle{a}|Q_{t}\}_{a}$ and obtain $A=\{a_{i}\}_{i=1}^{m}$.
6  Measure $m$ copies of $\mathcal{V}(|{\psi_{t}}\rangle\!\langle{\psi_{t}}|)$ in the basis $\{Q_{t}^{\dagger}|{b}\rangle\!\langle{b}|Q_{t}\}_{b}$ and obtain $B=\{b_{i}\}_{i=1}^{m}$.
7  Compute the $\tilde{g}_{t}=\sum_{i,j=1}^{m}\mathbb{1}(a_{i},b_{j})/m^{2}$, where $\mathbb{1}(a,b)=1$ if $a=b$; otherwise 0.
8  Compute
$\tilde{\omega}_{t}=\frac{(d+1)^{2}}{d}\tilde{g}_{t}-\frac{d+2}{d}.$ (28)
Return ${\omega}\mathrel{\mathop{\ordinarycolon}}=\sum_{t}\tilde{\omega}_{t}/T$.
Algorithm 1 Distributed Similarity Estimation for Unitary Channels with Incoherent Access
For each SPAM setting $\{Q,\psi\}$, we define the following two probabilities:
$p_{Q,\psi}(a)\mathrel{\mathop{\ordinarycolon}}=\langle{a}|Q\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle,\quad q_{Q,\psi}(b)\mathrel{\mathop{\ordinarycolon}}=\langle{b}|Q\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{b}\rangle,$ (29)
and their classical inner product function
$g(Q,\psi)\mathrel{\mathop{\ordinarycolon}}=\sum_{a}p_{Q,\psi}(a)q_{Q,\psi}(a).$ (30)
Then, the expectation of $\tilde{g}$ can be represented as
$\mathbb{E}_{Q,\psi,A,B}\;\tilde{g}$ $=\mathbb{E}_{Q,\psi}g(Q,\psi)$ (31)
$=\sum_{a}\langle{a}|Q\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle\langle{a}|Q\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle$ (32)
$=\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(\sum_{a}\mathbb{E}_{Q}Q^{\dagger\otimes 2}|{aa}\rangle\!\langle{aa}|Q^{\otimes 2}\right)\left(U|{\psi}\rangle\!\langle{\psi}|U^{\dagger}\otimes V|{\psi}\rangle\!\langle{\psi}|V^{\dagger}\right)\right]$ (33)
$=\frac{1}{d+1}+\frac{1}{d+1}\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(U|{\psi}\rangle\!\langle{\psi}|U^{\dagger}\right)\left(V|{\psi}\rangle\!\langle{\psi}|V^{\dagger}\right)\right],$ (34)
where we have used Lemma 7 and the property of unitary $2$-design. For the last term, we have
$\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(U|{\psi}\rangle\!\langle{\psi}|U^{\dagger}\right)\left(V|{\psi}\rangle\!\langle{\psi}|V^{\dagger}\right)\right]$ $=\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)\left(\mathbb{E}_{\psi}|{\psi}\rangle\!\langle{\psi}|^{\otimes 2}\right)\right]$ (35)
$=\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)\frac{\Pi_{\rm sym}^{(d,2)}}{\kappa_{2}}\right]=\frac{\operatorname{Tr}^{2}[U^{\dagger}V]+d}{d(d+1)}.$ (36)
Therefore, we have
$\quad\mathbb{E}_{Q,\psi,A,B}\;\tilde{g}=\mathbb{E}_{Q,\psi}g(Q,\psi)=\frac{1}{d+1}+\frac{1}{d}\frac{\operatorname{Tr}^{2}[U^{\dagger}V]+d}{d(d+1)},$ (37)
$\Rightarrow$ $\quad\frac{(d+1)^{2}}{d}\mathbb{E}\tilde{g}-\frac{d+2}{d}=\frac{\operatorname{Tr}^{2}[U^{\dagger}V]}{d^{2}}$ (38)
$\Rightarrow$ $\quad\mathbb{E}\tilde{\omega}_{t}=\frac{\operatorname{Tr}^{2}[U^{\dagger}V]}{d^{2}}.$ (39)
Now, we consider the variance of estimator so that we can obtain the query complexity of Algorithm 1. The following analysis is similar to [15]. With the definition of $\omega$}, the variance of $\omega$} is given by
[…]
Lemma 9.
The variance of $\tilde{g}$ with respect to SPAM setting $\{Q,\psi\}$ and measurement results $S\mathrel{\mathop{\ordinarycolon}}=\{A,B\}$ can be upper bounded by
$\mathbb{V}(\tilde{g})\leq\mathcal{O}\left(\frac{1}{m^{2}d}+\frac{1}{md^{2}}+\frac{1}{d^{3}}\right).$ (41)
Proof of Lemma 9.
It should be note that $\tilde{g}$ is a function of SPAM setting $\{Q,\psi\}$ and measurement results $S\mathrel{\mathop{\ordinarycolon}}=\{A,B\}$. Thus, we can write $\tilde{g}$ as $\tilde{g}(Q,\psi,S)$ and the law of total variance gives [15]
$\mathbb{V}(\tilde{g}(Q,\psi,S))=\mathbb{E}_{Q,\psi}\mathbb{V}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right]\mathbb{V}_{Q,\psi}+\mathbb{E}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right].$ (42)
We consider this two terms respectively as follows.
  1. 1.
    For the first term, with Lemma 14 in [15], we have
    $\mathbb{E}_{Q,\psi}\mathbb{V}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right]$ $\leq\mathbb{E}_{Q,\psi}\left[\frac{g(Q,\psi)}{m^{2}}+\frac{1}{m}\sum_{a}\left(p_{Q,\psi}^{2}(a)q_{Q,\psi}(a)+p_{Q,\psi}(a)q_{Q,\psi}^{2}(a)\right)\right].$ (43)
    As shown in Eq. (37), we have
    $\mathbb{E}_{Q,\psi}\frac{g(Q,\psi)}{m^{2}}=\frac{1}{m^{2}(d+1)}+\frac{1}{m^{2}d}\frac{\operatorname{Tr}^{2}[U^{\dagger}V]+d}{d(d+1)}=\mathcal{O}\left(\frac{1}{m^{2}d}\right).$ (44)
    Additionally, with Eq. (187) in [15], we have
    $\mathbb{E}_{Q}p_{Q,\psi}^{2}(a)q_{Q,\psi}(a)$ $=\mathbb{E}_{Q}\sum_{a}\langle{a}|Q\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle^{2}\langle{a}|Q\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle$ (45)
    $=d\mathbb{E}_{\phi}\langle{\phi}|\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}\langle{\phi}|Q\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{\phi}\rangle=\mathcal{O}\left(\frac{1}{d^{2}}\right),$ (46)
    where we use the property of unitary $3$-design. Thus, we have
    $\mathbb{E}_{Q,\psi}p_{Q,\psi}^{2}(a)q_{Q,\psi}(a)=\mathcal{O}\left(\frac{1}{d^{2}}\right),\quad\mathbb{E}_{Q,\psi}p_{Q,\psi}(a)q_{Q,\psi}^{2}(a)=\mathcal{O}\left(\frac{1}{d^{2}}\right).$ (47)
    Therefore, we have
    $\mathbb{E}_{Q,\psi}\mathbb{V}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right]\leq\mathcal{O}\left(\frac{1}{m^{2}d}+\frac{1}{md^{2}}\right).$ (48)
  2. 2.
    For the second term, we have
    $\mathbb{V}_{Q,\psi}\mathbb{E}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right]$ $=\mathbb{V}_{Q,\psi}\;g(Q,\psi)=\mathbb{E}_{Q,\psi}\;g^{2}(Q,\psi)-\left[\mathbb{E}_{Q,\psi}\;g(Q,\psi)\right]^{2}$ (49)
    $=\mathbb{E}_{Q,\psi}\;g^{2}(Q,\psi)-\mathbb{E}_{\psi}\left(\frac{1+\operatorname{Tr}[\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)]}{d+1}\right)^{2}.$ (50)
    We have
    $\mathbb{E}_{Q}\;g^{2}(Q,\psi)=\mathbb{E}_{Q}\;\left(\sum_{a}\langle{a}|Q\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle\langle{a}|Q\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle\right)^{2}$ (51)
    $=$ $\;d\;\mathbb{E}_{\phi}\langle{\phi}|\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}\langle{\phi}|\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}+d(d-1)\mathbb{E}_{\phi,\phi^{\bot}}\langle{\phi\phi^{\bot}}|\mathcal{U}^{\otimes 2}(|{\psi}\rangle\!\langle{\psi}|)|{\phi\phi^{\bot}}\rangle\langle{\phi\phi^{\bot}}|\mathcal{V}^{\otimes 2}(|{\psi}\rangle\!\langle{\psi}|)|{\phi\phi^{\bot}}\rangle,$ (52)
    where $\phi^{\bot}$ is randomly sampled from the orthogonal space of $\phi$, i.e., $\langle\phi|\phi^{\bot}\rangle=0$. As shown in [15, Lemma 23 and Eq. (194)], we have
    $d\;\mathbb{E}_{\phi}\langle{\phi}|\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}\langle{\phi}|\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}=d\mathcal{O}\left(\frac{1}{d^{4}}\right)=\mathcal{O}\left(\frac{1}{d^{3}}\right),$ (53)
    $d(d-1)\mathbb{E}_{\phi,\phi^{\bot}}\langle{\phi\phi^{\bot}}|\mathcal{U}^{\otimes 2}(|{\psi}\rangle\!\langle{\psi}|)|{\phi\phi^{\bot}}\rangle\langle{\phi\phi^{\bot}}|\mathcal{V}^{\otimes 2}(|{\psi}\rangle\!\langle{\psi}|)|{\phi\phi^{\bot}}\rangle=\frac{(1+\operatorname{Tr}[\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)])^{2}}{d(d+1)}+\mathcal{O}\left(\frac{1}{d^{4}}\right).$ (54)
    Therefore, we have
    $\mathbb{V}_{Q,\psi}\mathbb{E}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right]$ $\leq\mathcal{O}\left(\frac{1}{d^{3}}\right)+\frac{1}{d(d+1)^{2}}\mathbb{E}_{\psi}\left(1+\operatorname{Tr}[\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)]\right)^{2}=\mathcal{O}\left(\frac{1}{d^{3}}\right).$ (55)
Combining the above two terms, we have
$\mathbb{V}(\tilde{g}(Q,\psi,S))\leq\mathcal{O}\left(\frac{1}{m^{2}d}+\frac{1}{md^{2}}+\frac{1}{d^{3}}\right).$ (56)
With Lemma 9 and Eq. (40), the variance of estimator $\omega$ can be bounded by
$\mathbb{V}(\tilde{\omega}_{t})\leq\mathcal{O}\left(\frac{d}{m^{2}}+\frac{1}{m}+\frac{1}{d}\right)\quad\Rightarrow\quad\mathbb{V}({\omega})\leq\mathcal{O}\left(\frac{d}{Tm^{2}}+\frac{1}{Tm}+\frac{1}{Td}\right).$ (57)
Therefore, to achieve $\varepsilon$ additive error, the required query times must satisfy
$Tm=\mathcal{O}\left(\max\left\{\frac{1}{\varepsilon^{2}},\frac{\sqrt{d}}{\varepsilon}\right\}\right).$ (58)
B.2 Coherent Access: Proof of Theorem 2
The algorithm solving DSEU with coherent access is summarized in Algorithm 2. In the following, we will prove Theorem 2, which concludes the query complexity of Algorithm 2.
Input: $T$ queries of unknown unitary channels $\mathcal{U}$ and $\mathcal{V}$ acting on $d$-dimension Hilbert space $\mathcal{H}$.
Output: an estimation of $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$.
1
2Randomly generate pure states $|{\psi}\rangle^{\otimes T}$ on each devices, where $|{\psi}\rangle$ is sampled from a 44-design state ensemble.
3Apply the unitary channels $\mathcal{U}^{\otimes T}$ and $\mathcal{V}^{\otimes T}$ on $|{\psi}\rangle^{\otimes T}$.
4Measure $(\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes T}$ with the POVM $\mathcal{M}_{T}$ and obtains result $|{\phi_{A}}\rangle$.
5Measure $(\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes T}$ with the POVM $\mathcal{M}_{T}$ and obtains result $|{\phi_{B}}\rangle$.
Compute $\tilde{f}=|\langle\phi_{A}|\phi_{B}\rangle|^{2}$ and return
$\tilde{\chi}\mathrel{\mathop{\ordinarycolon}}=\frac{(d+1)(d+T)^{2}}{T^{2}d}\tilde{f}-\frac{(d+1)(d+2T)+T^{2}}{T^{2}d}.$ (59)
Algorithm 2 Distributed Similarity Estimation for Unitary Channels with Coherent Access
To obtain the expectation of $\tilde{\chi}$, we compute the expectation of $\tilde{f}$ first:
$\mathbb{E}\tilde{f}$ $=\mathbb{E}|\langle\phi_{A}|\phi_{B}\rangle|^{2}=\mathbb{E}\operatorname{Tr}\left[\left(|{\phi_{A}}\rangle\!\langle{\phi_{A}}|\right)\left(|{\phi_{B}}\rangle\!\langle{\phi_{B}}|\right)\right]$ (60)
$=\int\operatorname{Tr}\left[\left(\frac{I+TU|{\psi}\rangle\!\langle{\psi}|U^{\dagger}}{d+T}\right)\left(\frac{I+TV|{\psi}\rangle\!\langle{\psi}|V^{\dagger}}{d+T}\right)\right]{\rm d}\psi$ Lemma 10
$=\frac{d+2T}{(d+T)^{2}}+\frac{T^{2}}{(d+T)^{2}}\int\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)|{\psi}\rangle\!\langle{\psi}|^{\otimes 2}\right]{\rm d}\psi$ (61)
$=\frac{d+2T}{(d+T)^{2}}+\frac{T^{2}\left(\operatorname{Tr}^{2}[U^{\dagger}V]+d\right)}{d(d+1)(d+T)^{2}}.$ Lemma 7
As we can see, the proof relies on Lemma 10, which requires that the output states remain pure. For a general quantum channel, the output states might be mixed, and the key identities in the lemma no longer hold. Therefore, the coherent algorithm is applicable only when the unknown channels are unitary channels. Lastly, we can prove $\tilde{\chi}$ is an unbiased estimator for $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$,
$\mathbb{E}\tilde{\chi}=\frac{(d+1)(d+T)^{2}}{T^{2}d}\mathbb{E}\tilde{f}-\frac{(d+1)(d+2T)+T^{2}}{T^{2}d}=\frac{\operatorname{Tr}^{2}[U^{\dagger}V]}{d^{2}}.$ (62)
We now consider the variance of the estimator $\tilde{\chi}$. Let $\tilde{\chi}=X\tilde{f}-Y$, where
$X=\frac{(d+1)(d+T)^{2}}{T^{2}d},\quad Y=\frac{(d+1)(d+2T)+T^{2}}{T^{2}d}.$ (63)
Then, we have
[…]
It is suffices to consider $\mathbb{E}\tilde{f}^{2}$. With [15, Proof of Lemma 5, Eqs. (165) and (166)], we have
$X^{2}\mathbb{E}\tilde{f}^{2}$ $\leq\left(\frac{d+1}{d}\right)^{2}\mathbb{E}_{\psi}\left(f_{\psi}^{2}+\frac{8f_{\psi}-2f_{\psi}^{2}}{T}+\frac{2df_{\psi}+f_{\psi}^{2}+8+2d}{T^{2}}+\frac{8d+4}{T^{3}}+\frac{2d^{2}+2d}{T^{4}}\right),$ (65)
where we define
$f_{\psi}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}\left[\left(U|{\psi}\rangle\!\langle{\psi}|U^{\dagger}\right)\left(V|{\psi}\rangle\!\langle{\psi}|V^{\dagger}\right)\right].$ (66)
With Lemma 7, we have
$\mathbb{E}_{\psi}f_{\psi}=\frac{\operatorname{Tr}^{2}[U^{\dagger}V]+d}{d(d+1)}\leq 1,$ (67)
and
$\mathbb{E}_{\psi}f_{\psi}^{2}$ $=\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)^{\otimes 2}|{\psi}\rangle\!\langle{\psi}|^{\otimes 4}\right]=\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)^{\otimes 2}\frac{\Pi_{\rm sym}^{(4)}}{\kappa_{4}}\right]$ (68)
$=\frac{1}{d(d+1)(d+2)(d+3)}\sum_{\pi\in\mathcal{S}_{4}}\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)^{\otimes 2}P_{\pi}\right]\leq\frac{\operatorname{Tr}^{4}[U^{\dagger}V]}{d^{4}}+\mathcal{O}\left(\frac{1}{d}\right)$ (69)
Thus, the variance of $\tilde{\chi}$ is upper bounded by
$\mathbb{V}(\tilde{\chi})\leq\mathcal{O}\left(\frac{1}{T}+\frac{d}{T^{2}}+\frac{d}{T^{3}}+\frac{d^{2}}{T^{4}}\right)+\mathcal{O}\left(\frac{d}{T^{2}}+\frac{1}{T}+\frac{1}{d}\right)=\mathcal{O}\left(\frac{1}{T}+\frac{d}{T^{2}}+\frac{d^{2}}{T^{4}}\right).$ (70)
Therefore, to achieve $\varepsilon$ additive error, we require
$T\geq\mathcal{O}\left(\max\left\{\frac{1}{\varepsilon^{2}},\frac{\sqrt{d}}{\varepsilon}\right\}\right).$ (71)
Lemma 10 (Lemmas 13 and 14 in [42]).
For measurement $\mathcal{M}_{s}$ on pure state $|{\psi}\rangle\!\langle{\psi}|^{\otimes s}$, the expectation of measurement result is
$\mathbb{E}|{\phi}\rangle\!\langle{\phi}|=\frac{I+s|{\psi}\rangle\!\langle{\psi}|}{d+s}.$ (72)
Additionally, we have
$\mathbb{E}|{\phi}\rangle\!\langle{\phi}|^{\otimes 2}=\frac{2}{(d+s)(d+s+1)}\left[(I+s|{\psi}\rangle\!\langle{\psi}|)^{\otimes 2}-\frac{s(s+1)}{2}|{\psi}\rangle\!\langle{\psi}|^{\otimes 2}\right]\Pi_{\rm sym}^{(2)}.$ (73)
Appendix C Lower Bound
In this section, we establish lower bounds for completing DSEU with both incoherent and coherent accesses. Here, we consider two channels $\mathcal{U}$ and $\mathcal{V}$ acting on $n$-qubit subsystems of a Hilbert space $\mathcal{H}\simeq\mathcal{H}_{\rm main}\otimes\mathcal{H}_{\rm aux}$, where $\mathcal{H}_{\rm main}\simeq(\mathbb{C}^{2})^{\otimes n}$ is the ‘main system’ comprising $n$ qubits and $\mathcal{H}_{\rm aux}\simeq(\mathbb{C}^{2})^{\otimes n^{\prime}}$ is an ‘auxiliary system’ of $n^{\prime}$ qubits. It is convenient to define $d=2^{n}$ and $d^{\prime}=2^{n^{\prime}}$. Additionally, as pointed in [37, Remark 4.19], we only need to consider rank-$1$ POVM on two quantum systems.
Our general approach is to reduce the prediction task to a two-hypothesis distinguishing problem. In particular, to find the lower bound of completing DSEU, we consider the following distinguishing problem.
Problem 1 (Distinguishing Problem).
We want to distinguish the following two cases:
  1. 1.
    Two quantum devices perform the same unitary $U$, which is a Haar random unitary;
  2. 2.
    Two quantum devices independently perform two unitaries $U$ and $V$, which are two independent Haar random unitaries.
If a learning algorithm can estimate values of $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$, then, we can use this learning algorithm to complete the above distinguishing problem. Hence, a lower bound for this distinguishing problem also gives a lower bound for DSEU. To bound the success probability of solving the above distinguishing problem, we use Le Cam’s two-point method as follows.
Lemma 11 (Le Cam’s two-point method [43]).
Let $p^{U,U}(\ell)$ and $p^{U,V}(\ell)$ be the probabilities of obtaining measurement outcome $\ell$ under Cases 1 and 2 of Problem 1, respectively. Then, the probability that the learning algorithm correctly solves the distinguishing problem in Problem 1 is upper bounded by total variation distance (TVD), defined as
$\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{2}\sum_{\ell}\left|\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)\right|.$ (74)
In the remainder of this section, we analyze TVD between $p^{U,U}(\ell)$ and $p^{U,V}(\ell)$, respectively, thereby establishing the corresponding lower bounds.
C.1 Incoherent Access
Here we will prove the lower bound for completing DSEU with incoherent access. First, we introduce the tree representation of channel learning with incoherent access, which is a powerful tool in proving the lower bound of learning task [37, 38, 44, 45, 46].
C.1.1 Proof of Theorem 3
Definition 12 (Tree Representation for Learning Quantum Channels [37]).
Consider two channels $\mathcal{U}$ and $\mathcal{V}$ acting on $n$-qubit subsystems of a Hilbert space $\mathcal{H}$. A learning algorithm with incoherent access can be represented as a rooted tree $\mathcal{T}$ of depth $T$ such that each node encodes all measurement outcomes the algorithm has received thus far. The tree has the following properties:
  • Each node $u$ has an associated probability $p^{U,V}(u)$.
  • The root of the tree $r$ has an associated probability $p^{U,V}(r)=1$.
  • At each non-leaf node $u$, we prepare a state $|{\phi_{u,1}}\rangle\otimes|{\phi_{u,2}}\rangle$ on $\mathcal{H}^{\otimes 2}$, apply channels $\mathcal{U}$ and $\mathcal{V}$ onto two $n$-qubit subsystems, and measure a rank-$1$ POVM $\{w_{v}^{u}d^{2}d^{\prime 2}|{\psi^{u}_{v,1}}\rangle\!\langle{\psi^{u}_{v,1}}|\otimes|{\psi^{u}_{v,2}}\rangle\!\langle{\psi^{u}_{v,2}}|\}_{v}$ (which can depend on $u$) on the entire system to obtain a classical outcome $v$. Each child node $v$ of the node $u$ corresponds to a particular POVM outcome $v$ and is connected by the edge $e_{u,v}$. We refer to the set of child node of node $u$ a ${\rm child}(u)$.
  • If $v$ is a child node of $u$, then
    $p^{U,V}(v)=p^{U,V}(u)w_{v}^{u}d^{2}d^{\prime 2}\operatorname{Tr}\left[\left(\bigotimes_{i=1}^{2}|{\psi^{u}_{v,i}}\rangle\!\langle{\psi^{u}_{v,i}}|\right)(\mathcal{U}\otimes\mathcal{I}_{\rm aux}\otimes\mathcal{V}\otimes\mathcal{I}_{\rm aux})\left(\bigotimes_{i=1}^{2}|{\phi_{u,i}}\rangle\!\langle{\phi_{u,i}}|\right)\right].$ (75)
  • Each root-to-leaf path is of length $T$. For a leaf of corresponding to node $\ell$, $p^{U,V}(\ell)$ is the probability that the classical memory is in state $\ell$ after the learning procedure.
Based on the above tree representation and distinguishing problem, we are ready to prove the lower bound for completing DSEU with incoherent access, i.e., Theorem 3. As shown in Lemma 11, our goal is to obtain an upper bound for $\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}$. With triangle inequality, we have the following upper bound for each leaf $\ell$,
$\left|\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)\right|$ $\leq\left|\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)-p^{\mathcal{D}}(\ell)\right|+\left|p^{\mathcal{D}}(\ell)-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)\right|,$ (76)
$\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}$ $\leq\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-p^{\mathcal{D}}\right\rVert_{{\rm TV}}+\left\lVert p^{\mathcal{D}}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}},$ (77)
where $\mathcal{D}(A)\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}[A]I/d$ is the completely depolarizing channel and $p^{\mathcal{D}}(\ell)$ is the probability that both two quantum channels are $\mathcal{D}$. With Lemma 13, we can find the upper bounds for two terms in Eq. (77) and have
$\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}$ $\leq\frac{6T^{2}}{d},$ (78)
which hints that we require $T=\Omega(\sqrt{d})$. Therefore, we complete the proof of Theorem 3.
C.1.2 Technical Lemmas
In the following, we prove the following key lemma.
Lemma 13.
For two terms defined in Eq. (77), we have the following upper bounds,
$\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-p^{\mathcal{D}}\right\rVert_{{\rm TV}}$ $\leq\frac{4T^{2}}{d},\quad\left\lVert p^{\mathcal{D}}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}\leq\frac{2T^{2}}{d}.$ (79)
Proof of Lemma 13.
Before the proof, we introduce some useful notation. Each root-to-leaf path in the learning tree $\mathcal{T}$ is uniquely specified by a sequence of vertices $v_{0},v_{1},\cdots,v_{T}$, where $v_{0}=r$ is the root and $v_{T}=\ell$ is a leaf. Since $\ell$ uniquely determines the shortest path back to the root, specifying the leaf $\ell$ is equivalent to specifying the entire path $v_{0}=r,v_{1},\cdots,v_{T-1},v_{T}=\ell$. Therefore, the probabilities of reaching leaf $\ell$ can be expressed as
$p^{\mathcal{D}}(\ell)$ $=\prod_{t=1}^{T}w_{t}d^{2}d^{\prime 2}\langle{\psi_{t,1}}|\left(\mathcal{D}\otimes\mathcal{I}_{\rm aux}\right)(|{\phi_{t,1}}\rangle\!\langle{\phi_{t,1}}|)|{\psi_{t,1}}\rangle\langle{\psi_{t,2}}|\left(\mathcal{D}\otimes\mathcal{I}_{\rm aux}\right)(|{\phi_{t,2}}\rangle\!\langle{\phi_{t,2}}|)|{\psi_{t,2}}\rangle,$ (80)
$p^{U,V}(\ell)$ $=\prod_{t=1}^{T}w_{t}d^{2}d^{\prime 2}\langle{\psi_{t,1}}|\left(\mathcal{U}\otimes\mathcal{I}_{\rm aux}\right)(|{\phi_{t,1}}\rangle\!\langle{\phi_{t,1}}|)|{\psi_{t,1}}\rangle\langle{\psi_{t,2}}|\left(\mathcal{V}\otimes\mathcal{I}_{\rm aux}\right)(|{\phi_{t,2}}\rangle\!\langle{\phi_{t,2}}|)|{\psi_{t,2}}\rangle,$ (81)
similar to $p^{U,U}$. As shown in [38], we can decompose the $|{\phi_{t,i}}\rangle$ and $|{\psi_{t,i}}\rangle$ as
$|{\phi_{t,i}}\rangle=\sum_{j=0}^{d^{\prime}-1}|{\phi_{t,i,j}}\rangle\otimes|{j}\rangle,\quad|{\psi_{t,i}}\rangle=\sum_{j=0}^{d^{\prime}-1}|{\psi_{t,i,j}}\rangle\otimes|{j}\rangle,\quad i=1,2,$ (82)
where the first tensor factor corresponds to the main system $\mathcal{H}_{\rm main}$ and the second to the auxiliary system $\mathcal{H}_{\rm aux}$. Note that the vectors $|{\phi_{t,i,j}}\rangle$ and $|{\psi_{t,i,j}}\rangle$ are not required to be normalized. For convenience, let $\bm{j}\mathrel{\mathop{\ordinarycolon}}=(\bm{j}_{1},\cdots,\bm{j}_{T})$ range over all sequences of length $T$, with each $j_{t}\in{0,\ldots,d^{\prime}-1}$. Define $W_{\ell}\mathrel{\mathop{\ordinarycolon}}=\prod w_{t}$, and
$|{\Phi_{\ell,i,\bm{j}}}\rangle\mathrel{\mathop{\ordinarycolon}}=\bigotimes_{t=1}^{T}|{\phi_{t,i,\bm{j}_{t}}}\rangle,\quad|{\Psi_{\ell,i,\bm{j}}}\rangle\mathrel{\mathop{\ordinarycolon}}=\bigotimes_{t=1}^{T}|{\psi_{t,i,\bm{j}_{t}}}\rangle,\quad i=1,2,$ (83)
we have
$\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)$ $=\mathbb{E}_{U,V\sim\mu_{H}}\sum_{\bm{j},\bm{k},\bm{x},\bm{y}}\prod_{t=1}^{T}w_{t}d^{2}d^{\prime 2}\langle{\psi_{t,1,\bm{j}_{t}}}|\mathcal{U}(|{\phi_{t,1,\bm{j}_{t}}}\rangle\!\langle{\phi_{t,1,\bm{k}_{t}}}|)|{\psi_{t,1,\bm{k}_{t}}}\rangle\langle{\psi_{t,2,\bm{x}_{t}}}|\mathcal{V}(|{\phi_{t,2,\bm{x}_{t}}}\rangle\!\langle{\phi_{t,2,\bm{y}_{t}}}|)|{\psi_{t,2,\bm{y}_{t}}}\rangle$ (84)
$=(dd^{\prime})^{2T}W_{\ell}\sum_{\bm{j},\bm{k},\bm{x},\bm{y}}\mathbb{E}_{U,V\sim\mu_{H}}\langle{\Psi_{t,1,\bm{j}}}|U^{\otimes T}|{\Phi_{t,1,\bm{j}}}\rangle\!\langle{\Phi_{t,1,\bm{k}}}|U^{\dagger\otimes T}|{\Psi_{t,1,\bm{k}}}\rangle\langle{\Psi_{t,2,\bm{x}}}|V^{\otimes T}|{\Phi_{t,2,\bm{x}}}\rangle\!\langle{\phi_{t,2,\bm{y}}}|V^{\dagger\otimes T}|{\Psi_{t,2,\bm{y}}}\rangle$
$=(dd^{\prime})^{2T}W_{\ell}\prod_{i=1}^{2}\left(\sum_{\bm{j},\bm{k}}\langle{\Psi_{\ell,i,\bm{j}}\otimes\Phi_{\ell,i,\bm{j}}^{*}}|J_{H}^{(T)}|{\Psi_{\ell,i,\bm{k}}\otimes\Phi_{\ell,i,\bm{k}}^{*}}\rangle\right),$ (85)
where $J_{H}^{(T)}$ is the Choi operator of Haar random channel defined in Eq. (18), Likewise, we have
$\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)$ $=(dd^{\prime})^{2T}W_{\ell}\sum_{\bm{j},\bm{k}}\langle{\Psi_{\ell,\bm{j}}\otimes\Phi_{\ell,\bm{j}}^{*}}|J_{H}^{(2T)}|{\Psi_{\ell,\bm{k}}\otimes\Phi_{\ell,\bm{k}}^{*}}\rangle,$ (86)
$p^{\mathcal{D}}(\ell)$ $=d^{\prime 2T}W_{\ell}\sum_{\bm{j},\bm{k}}\langle\Psi_{\ell,\bm{j}}|\Psi_{\ell,\bm{k}}\rangle\langle\Phi_{\ell,\bm{k}}|\Phi_{\ell,\bm{j}}\rangle,$ (87)
where
$|{\Phi_{\ell,\bm{j}}}\rangle\mathrel{\mathop{\ordinarycolon}}=|{\Phi_{\ell,1,\bm{j}^{\prime}}}\rangle\otimes|{\Phi_{\ell,2,\bm{j}^{\prime\prime}}}\rangle,\quad|{\Psi_{\ell,\bm{k}}}\rangle\mathrel{\mathop{\ordinarycolon}}=|{\Psi_{\ell,1,\bm{k}^{\prime}}}\rangle\otimes|{\Psi_{\ell,2,\bm{k}^{\prime\prime}}}\rangle,$ (88)
$\bm{j}^{\prime}=(\bm{j}_{1},\cdots,\bm{j}_{T})$, $\bm{j}^{\prime\prime}=(\bm{j}_{T+1},\cdots,\bm{j}_{2T})$, similar to $\bm{k}^{\prime}$ and $\bm{k}^{\prime\prime}$. Now, we are ready to prove the upper bound for two terms defined in Eq. (77) as follows.
  1. 1.
    For the first term, we can also write it in the following form.
    $\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-p^{\mathcal{D}}\right\rVert_{{\rm TV}}$ $=\frac{1}{2}\sum_{\ell}\left|\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)-p^{\mathcal{D}}(\ell)\right|$ (89)
    $=\sum_{\ell\mathrel{\mathop{\ordinarycolon}}p^{\mathcal{D}}(\ell)\geq\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)}p^{\mathcal{D}}(\ell)\left[1-\frac{\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)}{p^{\mathcal{D}}(\ell)}\right].$ (90)
    Therefore, we can focus on the lower bound for $\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)/p^{\mathcal{D}}(\ell)$ and have
    $\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)$ $=(dd^{\prime})^{2T}W_{\ell}\sum_{\bm{j},\bm{k}}\langle{\Psi_{\ell,\bm{j}}\otimes\Phi_{\ell,\bm{j}}^{*}}|J_{H}^{(2T)}|{\Psi_{\ell,\bm{k}}\otimes\Phi_{\ell,\bm{k}}^{*}}\rangle$ (91)
    $\geq\frac{(dd^{\prime})^{2T}W_{\ell}}{d(d+1)\cdots(d+2T-1)}\sum_{\bm{j},\bm{k}}\langle{\Psi_{\ell,\bm{j}}\otimes\Phi_{\ell,\bm{j}}^{*}}|\left(\sum_{\sigma\in\mathcal{S}_{2T}}P_{\sigma}\otimes P_{\sigma}\right)|{\Psi_{\ell,\bm{k}}\otimes\Phi_{\ell,\bm{k}}^{*}}\rangle$ Lemma 8
    $=\frac{(dd^{\prime})^{2T}W_{\ell}}{d(d+1)\cdots(d+2T-1)}\sum_{\bm{j},\bm{k}}\langle{\Psi_{\ell,\bm{j}}}|\left(\sum_{\sigma\in\mathcal{S}_{2T}}P_{\sigma}|{\Psi_{\ell,\bm{k}}}\rangle\!\langle{\Phi_{\ell,\bm{k}}}|P_{\sigma}^{\dagger}\right)|{\Phi_{\ell,\bm{j}}}\rangle.$ (92)
    Define $\mathcal{P}_{\pi}\mathrel{\mathop{\ordinarycolon}}|X_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\otimes|X_{2}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\otimes\cdots\otimes|X_{k}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\mapsto|X_{\pi^{-1}(1)}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\otimes|X_{\pi^{-1}(2)}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\otimes\cdots\otimes|X_{\pi^{-1}(k)}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ for $\pi\in\mathcal{S}_{k}$ [38], then, we have
    $\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)$ $\geq\frac{(dd^{\prime})^{2T}W_{\ell}}{d(d+1)\cdots(d+2T-1)}\sum_{\bm{j},\bm{k}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}|{\Psi_{\ell,\bm{j}}}\rangle\!\langle{\Phi_{\ell,\bm{j}}}||\left(\sum_{\sigma\in\mathcal{S}_{2T}}\mathcal{P}_{\sigma}\right)||{\Psi_{\ell,\bm{k}}}\rangle\!\langle{\Phi_{\ell,\bm{k}}}|\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ (93)
    $\geq\frac{(dd^{\prime})^{2T}W_{\ell}}{d(d+1)\cdots(d+2T-1)}\sum_{\bm{j},\bm{k}}\langle\Psi_{\ell,\bm{j}}|\Psi_{\ell,\bm{k}}\rangle\langle\Phi_{\ell,\bm{k}}|\Phi_{\ell,\bm{j}}\rangle.$ Lemma 14
    Thus, for each leaf $\ell$, we have
    $\frac{\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)}{p^{\mathcal{D}}(\ell)}\geq\frac{d^{2T}}{d(d+1)\cdots(d+2T-1)}=\prod_{t=1}^{2T}\left(1+\frac{t-1}{d}\right)^{-1}\geq\prod_{t=1}^{2T}\left(1-\frac{t-1}{d}\right)\geq\left(1-\frac{2T}{d}\right)^{2T}\geq 1-\frac{4T^{2}}{d}.$ (94)
    Therefore, we have
    $\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-p^{\mathcal{D}}\right\rVert_{{\rm TV}}$ $=\sum_{\ell\mathrel{\mathop{\ordinarycolon}}p^{\mathcal{D}}(\ell)\geq\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)}p^{\mathcal{D}}(\ell)\left[1-\frac{\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)}{p^{\mathcal{D}}(\ell)}\right]\leq\frac{4T^{2}}{d}.$ (95)
  2. 2.
    Likewise, we have
    $\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)$ $=(dd^{\prime})^{2T}W_{\ell}\prod_{i=1}^{2}\left(\sum_{\bm{j},\bm{k}}\langle{\Psi_{\ell,i,\bm{j}}\otimes\Phi_{\ell,i,\bm{j}}^{*}}|J_{H}^{(T)}|{\Psi_{\ell,i,\bm{k}}\otimes\Phi_{\ell,i,\bm{k}}^{*}}\rangle\right)$ (96)
    $\geq\left(\frac{(dd^{\prime})^{T}}{d(d+1)\cdots(d+T-1)}\right)^{2}W_{\ell}\prod_{i=1}^{2}\left(\sum_{\bm{j},\bm{k}}\langle\Psi_{\ell,i,\bm{j}}|\Psi_{\ell,i,\bm{k}}\rangle\langle\Phi_{\ell,i,\bm{k}}|\Phi_{\ell,i,\bm{j}}\rangle\right)$ (97)
    $=\left(\frac{(dd^{\prime})^{T}}{d(d+1)\cdots(d+T-1)}\right)^{2}W_{\ell}\sum_{\bm{j},\bm{k}}\langle\Psi_{\ell,\bm{j}}|\Psi_{\ell,\bm{k}}\rangle\langle\Phi_{\ell,\bm{k}}|\Phi_{\ell,\bm{j}}\rangle$ (98)
    Thus, for each leaf $\ell$, we have
    $\frac{\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)}{p^{\mathcal{D}}(\ell)}$ $\geq\left(\frac{d^{T}}{d(d+1)\cdots(d+T-1)}\right)^{2}\geq\left(1-\frac{T^{2}}{d}\right)^{2}.$ (99)
    Therefore, we have
    $\left\lVert p^{\mathcal{D}}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}\leq 1-\left(1-\frac{T^{2}}{d}\right)^{2}\leq\frac{2T^{2}}{d}.$ (100)
Lemma 14 (Lemma 5.12 in [37] and Lemma 3 in [38]).
For any product vector $|{X}\rangle=\bigotimes_{t=1}^{k}|{x_{t}}\rangle$, we have
$\langle{X}|\sum_{\sigma\in\mathcal{S}_{k}}P_{\sigma}|{X}\rangle\geq\langle X|X\rangle.$ (101)
C.2 Coherent Access
We now consider the lower bound for completing DSEU with coherent access.
C.2.1 Proof of Theorem 4
We first establish some notations. Recall that any learning algorithms with coherent access can be described as follows. Consider two unknown unitary channels $U$ and $V$ acting on two separate quantum systems. On these devices, we input the initial state $|{0}\rangle\otimes|{0}\rangle$ on $\mathcal{H}^{\otimes 2}$ and interleave the channel queries with a sequence of $T$ adaptive data-processing channels $\mathcal{C}_{t}=\mathcal{C}_{t,1}\otimes\mathcal{C}_{t,2}$, $t=1,\cdots,T$. Assuming the auxiliary system is large enough, each data-processing channel can be represented as a unitary $W_{t}=W_{t,1}\otimes W_{t,2}$ [51]. After $T$ queries, the resulting states on the two devices can then be expressed as
$|{\psi_{1}}\rangle$ $=\prod_{t=1}^{T}\left[\left(U\otimes I_{\rm aux}\right)W_{t,1}\right]|{0}\rangle,\quad|{\psi_{2}}\rangle=\prod_{t=1}^{T}\left[\left(V\otimes I_{\rm aux}\right)W_{t,2}\right]|{0}\rangle.$ (102)
As shown in [47], they can also be written as
$|{\psi_{1}}\rangle=\left(I_{\mathcal{H}}\otimes\langle{\Phi}|\right)\left(I_{\mathcal{H}}\otimes U^{\otimes T}\otimes I_{\rm main}^{\otimes T}\right)|{\Psi_{I,1}}\rangle,\quad|{\psi_{2}}\rangle=\left(I_{\mathcal{H}}\otimes\langle{\Phi}|\right)\left(I_{\mathcal{H}}\otimes V^{\otimes T}\otimes I_{\rm main}^{\otimes T}\right)|{\Psi_{I,2}}\rangle,$ (103)
where $|{\Phi}\rangle=\sum_{i=0}^{d^{T}-1}|{ii}\rangle$ is the maximally entangled state on system $\mathcal{H}_{\rm main}^{\otimes T}\otimes\mathcal{H}_{\rm main}^{\otimes T}$ and
$|{\Psi_{I,i}}\rangle=\sum_{\begin{subarray}{c}x_{1},\cdots,x_{T}\in\{0,1\}^{n}\\ y_{1},\cdots,y_{T}\in\{0,1\}^{n}\end{subarray}}\prod_{t=1}^{T}\overbrace{\left[\left(|{y_{t}}\rangle\!\langle{x_{t}}|\otimes I_{\rm aux}\right)W_{t,i}\right]|{0}\rangle}^{\text{$\mathcal{H}$}}\otimes\overbrace{|{x_{1}\cdots x_{T}}\rangle\otimes|{y_{1}\cdots y_{T}}\rangle}^{\text{$\mathcal{H}_{\rm main}^{\otimes T}\otimes\mathcal{H}_{\rm main}^{\otimes T}$}}.$ (104)
Lastly, we can perform an adaptive POVM $\{E_{\ell}\mathrel{\mathop{\ordinarycolon}}=E_{\ell,1}\otimes E_{\ell,2}\}_{\ell}$. Based on the above definition, we can define the following two probabilities in the following form:
$p^{U,V}(\ell)$ $\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}\left[\left(E_{\ell,1}\otimes J_{U}^{(T)}\right)\Psi_{I,1}\right]\operatorname{Tr}\left[\left(E_{\ell,2}\otimes J_{V}^{(T)}\right)\Psi_{I,2}\right],$ (105)
$p^{U,U}(\ell)$ $\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}\left[\left(E_{\ell,1}\otimes J_{U}^{(T)}\right)\Psi_{I,1}\right]\operatorname{Tr}\left[\left(E_{\ell,2}\otimes J_{U}^{(T)}\right)\Psi_{I,2}\right].$ (106)
where $J_{U}^{(T)}$ is the Choi operator of unitary $U^{\otimes T}$, as defined in Eq. (12), and similar to $J_{V}^{(T)}$. With the definition of Haar random channel, we have
$\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)=\operatorname{Tr}\left[\left(E_{\ell,1}\otimes J_{H}^{(T)}\right)\Psi_{I,1}\right]\operatorname{Tr}\left[\left(E_{\ell,2}\otimes J_{H}^{(T)}\right)\Psi_{I,2}\right],$ (107)
where $J_{H}^{(T)}$ is defined in Eq. (18). As we do in the proof of lower bound for incoherent access, we define a intermediate probability
$p^{a}(\ell)\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}\left[\left(E_{\ell,1}\otimes J_{a}^{(T)}\right)\Psi_{I,1}\right]\operatorname{Tr}\left[\left(E_{\ell,2}\otimes J_{a}^{(T)}\right)\Psi_{I,2}\right],$ (108)
where $J_{a}^{(T)}$ is the Choi operator of $\mathcal{E}_{a}^{(T)}$, defined in Eq. (25). Then, with Lemma 15, we have
$\left\lVert\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)-\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)\right\rVert_{\rm TV}$ $\leq\left\lVert\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)-p^{a}(\ell)\right\rVert_{\rm TV}+\left\lVert p^{a}(\ell)-\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)\right\rVert_{\rm TV}$ (109)
$\leq\frac{T^{2}/2d}{1-T^{2}/2d}+\frac{4T^{2}}{d},$ (110)
which hints that we require $T=\Omega(\sqrt{d})$. Therefore, we complete the proof of Theorem 4.
C.2.2 Technical Lemmas
In the following, we prove key lemmas used in the above proof.
Lemma 15.
For two terms defined in Eq. (109), we have the following upper bound,
$\left\lVert\mathbb{E}p^{U,V\sim\mu_{H}}(\ell)-p^{a}(\ell)\right\rVert_{\rm TV}$ $\leq\frac{T^{2}/2d}{1-T^{2}/2d},\quad\left\lVert p^{a}(\ell)-\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)\right\rVert_{\rm TV}\leq\frac{4T^{2}}{d}.$ (111)
Proof of Lemma 15.
First, with the definition of $J_{a}^{(T)}$ (Eq. (25)), we have
$\quad\left|\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)-p^{a}(\ell)\right|$ (112)
$\leq\operatorname{Tr}\left[\left(E_{\ell_{1}}\otimes J_{H}^{(T)}\right){\Psi_{I,1}}\right]\left|\operatorname{Tr}\left[\left(E_{\ell_{2}}\otimes\left(J_{H}^{(a)}-J_{a}^{(T)}\right)\right){\Psi_{I,2}}\right]\right|+\left|\operatorname{Tr}\left[\left(E_{\ell_{1}}\otimes\left(J_{H}^{(T)}-J_{a}^{(T)}\right)\right){\Psi_{I,1}}\right]\right|\operatorname{Tr}\left[\left(E_{\ell_{2}}\otimes J_{a}^{(T)}\right){\Psi_{I,2}}\right]$ (113)
$\leq\frac{T^{2}/2d}{1-T^{2}/2d}\left\{\operatorname{Tr}\left[\left(E_{\ell_{1}}\otimes J_{H}^{(T)}\right){\Psi_{I,1}}\right]\operatorname{Tr}\left[\left(E_{\ell_{2}}\otimes J_{a}^{(T)}\right){\Psi_{I,2}}\right]+\operatorname{Tr}\left[\left(E_{\ell_{1}}\otimes J_{a}^{(T)}\right){\Psi_{I,1}}\right]\operatorname{Tr}\left[\left(E_{\ell_{2}}\otimes J_{H}^{(T)}\right){\Psi_{I,2}}\right]\right\}.$ (114)
Thus, we have
$\sum_{\ell}\left|\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V\sim\mu_{H}}(\ell)-p^{a}(\ell)\right|$ $\leq 2\frac{T^{2}/2d}{1-T^{2}/2d}\quad\Rightarrow\quad\left\lVert\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V\sim\mu_{H}}(\ell)-p^{a}(\ell)\right\rVert_{\rm TV}\leq\frac{T^{2}/2d}{1-T^{2}/2d}.$ (115)
For the second term, we have
$\left\lVert p^{a}(\ell)-\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)\right\rVert_{\rm TV}$ $=\sum_{\ell\mathrel{\mathop{\ordinarycolon}}p^{a}(\ell)\geq\mathbb{E}p^{U,U}(\ell)}p^{a}(\ell)\left[1-\frac{\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)}{p^{a}(\ell)}\right]$ (116)
$=\sum_{\ell\mathrel{\mathop{\ordinarycolon}}p^{a}(\ell)\geq\mathbb{E}p^{U,U}(\ell)}\operatorname{Tr}\left[M_{\ell_{1}}J_{a}^{(T)}\right]\operatorname{Tr}\left[M_{\ell_{2}}J_{a}^{(T)}\right]\left[1-\frac{\operatorname{Tr}\left[M_{\ell_{2}}\Phi_{\ell}^{(2T)}\right]}{\operatorname{Tr}\left[M_{\ell_{2}}J_{a}^{(T)}\right]}\right].$ (117)
Here we define a “measure-and-prepare” map
$\Phi_{\ell}^{(2T)}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mu_{H}}\frac{\operatorname{Tr}\left[M_{\ell_{1}}J_{U}^{(T)}\right]}{\operatorname{Tr}\left[M_{\ell_{1}}J_{a}^{(T)}\right]}J_{U}^{(T)},\quad M_{\ell_{i}}\mathrel{\mathop{\ordinarycolon}}=\Pi_{{\rm sym}}^{\prime(d,T)}\operatorname{Tr}_{1}\left[\left(E_{\ell_{i}}\otimes I\right)\Psi_{I,i}\right]\Pi_{{\rm sym}}^{\prime(d,T)},\quad i=1,2,$ (118)
where $\Pi_{{\rm sym}}^{\prime(d,T)}\mathrel{\mathop{\ordinarycolon}}=\sum_{\pi\in\mathcal{S}_{T}}P_{\pi}\otimes P_{\pi}/T!$ is the projector onto the symmetric subspace of $\mathcal{H}_{\rm main}^{\otimes T}\otimes\mathcal{H}_{\rm main}^{\otimes T}$ [47]. Thus, with the definition, $M_{\ell_{i}}$ is in the symmetric subspace of $\mathcal{H}_{\rm main}^{\otimes T}\otimes\mathcal{H}_{\rm main}^{\otimes T}$ and we have [48, 47]
$J_{a}^{(T)}=\frac{T!}{d^{T}}\Pi_{{\rm sym}}^{\prime(d,T)}\quad\Rightarrow\quad\frac{\operatorname{Tr}\left[M_{\ell_{2}}\Phi_{\ell}^{(2T)}\right]}{\operatorname{Tr}\left[M_{\ell_{2}}J_{a}^{(T)}\right]}=\frac{d^{2T}}{(T!)^{2}\operatorname{Tr}\left[M_{\ell_{1}}\right]\operatorname{Tr}\left[M_{\ell_{2}}\right]}\mathbb{E}_{U\sim\mu_{H}}\operatorname{Tr}\left[M_{\ell_{1}}J_{U}^{(T)}\right]\operatorname{Tr}\left[M_{\ell_{2}}J_{U}^{(T)}\right].$ (119)
With Lemma 16, we have
$\frac{\operatorname{Tr}\left[M_{\ell_{2}}\Phi_{\ell}^{(2T)}\right]}{\operatorname{Tr}\left[M_{\ell_{2}}J_{a}^{(T)}\right]}\geq\frac{d^{2T}}{d(d+1)\cdots(d+2T-1)}\geq 1-\frac{4T^{2}}{d}.$ (120)
Therefore, we have
$\left\lVert p^{a}(\ell)-\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)\right\rVert_{\rm TV}\leq\frac{4T^{2}}{d}.$ (121)
Lemma 16.
For quantum states $\rho,\sigma$ in the symmetric subspace of $\mathcal{H}^{\otimes T}\otimes\mathcal{H}^{\otimes T}$, we have
$\mathbb{E}_{U\sim\mu_{H}}\operatorname{Tr}\left[\rho\cdot J_{U}^{(T)}\right]\operatorname{Tr}\left[\sigma\cdot J_{U}^{(T)}\right]\geq\frac{(T!)^{2}}{d(d+1)\cdots(d+2T-1)}$ (122)
Figure 2: The visualization of key proof steps. (a) The visualization of permutation map $\mathcal{R}$ (defined in Eq. (123)) when $T=3$. (b) The visualization of of $\mathcal{R}^{\dagger}$ when $T=3$. (c) The visualization of Eq. (126) when $T=2$. (d) The visualization of Eq. (128) when $T=2$.
Proof of Lemma 16.
For convenience, we also provide the visualization of the key proof steps in Fig. 2. Now, we begin our proof. First, we define the following permutation map,
$\mathcal{R}\mathrel{\mathop{\ordinarycolon}}\mathcal{D}\left(\mathcal{H}^{\otimes T}\otimes\mathcal{H}^{\otimes T}\right)\to\mathcal{D}\left((\mathcal{H}^{\otimes 2})^{\otimes T}\right),\quad|{\alpha}\rangle^{\otimes T}\otimes|{\beta}\rangle^{\otimes T}\mapsto\left(|{\alpha}\rangle\otimes|{\beta}\rangle\right)^{\otimes T},$ (123)
which can map a state in the symmetric subspace of $\mathcal{H}^{\otimes T}\otimes\mathcal{H}^{\otimes T}$ to the symmetric subspace of $(\mathcal{H}^{\otimes 2})^{\otimes T}$. This map $\mathcal{R}$ is visualized in Fig 2(a). With this definition, we have
$\mathcal{R}\left(\Pi_{\rm sym}^{\prime(d,T)}\right)=\Pi_{\rm sym}^{(d^{2},T)},\quad\mathcal{R}\left(J_{U}^{(T)}\right)=\left[(U\otimes I)\Phi(U\otimes I)^{\dagger}\right]^{\otimes T}.$ (124)
Then, we can observe that the space of density matrices on the symmetric subspace of $(\mathcal{H}^{\otimes 2})^{\otimes T}$ can is spanned by $\alpha^{\otimes T}$, where $\alpha\in\mathcal{D}(\mathcal{H}^{\otimes 2})$ [52, 48]. Thus, to compute the lower bound, it suffices to calculate the following function
$f(\alpha,\beta)$ $\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mu_{H}}\operatorname{Tr}\left[\beta^{\otimes T}\mathcal{R}\left(J_{U}^{(T)}\right)\right]\operatorname{Tr}\left[\alpha^{\otimes T}\mathcal{R}\left(J_{U}^{(T)}\right)\right]=\mathbb{E}_{U\sim\mu_{H}}\operatorname{Tr}\left[\left(\mathcal{R}^{\dagger}(\alpha^{\otimes T})\otimes\mathcal{R}^{\dagger}(\beta^{\otimes T})\right)\left(J_{U}^{(T)}\otimes J_{U}^{(T)}\right)\right],$ (125)
where $\mathcal{R}^{\dagger}$ is the inverse map of $\mathcal{R}$ and is visualized in Fig. 2(b). With the permutation operations shown in Fig. 2(c), we can write $f(\alpha,\beta)$ as
$f(\alpha,\beta)$ $=\operatorname{Tr}\left[\mathbb{F}_{2,3}\left(\mathcal{R}^{\dagger}(\alpha^{\otimes T})\otimes\mathcal{R}^{\dagger}(\beta^{\otimes T})\right)\mathbb{F}_{2,3}\left(\sum_{\pi,\sigma\in\mathcal{S}_{2T}}{\rm Wg}_{\pi,\sigma}P_{\pi}\otimes P_{\sigma}\right)\right],$ (126)
where $\mathbb{F}_{2,3}$ is acting on the second and third $\mathcal{H}^{\otimes T}$. With Lemma 8, we have
$f(\alpha,\beta)$ $\geq\frac{1}{d(d+1)\cdots(d+2T-1)}\operatorname{Tr}\left[\mathbb{F}_{2,3}\left(\mathcal{R}^{\dagger}(\alpha^{\otimes T})\otimes\mathcal{R}^{\dagger}(\beta^{\otimes T})\right)\mathbb{F}_{2,3}\left(\sum_{\pi\in\mathcal{S}_{2T}}P_{\pi}\otimes P_{\pi}\right)\right]$ (127)
$=\frac{(2T!)}{d(d+1)\cdots(d+2T-1)}\operatorname{Tr}\left[\Pi_{\rm sym}^{(d^{2},2T)}\left(\alpha^{\otimes T}\otimes\beta^{\otimes T}\right)\right]$ (128)
$=\frac{(2T!)}{d(d+1)\cdots(d+2T-1)}\sum_{s=0}^{T}\frac{\binom{T}{s}^{2}}{\binom{2T}{T}}x^{s}\geq\frac{(T!)^{2}}{d(d+1)\cdots(d+2T-1)},$ (129)
where $x\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}[\alpha\beta]$ and Eq. (128) is visualized in Fig. (2)(d). The term on the last line is the probability that a random $\pi\in\mathcal{S}_{2T}$ satisfies $|\pi(\{1,\cdots,T\})\cap\{1,\cdots,T\}|=s$ [48]. The corresponding physical meaning is that map the first $T$ $\alpha$ terms to $2T$ positions, the overlap term will be $x$, the other terms are all $1$. This is equivalent to the probability that when $T$ balls are drawn without replacement from a bucket of $T$ white balls and $T$ black balls, that the resulting sample contains TsT-s white balls and $s$ black balls. ∎
Appendix D Similarity Estimation with Independent Classical Shadow
In this section, we consider completing DSEU with independent classical shadow estimation for unitary channels (CSEU) [34]. In Appendix D.1, we describe the algorithm to complete DSEU and analyze the corresponding query complexity given in Theorem 5 of the main text. In Appendix D.2, we summarize several lemmas used in the proof.
D.1 Algorithm & Proof of Theorem 5
We first briefly introduce CSEU. Given a unknown unitary channel $\mathcal{U}$, we can obtain the following classical snapshot by inputting state $|{\psi}\rangle^{\otimes s}$ and performing POVM $\mathcal{M}_{s}$ (defined in Eq. (5)),
$\tilde{X}={\rm Snap}(\psi,\phi,s)\mathrel{\mathop{\ordinarycolon}}=\frac{d(d+1)(d+s){\phi}\otimes{\psi}^{T}-(d+1+s)(I\otimes I)}{s},$ (130)
where $|{\psi}\rangle$ is sampled from a state $4$-design ensemble, and $|{\phi}\rangle$ is the measurement result of POVM $\mathcal{M}_{s}$ . As shown in [34, Eq. (10)], the expectation of $\tilde{X}$ is Choi operator of unitary $U$, i.e.,
$\mathbb{E}\tilde{X}=J_{U}.$ (131)
See more details regarding CSEU in [34].
We can complete DSEU via CSEU as follows. First, we independently perform CSEU on each quantum devices, and obtain two set of classical snapshots of $\mathcal{U},\mathcal{V}$, labeled as $\{\tilde{X}\}$ and $\{\tilde{Y}\}$. Then, we can estimate the similarity $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$ with these independent classical snapshots. We summarize this algorithm in Algorithm 3. In the following, we prove that this algorithm can indeed complete DSEU and analyze its query complexity.
Input: number of SPAM settings $T$,
    the size of symmetric collective measurement $s$,
    TsTs queries of unknown unitary channels $\mathcal{U}$ and $\mathcal{V}$ acting on $d$-dimension Hilbert space $\mathcal{H}$.
Output: an estimation of $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$.
1 for $t=1,\cdots,T$ do
2
3  Randomly generate pure states $|{\psi_{t,A}}\rangle^{\otimes s}$ and apply the unitary channels $\mathcal{U}^{\otimes s}$.
4  Measure $(\mathcal{U}(|{\psi_{t,A}}\rangle\!\langle{\psi_{t,A}}|))^{\otimes s}$ with the POVM $\mathcal{M}_{s}$ and obtains result $|{\phi_{t,A}}\rangle$.
5  Randomly generate pure states $|{\psi_{t,B}}\rangle^{\otimes s}$ and apply the unitary channels $\mathcal{V}^{\otimes s}$.
6  Measure $(\mathcal{V}(|{\psi_{t,B}}\rangle\!\langle{\psi_{t,B}}|))^{\otimes s}$ with the POVM $\mathcal{M}_{s}$ and obtains result $|{\phi_{t,B}}\rangle$.
7  Compute and store the classical snapshots with Eq. (130):
$\tilde{X}_{t}={\rm Snap}\left(\phi_{t,A},\psi_{t,A},s\right),\quad\tilde{Y}_{t}={\rm Snap}\left(\phi_{t,B},\psi_{t,B},s\right).$ (132)
Return
$\tilde{\gamma}=\frac{1}{T^{2}d^{2}}\sum_{i,j=1}^{T}\operatorname{Tr}\left[\tilde{X}_{i}^{\dagger}\tilde{Y}_{j}\right].$ (133)
Algorithm 3 Distributed Similarity Estimation for Unitary Channels with CSEU
First, we prove that $\tilde{\gamma}$ is an unbiased estimator of $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$:
$\mathbb{E}\tilde{\gamma}$ $=\frac{1}{d^{2}}\mathbb{E}\operatorname{Tr}\left[\tilde{X}^{\dagger}\tilde{Y}\right$] (134)
$=\frac{1}{d^{2}}\operatorname{Tr}\left[J_{\mathcal{U}}^{\dagger}J_{\mathcal{V}}\right]$ (135)
$=\frac{1}{d^{2}}\sum_{ijkl}\operatorname{Tr}\left[\left(U|{j}\rangle\!\langle{i}|U^{\dagger}\otimes|{j}\rangle\!\langle{i}|\right)\left(V|{k}\rangle\!\langle{l}|V^{\dagger}\otimes|{k}\rangle\!\langle{l}|\right)\right]$ (136)
$=\frac{1}{d^{2}}\sum_{ij}\operatorname{Tr}\left[U|{j}\rangle\!\langle{i}|U^{\dagger}V|{i}\rangle\!\langle{j}|V^{\dagger}\right]$ (137)
$=\frac{\operatorname{Tr}^{2}\left[U^{\dagger}V\right]}{d^{2}}.$ (138)
Then, we analyze the query complexity of this classical shadow based algorithm. The variance of estimator $\tilde{\gamma}$ is given by
[…]
$=\frac{1}{T^{4}d^{4}}\mathbb{E}\left(\sum_{i,j,k,l}^{T}\operatorname{Tr}\left[\tilde{X}_{i}^{\dagger}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}^{\dagger}\tilde{Y}_{l}\right]\right)-\frac{\operatorname{Tr}^{4}[U^{\dagger}V]}{d^{4}}.$ (140)
After expanding the expectation in the above equation, there are four terms as follows.
  1. 1.
    $i=k$ and $j=l$, there are $T^{2}$ terms and with Lemma 18, we have
    $\mathbb{E}\left(\sum_{i,j}\operatorname{Tr}^{2}\left[\tilde{X}_{i}^{\dagger}\tilde{Y}_{j}\right]\right)=T^{2}\mathbb{E}\operatorname{Tr}^{2}\left[X^{\dagger}Y\right]=T^{2}\mathcal{O}\left(\frac{d^{4}(d+s)^{2}}{s^{2}}\right).$ (141)
  2. 2.
    For $i\neq k$ and $j\neq l$ case, there are $T^{2}(T-1)^{2}$ terms and we have
    $\mathbb{E}\left(\sum_{i\neq k,j\neq l}\operatorname{Tr}\left[\tilde{X}_{i}^{\dagger}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}^{\dagger}\tilde{Y}_{l}\right]\right)=T^{2}(T-1)^{2}\operatorname{Tr}^{4}[U^{\dagger}V].$ (142)
  3. 3.
    For $i=k$ and $j\neq l$ case, there are $T^{2}(T-1)$ terms and with Lemma 19 we have
    $\mathbb{E}\left(\sum_{i=k,j\neq l}\operatorname{Tr}\left[\tilde{X}_{i}^{\dagger}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}^{\dagger}\tilde{Y}_{l}\right]\right)=T^{2}(T-1)\mathbb{E}\operatorname{Tr}\left[X^{\dagger}\tilde{Y}_{1}\right]\operatorname{Tr}\left[X^{\dagger}\tilde{Y}_{2}\right]\leq\mathcal{O}\left(T^{3}d^{4}\right).$ (143)
  4. 4.
    For $i\neq k$ and $j=l$ case, there are $T^{2}(T-1)$ terms and with Lemma 19 we have
    $\mathbb{E}\left(\sum_{i\neq k,j=l}\operatorname{Tr}\left[\tilde{X}_{i}^{\dagger}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}^{\dagger}\tilde{Y}_{l}\right]\right)=T^{2}(T-1)\mathbb{E}\operatorname{Tr}\left[\tilde{X}_{1}^{\dagger}Y\right]\operatorname{Tr}\left[\tilde{X}_{2}^{\dagger}Y\right]\leq\mathcal{O}\left(T^{3}d^{4}\right).$ (144)
Therefore, we have
[…]
$\leq\mathcal{O}\left(\frac{(d+s)^{2}}{T^{2}s^{2}}+\frac{1}{T}\right)=\mathcal{O}\left(\frac{d^{2}}{T^{2}s^{2}}+\frac{1}{T}\right).$ (146)
Thus, to achieve $\varepsilon$ additive error, we require the query times satisfy
$Ts=\mathcal{O}\left(\max\left\{\frac{d}{\varepsilon},\frac{1}{\varepsilon^{2}}\right\}\right).$ (147)
Therefore, we complete the proof of Theorem 5.
D.2 Technical Lemmas
Lemma 17 (Lemma D1 in [34]).
Suppose that $\phi$ and $\psi$ are random input states and measurement outcome of learning unitary channel $\mathcal{U}$, then we have
$\mathbb{E}\phi^{\otimes 2}\otimes\psi^{\otimes 2}$ $=\frac{2}{(d+s)(d+s+1)}\sum_{i=1}^{4}\Delta_{U,i},$ (148)
where
$\Delta_{U,1}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{1}{\kappa_{2}}\Pi_{\rm sym}^{(2)}\otimes\Pi_{\rm sym}^{(2)},$ (149)
$\Delta_{U,2}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{s}{\kappa_{3}}(I\otimes U\otimes I\otimes I)\left[I_{1}\otimes\left(\Pi_{\rm sym}^{(3)}\right)_{2,3,4}\right](I\otimes U^{\dagger}\otimes I\otimes I)\left(\Pi_{\rm sym}^{(2)}\otimes I\otimes I\right),$ (150)
$\Delta_{U,3}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{s}{\kappa_{3}}(U\otimes I\otimes I\otimes I)\left[I_{2}\otimes\left(\Pi_{\rm sym}^{(3)}\right)_{1,3,4}\right](U^{\dagger}\otimes I\otimes I\otimes I)\left(\Pi_{\rm sym}^{(2)}\otimes I\otimes I\right),$ (151)
$\Delta_{U,4}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{s(s-1)}{2\kappa_{4}}(U\otimes U\otimes I\otimes I)\Pi_{\rm sym}^{(4)}(U^{\dagger}\otimes U^{\dagger}\otimes I\otimes I).$ (152)
Lemma 18.
Suppose $X$ and $Y$ are classical snapshot defined in Eq. (130) with POVM $\mathcal{M}_{s}$ for unitary channels $\mathcal{U}$ and $\mathcal{V}$, we have
$\mathbb{E}\operatorname{Tr}^{2}\left[X^{\dagger}Y\right]=\mathcal{O}\left(\frac{d^{4}(d+s)^{2}}{s^{2}}\right).$ (153)
Proof.
With Lemma 10 and the definition of $X$, $Y$, we have
$\mathbb{E}\operatorname{Tr}^{2}\left[X^{\dagger}Y\right]=\frac{1}{s^{2}}\operatorname{Tr}^{2}\left[\left(d(d+1)(d+s){\phi_{1}}\otimes{\psi_{1}}^{T}-(d+1+s)I\otimes I\right)\left(d(d+1)(d+s){\phi_{2}}\otimes{\psi_{2}}^{T}-(d+1+s)I\otimes I\right)\right]$ (154)
$=\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{1}\phi_{2}\otimes\psi_{2}\psi_{2}\right]-\frac{d(d+1)(d+s)(d+1+s)}{s^{2}}\left(\sum_{i=1}^{2}\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{i}\otimes\psi_{i}\right]\right)+\frac{d^{4}(d+1+s)^{2}}{s^{2}}.$ (155)
We consider the first two terms respectively as follows.
  1. 1.
    For the first term, with Lemma 17, there are $16$ terms,
    $\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{1}\phi_{2}\otimes\psi_{2}\psi_{2}\right]$ $=\mathbb{E}\operatorname{Tr}\left[\left(\phi_{1}^{\otimes 2}\otimes\psi_{1}^{\otimes 2}\right)\left(\phi_{2}^{\otimes 2}\otimes\psi_{2}^{\otimes 2}\right)\right]=\frac{4}{(d+s)^{2}(d+s+1)^{2}}\sum_{k,l=1}^{4}\operatorname{Tr}\left[\Delta_{U,k}\Delta_{V,l}\right]$ (156)
    We analyze these terms as follows
    $\operatorname{Tr}[\Delta_{U,1}\Delta_{V,1}]$ $=\frac{\operatorname{Tr}^{2}\left[\Pi_{\rm sym}^{(2)}\right]}{\kappa_{2}^{2}}=1,$ (157)
    $\operatorname{Tr}[\Delta_{U,1}\Delta_{V,2}]$ $=\operatorname{Tr}[\Delta_{U,2}\Delta_{V,1}]=\operatorname{Tr}[\Delta_{U,1}\Delta_{V,3}]=\operatorname{Tr}[\Delta_{U,3}\Delta_{V,1}]=\frac{s}{\kappa_{2}\kappa_{3}}\frac{\kappa_{2}\kappa_{3}}{d}=\frac{s}{d},$ (158)
    $\operatorname{Tr}\left[\Delta_{U,1}\Delta_{V,4}\right]$ $=\operatorname{Tr}\left[\Delta_{U,4}\Delta_{V,1}\right]=\frac{s(s-1)}{d(d+1)}=\mathcal{O}\left(\frac{s^{2}}{d^{2}}\right),$ (159)
    $\operatorname{Tr}\left[\Delta_{U,2}\Delta_{V,2}\right]$ $=\operatorname{Tr}\left[\Delta_{U,3}\Delta_{V,3}\right]=\frac{s^{2}(d^{2}+2d+\operatorname{Tr}^{2}[U^{\dagger}V])}{d^{2}(d+1)^{2}}=\mathcal{O}\left(\frac{s^{2}}{d^{2}}\right),$ (160)
    $\operatorname{Tr}\left[\Delta_{U,2}\Delta_{V,3}\right]$ $=\operatorname{Tr}\left[\Delta_{U,3}\Delta_{V,2}\right]=\frac{s^{2}(d^{2}+2d+\operatorname{Tr}^{2}[U^{\dagger}V])}{d^{2}(d+1)^{2}}=\mathcal{O}\left(\frac{s^{2}}{d^{2}}\right),$ (161)
    $\operatorname{Tr}\left[\Delta_{U,2}\Delta_{V,4}\right]$ $\leq\frac{s^{2}(s-1)d^{4}}{2\kappa_{3}\kappa_{4}}=\mathcal{O}\left(\frac{s^{3}}{d^{3}}\right),\quad\operatorname{Tr}\left[\Delta_{U,4}\Delta_{V,2}\right]\leq\mathcal{O}\left(\frac{s^{3}}{d^{3}}\right),$ (162)
    $\operatorname{Tr}\left[\Delta_{U,3}\Delta_{V,4}\right]$ $\leq\mathcal{O}\left(\frac{s^{3}}{d^{3}}\right),\quad\operatorname{Tr}\left[\Delta_{U,4}\Delta_{V,3}\right]\leq\mathcal{O}\left(\frac{s^{3}}{d^{3}}\right),\quad\operatorname{Tr}\left[\Delta_{U,4}\Delta_{V,3}\right]\leq\frac{s^{2}(s-1)^{2}d^{4}}{4\kappa_{4}}=\mathcal{O}\left(\frac{s^{4}}{d^{4}}\right).$ (163)
    Therefor, for the first term, we have
    $\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{1}\phi_{2}\otimes(\psi_{2}\psi_{2})^{T}\right]$ $\leq\frac{4d^{2}(d+1)^{2}}{s^{2}(d+s+1)^{2}}\mathcal{O}\left(1+\frac{s}{d}+\frac{s^{2}}{d^{2}}+\frac{s^{3}}{d^{3}}+\frac{s^{4}}{d^{4}}\right)$ (164)
    $=\mathcal{O}\left(\frac{d^{4}}{s^{2}(d+s)^{2}}+\frac{s^{2}}{(d+s)^{2}}\right)$ (165)
  2. 2.
    For the second term, with Lemma 17, we have
    $\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{1}\otimes\psi^{T}_{1}\right]$ $=\mathbb{E}\operatorname{Tr}\left[\phi_{1}^{\otimes 2}\otimes\psi_{1}^{\otimes 2}\right]=\frac{2}{(d+s)(d+s+1)}\sum_{i=1}^{4}\operatorname{Tr}\left[\Delta_{U,i}\right]$ (166)
    $=\frac{2}{(d+s)(d+s+1)}\left[\kappa_{2}+\frac{2\cdot d(d+1)^{2}(d+2)}{12\cdot\kappa_{3}}+\frac{s(s-1)}{2}\right]=\mathcal{O}(1).$ (167)
    Likewise, we have $\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{1}\otimes\psi^{T}_{1}\right]=\mathcal{O}(1)$.
Therefore, we have
$\mathbb{E}\operatorname{Tr}^{2}\left[X^{\dagger}Y\right]$ $\leq\mathcal{O}\left(\frac{d^{4}}{s^{2}(d+s)^{2}}+\frac{s^{2}}{(d+s)^{2}}\right)+\frac{d^{4}(d+1+s)^{2}}{s^{2}}=\mathcal{O}\left(\frac{d^{4}(d+s)^{2}}{s^{2}}\right).$ (168)
Lemma 19.
[…]
$\mathbb{E}\operatorname{Tr}\left[X_{1}^{\dagger}Y\right]\operatorname{Tr}\left[X_{2}^{\dagger}Y\right]$ $\leq\mathcal{O}\left(d^{4}\right).$ (169)
[…]
$\mathbb{E}\operatorname{Tr}\left[X^{\dagger}Y_{1}\right]\operatorname{Tr}\left[X^{\dagger}Y_{2}\right]$ $\leq\mathcal{O}\left(d^{4}\right).$ (170)
Proof.
With the definition in Eq. (130), we have
$\;\mathbb{E}\operatorname{Tr}\left[X_{1}^{\dagger}Y\right]\operatorname{Tr}\left[X_{2}^{\dagger}Y\right]=\operatorname{Tr}\left[\left(\mathbb{E}X\right)^{\otimes 2}\mathbb{E}Y^{\otimes 2}\right]$ (171)
$=\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\operatorname{Tr}\left[J_{U}^{\otimes 2}\mathbb{E}\left(\phi_{2}\otimes\psi_{2}^{T}\right)^{\otimes 2}\right]-\frac{2d^{2}(d+1)(d+s)(d+s+1)}{s^{2}}\operatorname{Tr}\left[J_{U}\mathbb{E}\phi_{2}\otimes\psi_{2}^{T}\right]+\frac{d^{2}(d+s+1)^{2}}{s^{2}}.$ (172)
With Lemma 1 in [34], we have
$\operatorname{Tr}\left[J_{U}\mathbb{E}\phi_{2}\otimes\psi_{2}^{T}\right]$ $=\frac{d+s+1}{(d+1)(d+s)}+\frac{s}{d(d+1)(d+s)}\operatorname{Tr}\left[J_{U}(V\otimes V^{\dagger})\mathbb{F}\right]$ (173)
$=\frac{d+s+1}{(d+1)(d+s)}+\frac{s}{d(d+1)(d+s)}\operatorname{Tr}^{2}\left[U^{\dagger}V\right].$ (174)
Thus, we have
$\mathbb{E}\operatorname{Tr}\left[X_{1}^{\dagger}Y\right]\operatorname{Tr}\left[X_{2}^{\dagger}Y\right]$ $=\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\operatorname{Tr}\left[J_{U}^{\otimes 2}\mathbb{E}\left(\phi_{2}\otimes\psi_{2}^{T}\right)^{\otimes 2}\right]-\frac{2d(d+s+1)}{s}\operatorname{Tr}^{2}\left[U^{\dagger}V\right]-\frac{d^{2}(d+s+1)^{2}}{s^{2}}.$ (175)
Now, we focus on the first term. With Lemma 17, we have
$\operatorname{Tr}\left[J_{U}^{\otimes 2}\mathbb{E}\left(\phi_{2}\otimes\psi_{2}^{T}\right)^{\otimes 2}\right]$ $=\sum_{i,j,k,l=0}^{d-1}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\mathbb{E}\phi^{\otimes 2}_{2}\otimes\psi^{\otimes 2}_{2}\right]$ (176)
$=\frac{2}{(d+s)(d+s+1)}\sum_{i=1}^{4}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\Delta_{V,i}\right]\leq\mathcal{O}\left(\frac{s^{2}}{(d+s)(d+s+1)}\right),$ (177)
with the following calculation,
$\sum_{i,j,k,l}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\Delta_{V,1}\right]$ $=\frac{s}{\kappa_{2}}\sum_{i,j,k,l}\operatorname{Tr}^{2}\left[\Pi_{\rm sym}^{(2)}|{jl}\rangle\!\langle{ik}|\right]=\frac{sd(d+1)}{2\cdot\kappa_{2}}=s,$ (178)
$\sum_{i,j,k,l}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\Delta_{V,2}\right]$ $=\frac{2s\left(\operatorname{Tr}^{2}[U^{\dagger}V]+d\right)(d+2)}{12\cdot\kappa_{3}}=\frac{s\operatorname{Tr}^{2}[U^{\dagger}V]}{d(d+1)}+\frac{s}{d+1}\leq s,$ (179)
$\sum_{i,j,k,l}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\Delta_{V,3}\right]$ $=\frac{2s\left(\operatorname{Tr}^{2}[U^{\dagger}V]+d\right)(d+2)}{12\cdot\kappa_{3}}=\frac{s\operatorname{Tr}^{2}[U^{\dagger}V]}{d(d+1)}+\frac{s}{d+1}\leq s,$ (180)
$\sum_{i,j,k,l}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\Delta_{V,4}\right]$ $\leq\frac{s(s-1)}{2}.$ (181)
Therefore, we have
$\mathbb{E}\operatorname{Tr}\left[X_{1}^{\dagger}Y\right]\operatorname{Tr}\left[X_{2}^{\dagger}Y\right]$ $\leq\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\mathcal{O}\left(\frac{s^{2}}{(d+s)(d+s+1)}\right)=\mathcal{O}\left(d^{4}\right).$ (182)
Likewise for $\mathbb{E}\operatorname{Tr}[X^{\dagger}Y_{1}]\operatorname{Tr}[X^{\dagger}Y_{2}]$. ∎