### v2

# Optimal Distributed Similarity Estimation of Quantum Channels

Abstract
We study distributed similarity estimation of quantum channels (DSEC), a primitive for cross-platform verification where two remote quantum devices are compared by estimating the inner product of their Choi states. We show that the optimal channel query complexity of DSEC for two $d$-dimensional quantum channels is $\Theta(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$, where $\varepsilon$ is the additive error. We first prove an information-theoretic lower bound with this scaling, which holds even in the strongest setting, allowing adaptive strategies, multiple rounds of classical communication, and coherent access with arbitrary ancillas. We then give a matching upper bound in the weakest setting, namely non-adaptive and ancilla-free incoherent access, via a randomized measurement protocol achieving this bound. Finally, we show that our protocol achieves a quadratic improvement over classical shadow baselines. Our results provide theoretically optimal and practical methods for cross-platform verification, quantum device benchmarking, and distributed quantum learning.
Introduction.— The engineering and physical realization of quantum computers are being actively pursued across a wide range of physical platforms [1, 2, 3]. To certify their performance, numerous protocols have been developed that compare experimentally generated quantum states or channels with known theoretical targets, including direct fidelity estimation [4, 5], randomized benchmarking [6, 7, 8], and quantum verification [9, 10, 11, 12]. However, these protocols rely on knowledge of the theoretical target, which may not always be available. A central challenge is how to directly compare unknown quantum states or channels produced on different platforms, possibly at different times and locations. This task, known as cross-platform verification, becomes especially pressing as we approach regimes of quantum advantage, where classical simulation of quantum systems becomes computationally intractable. For quantum states, the core step of cross-platform verification is distributed inner product estimation (DIPE). A variety of efficient algorithms have been developed for this task [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Notably, for $d$-dimensional quantum states, it has been shown that the sample complexity of completing this task is $\Theta(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$ [15], where $\varepsilon$ is the additive error.
Building upon the success of DIPE, a natural next question is cross-platform verification of quantum channels. The corresponding core task is estimating the inner product of their Choi states using only local operations and classical communication (LOCC), which we term distributed similarity estimation of quantum channels (DSEC). DSEC has attracted significant attention in several recent works [24, 25, 23], due to important applications in quantum information theory, including circuit equivalence checking [26, 27, 28] and quantum channel benchmarking [29, 30, 31, 32]. Despite its importance, fundamental questions about DSEC remain open: What is the lower bound for completing DSEC? Can we design optimal algorithms that achieve this bound? Several existing techniques can potentially address this task, including process tomography, extension of DIPE, and classical shadow for channels. However, all of these approaches face significant limitations. Process tomography is highly resource-intensive, requiring $\mathcal{O}(d^{4})$ queries [33, 34]. Directly extending DIPE by preparing Choi states and then applying DIPE algorithms leads to $\mathcal{O}(d)$ query complexity [15] and is experimentally challenging due to the need for preparing maximally entangled states. Classical shadow for channels, while sample-efficient in many cases [35, 36, 37], is far from optimal as we will discuss later.
In this work, we settle the optimal query complexity of DSEC by proving the lower bound and constructing an estimation protocol with matching upper bound based on randomized measurements [38]. First, we establish an information-theoretic lower bound, proving that any algorithm for DSEC must require at least $\Omega(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$ queries. Remarkably, this bound holds even in the strongest setting: multi-round LOCC, arbitrarily large ancillary systems, coherent access, and adaptive strategies. We further show that this lower bound is achievable. For unitary channels, we provide a coherent algorithm with query complexity $\mathcal{O}(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$. For general channels, we provide an algorithm requiring only non-adaptive strategies, incoherent access, and ancilla-free protocols. Surprisingly, by leveraging shared randomness between devices, this incoherent algorithm also achieves the same query complexity of $\mathcal{O}(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$. This highlights the power of shared randomness in distributed quantum learning and shows that the query complexity of DSEC is exactly $\Theta(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$. Finally, we compare our algorithms with independent classical shadow for channels [35, 36, 37]. We show that independent classical shadow requires $\mathcal{O}(\max\{d/\varepsilon,1/\varepsilon^{2}\})$ queries, whereas our protocols achieve a quadratic improvement. This square-root advantage arises from the shared randomness.
Figure 1: Learning models for distributed channel learning. (a) Learning access: (i) Incoherent access: the unknown channel is queried once before each measurement. (ii) Coherent access: the unknown channel is queried $T$ times before each measurement, with arbitrary intermediate quantum channels $\{\mathcal{C}_{t}\}$. (b) Shared randomness: (i) for both state preparation and measurement (SPAM) settings. (ii) only for state preparation. (iii) no shared randomness.
Problem Setup.— We now formally define the DSEC problem. Let $\mathcal{H}$ be the Hilbert space of a $d$-dimensional system. Let $\mathcal{E}$ and $\mathcal{F}$ be two unknown channels implemented on different (possibly distant) quantum devices. Define the Choi operator of $\mathcal{E}$ as $J_{\mathcal{E}}\mathrel{\mathop{\ordinarycolon}}=(\mathcal{E}\otimes\mathcal{I})(|{\Phi}\rangle\!\langle{\Phi}|)$, where $\mathcal{I}$ is the identity channel and $|{\Phi}\rangle\mathrel{\mathop{\ordinarycolon}}=\sum|{ii}\rangle$ is the unnormalized maximally entangled state in $\mathcal{H}^{\otimes 2}$ [39], and similarly for $J_{\mathcal{F}}$. The goal of DSEC is to estimate the inner product $\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}]/d^{2}\in[0,1]$ up to additive error $\varepsilon\in(0,1)$ with success probability at least $2/3$. This inner product can quantify the similarity between $\mathcal{E}$ and $\mathcal{F}$ via the max fidelity [24, 40]:
$\frac{\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}]}{\max\{\operatorname{Tr}[J_{\mathcal{E}}^{2}],\operatorname{Tr}[J_{\mathcal{F}}^{2}]\}},$ (1)
where $\operatorname{Tr}[J_{\mathcal{E}}^{2}]$, $\operatorname{Tr}[J_{\mathcal{F}}^{2}]$ describe the unitarity of $\mathcal{E}$, $\mathcal{F}$. Unitarity estimation has been considered in [41], and our algorithm can also serve as a new tool for it. In the following, we describe the learning models relevant to DSEC, focusing on two aspects: learning access and shared randomness.
Learning Access.
Following the classifications in [42, 43, 44, 41, 37, 45], we consider two types of learning access, illustrated in Fig. 1(a). (i) Incoherent access, also known as learning without quantum memory: In each query, the learner prepares an arbitrary quantum state, applies the unknown channel, and immediately measures the output state. The key restriction is that the unknown channel is queried exactly once before each measurement. (ii) Coherent access, also known as learning with quantum memory: In each experimental round, the learner prepares an arbitrary quantum state, may interleave multiple queries of the unknown channel with arbitrary intermediate quantum channels, and then applies an arbitrary measurement on the final output state. This allows quantum information to be preserved across multiple channel queries.
Shared Randomness.
Beyond learning access, shared randomness plays a crucial role in distributed learning with randomized measurements [38]. In many learning protocols, state preparation and measurement (SPAM) settings are essential components. Within the framework of randomized measurements, both initial states and measurement settings are sampled randomly to enhance practicality and sample efficiency. In the distributed scenario, allowing devices to share part or all of this randomness can significantly improve performance, as demonstrated in DIPE [15]. We consider three levels of shared randomness, illustrated in Fig. 1(b): (i) shared randomness for both SPAM settings, (ii) shared randomness only for state preparation, and (iii) no shared randomness between devices. A key finding of this work is that shared randomness provides substantial advantages in completing DSEC.
Lower Bound.— We now establish the lower bound for completing DSEC. Our strategy is to reduce DSEC to a simpler hypothesis testing problem and analyze its query complexity. Specifically, we consider the following distinguishing problem, which serves as a restricted version of DSEC.
Problem 1.
Distinguish between the following two cases:
Two devices perform the same unitary $U$, which is a Haar random unitary;
Two devices independently perform two unitaries $U$ and $V$, which are independent Haar random unitaries.
Let $J_{U}$ and $J_{V}$ denote the Choi operators of $U$ and $V$, respectively. If a learning algorithm can estimate the value of $\operatorname{Tr}[J_{U}J_{V}]/d^{2}$, then we can use this algorithm to complete Problem 1. Hence, a lower bound for completing Problem 1 directly implies a lower bound for completing DSEC. Moreover, this lower bound also applies to estimating other commonly used distance measures between quantum channels. In the following, we analyze the query complexity required to complete Problem 1 with incoherent and coherent accesses.
To analyze the query complexity of Problem 1, we employ Le Cam’s two-point method [46], which upper bounds the success probability of hypothesis testing in terms of the total variation distance (TVD) between the probability distributions of measurement outcomes under the two hypotheses. Let $p_{1}(\ell)$ and $p_{2}(\ell)$ be the probabilities of obtaining measurement result $\ell$ after querying the channels in cases (i) and (ii), respectively. Their TVD is defined as
$\left\lVert p_{1}-p_{2}\right\rVert_{\rm TV}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{2}\sum_{\ell}\left|p_{1}(\ell)-p_{2}(\ell)\right|.$ (2)
For reliable hypothesis testing, this TVD must be at least constantly large. Now, we are ready to prove the lower bounds for both learning access models.
Incoherent Access.
We first analyze the case of incoherent access, following the learning tree framework [44, 41, 47, 48, 49]. We show that incoherent access requires exponentially many queries in $\sqrt{d}$, even with multi-round LOCC, arbitrarily large ancillary systems, and adaptive SPAM settings.
Theorem 1.
Any algorithm with incoherent access, possibly utilizing multi-round LOCC, ancillary systems, and adaptive strategies, requires $\Omega(\sqrt{d})$ queries to complete Problem 1.
This theorem implies that any algorithm with incoherent access requires at least $\Omega(\sqrt{d})$ queries to complete DSEC. Notably, this scaling matches that of unitarity estimation [41], consistent with the fact that unitarity estimation is a special case of DSEC.
Proof sketch of Theorem 1.
Inspired by [41], we introduce an intermediate distribution $p_{\mathcal{D}}(\ell)$, corresponding to the probability of obtaining measurement result $\ell$ when both channels are replaced by the completely depolarizing channel $\mathcal{D}(A)\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}[A]I/d$. By the triangle inequality, the target TVD is upper bounded by two terms: $\left\lVert p_{1}-p_{\mathcal{D}}\right\rVert_{\rm TV}$ and $\left\lVert p_{\mathcal{D}}-p_{2}\right\rVert_{\rm TV}$. We prove that each term is bounded by $\mathcal{O}(T^{2}/d)$, where $T$ is the number of queries per device. Therefore, completing Problem 1 with incoherent access requires $T=\Omega(\sqrt{d})$ queries. The detailed proof is provided in Appendix B.1. ∎
Coherent Access.
We now turn to coherent access, where the learner may interleave multiple queries with arbitrary quantum channels before measurement. In various quantum learning tasks, coherent access often provides substantial advantages over incoherent access [44, 41, 37, 45]. A central question is whether such advantages exist for DSEC. Remarkably, we prove that even with this significantly more powerful learning access, completing Problem 1 still requires $\Omega(\sqrt{d})$ queries. This result holds even with multi-round LOCC, arbitrarily large ancillary systems, and adaptive SPAM settings, highlighting the fundamental hardness of DSEC.
Theorem 2.
Any algorithm with coherent access, possibly utilizing multi-round LOCC, ancillary systems, and adaptive strategies, requires $\Omega(\sqrt{d})$ queries to complete Problem 1.
Likewise, this theorem implies that any algorithm with coherent access requires at least $\Omega(\sqrt{d})$ queries to complete DSEC. Importantly, we will show that coherent access offers no asymptotic advantage for completing DSEC: our incoherent algorithm matches this lower bound.
Proof sketch of Theorem 2.
Suppose we query $T$ times on each device. Analyzing coherent access is generally subtle due to the arbitrary quantum channels that may appear between queries. To overcome this difficulty, inspired by [50], we reformulate the problem in terms of Choi operators. Specifically, we work with the Choi operator of $U^{\otimes T}$, denoted $J_{U}^{(T)}$. This formulation absorbs all interleaving channels into an enlarged effective input system. The two probabilities $p_{1}$ and $p_{2}$ can then be expressed via the following averaged Choi operators:
$\bar{J}_{1}^{(T)}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U}\;J_{U}^{(T)}\otimes J_{U}^{(T)},\quad\bar{J}_{2}^{(T)}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U,V}\;J_{U}^{(T)}\otimes J_{V}^{(T)}.$ (3)
Subsequently, as in the proof of lower bound for incoherent access, we introduce an intermediate probability distribution $p_{a}(\ell)$, defined as the probability of obtaining measurement result $\ell$ when both channels are replaced by the approximate $T$-design channel with Choi operator $J_{a}^{(T)}$ (see definition in [50], also Appendix Eq. (29)). We again analyze the TVDs: $\left\lVert p_{1}-p_{a}\right\rVert_{\rm TV}$ and $\left\lVert p_{a}-p_{2}\right\rVert_{\rm TV}$ separately. For the first term, we directly bound it using the relative error of the approximate $T$-design channel, and the corresponding result is $\mathcal{O}(T^{2}/d)$. For the second term, we relate it to the distance between the approximate $T$-design channel and a “measure-and-prepare” map. Using tools from quantum cloning theory [51], we show that this term is also bounded by $\mathcal{O}(T^{2}/d)$. Therefore, completing DSEC with coherent access requires $T=\Omega(\sqrt{d})$ queries. The detailed proof is provided in Appendix B.2. ∎
Lower Bound for DSEC.
Combining the above results with an additional analysis of the $\varepsilon$-dependence, we establish the complete lower bound for completing DSEC. The result, proven in Appendix B.3, is as follows.
Theorem 3.
Any algorithm, possibly utilizing multi-round LOCC, ancillary systems, and adaptive strategies, requires $\Omega(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$ queries to complete DSEC.
This theorem provides an information-theoretic lower bound that holds regardless of the learning access model.
Upper bound.— Having established the lower bound, we now present algorithms that achieve the optimal query complexity. We provide two algorithms: one for coherent access (applicable to unitary channels) and one for incoherent access (applicable to general channels).
Coherent Access.
We first consider the case where both unknown channels are unitary, i.e., $\mathcal{E}(\cdot)=U(\cdot)U^{\dagger}$ and $\mathcal{F}(\cdot)=V(\cdot)V^{\dagger}$. For this setting, we propose an algorithm based on symmetric collective measurements [52, 37]. Given a state $(s+2)$-design $\{|{\phi_{j}}\rangle\}_{j=1}^{L}$, the symmetric collective measurement $\mathcal{M}_{s}$ on $\mathcal{H}^{\otimes s}$ is defined as the following POVM:
$\mathcal{M}_{s}\mathrel{\mathop{\ordinarycolon}}=\left\{\frac{\kappa_{s}}{L}|{\phi_{j}}\rangle\!\langle{\phi_{j}}|^{\otimes T}\right\}_{j=1}^{L}\cup\left\{I-\Pi_{\rm sym}^{(d,s)}\right\},$ (4)
where $\kappa_{s}\mathrel{\mathop{\ordinarycolon}}=\binom{d+s-1}{s}$ and $\Pi_{\rm sym}^{(d,s)}$ is the projector onto the symmetric subspace of $\mathcal{H}^{\otimes s}$. Our algorithm proceeds as follows:
  1. 1.
    Randomly generate pure state $|{\psi}\rangle^{\otimes T}$ on each quantum device, where $|{\psi}\rangle$ from a state $4$-design ensemble.
  2. 2.
    Apply the unitary operators $U^{\otimes T}$ and $V^{\otimes T}$ on $|{\psi}\rangle^{\otimes T}$.
  3. 3.
    Measure $(U|{\psi}\rangle)^{\otimes T}$ and $(V|{\psi}\rangle)^{\otimes T}$ with the POVM $\mathcal{M}_{T}$. Record the measurement results on two devices as $|{\phi_{A}}\rangle$ and $|{\phi_{B}}\rangle$, and compute $\tilde{f}=|\langle\phi_{A}|\phi_{B}\rangle|^{2}$.
  4. 4.
    Return the unbiased estimator:
    $\tilde{\chi}\mathrel{\mathop{\ordinarycolon}}=\frac{(d+1)(d+T)^{2}}{T^{2}d}\tilde{f}-\frac{(d+1)(d+2T)+T^{2}}{T^{2}d}.$ (5)
This algorithm requires shared randomness only for state preparation and is described in pseudocode in Algorithm 1. We summarize the query complexity of this algorithm in the following theorem, which is proven in Appendix C.1.
Theorem 4.
For unitary channels $\mathcal{E}$ and $\mathcal{F}$, the expectation and variance of estimator $\tilde{\chi}$ defined in Eq. (5) are given by
$\mathbb{E}\tilde{\chi}=\frac{1}{d^{2}}\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}],\quad\mathbb{V}(\tilde{\chi})\leq\mathcal{O}\left(\frac{d}{m^{2}}+\frac{1}{m}+\frac{1}{d}\right).$ (6)
Consequently, the query complexity is $\mathcal{O}(\max\{1/\varepsilon^{2},\sqrt{d}/\varepsilon\})$.
Combined with Theorem 3, this coherent algorithm is optimal in both $d$ and $\varepsilon$ for unitary channels. However, this algorithm fundamentally requires unitary channels since the proof relies on pure output states (see Appendix C.1 for details). While random purification channels [53, 54, 34] could potentially extend this result to general channels, we pursue a different approach: we show that incoherent access already suffices to achieve optimality.
Incoherent Access.
We now present our main algorithm, which works for general channels using only incoherent access. Let $\{|{a}\rangle\!\langle{a}|\}_{a}$ be the computational basis. The algorithm repeats the following procedure $T$ times independently:
  1. 1.
    Randomly generate a pure state $|{\psi}\rangle$ from a state $4$-design ensemble on each quantum device.
  2. 2.
    Apply the unknown channels $\mathcal{E}$ and $\mathcal{F}$ on each $|{\psi}\rangle$.
  3. 3.
    Measure two output states $\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)$ and $\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)$ in the basis $\{Q^{\dagger}|{a}\rangle\!\langle{a}|Q\}_{a}$ $m$ times, where $Q$ is randomly sampled from a unitary $4$-design ensemble.
  4. 4.
    Record the measurement results as $\{a_{i}\}_{i=1}^{m}$ and $\{b_{i}\}_{i=1}^{m}$, respectively. Compute
    $\tilde{g}=\frac{1}{m^{2}}\sum_{i,j=1}^{m}\mathbb{1}(a_{i},b_{j}),$ (7)
    where $\mathbb{1}(a,b)=1$ if $a=b$; otherwise 0.
  5. 5.
    Return the unbiased estimator:
    $\tilde{\omega}=\frac{(d+1)^{2}}{d}\tilde{g}-\frac{d+2}{d}.$ (8)
This incoherent algorithm is summarized in Algorithm 2. Additionally, the corresponding query complexity is shown in the following theorem, which is provided in Appendix C.2.
Theorem 5.
For general channels $\mathcal{E}$ and $\mathcal{F}$, the expectation and variance of estimator $\tilde{\omega}$ defined in Eq. (8) are given by
$\mathbb{E}\tilde{\omega}=\frac{1}{d^{2}}\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}],\quad\mathbb{V}(\tilde{\omega})\leq\mathcal{O}\left(\frac{d}{m^{2}}+\frac{1}{m}+\frac{1}{d}\right).$ (9)
Consequently, the query complexity is $\mathcal{O}(\max\{1/\varepsilon^{2},\sqrt{d}/\varepsilon\})$.
Combined with Theorem 3, the query complexity of DSEC is exactly $\Theta(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$, as this incoherent algorithm matches the lower bound and applies to general channels.
This incoherent algorithm enjoys several notable advantages. First, it is ancilla-free and non-adaptive, making it immediately implementable on near-term devices. Second, it can also serve as a tool for unitarity estimation [41], since it works without the unitary assumption. Third, despite incoherent access being strictly weaker than coherent access in general, both algorithms achieve the same query complexity for DSEC. This surprising phenomenon parallels the result for DIPE [15], which shows that symmetric collective measurements do not reduce complexity in distributed learning.
The key to these advantages lies in the different types of shared randomness employed. The coherent algorithm shares only the randomness for state preparation, whereas the incoherent algorithm additionally shares measurement settings. This shared measurement randomness effectively boosts the performance of the incoherent algorithm, enabling it to achieve the optimal query complexity for general channels. This finding highlights that shared randomness, rather than coherent access or collective measurements, is the key resource for efficient distributed learning. On the other hand, the coherent algorithm has its own advantage: less shared randomness translates to less communication overhead. Thus, there exists a trade-off between classical communication resources and the power of the learning algorithm.
Classical Shadow.— To quantify the advantage of shared randomness, we compare our algorithms with independent classical shadow for channels [35, 36, 37], which use no shared randomness between devices. We adopt the approach of [37] as our baseline, since it achieves near-optimal performance for unitary channels and naturally extends to incoherent access for general channels.
The independent classical shadow approach for DSEC proceeds as follows (see Algorithm 3 for details). Suppose $\mathcal{E}(\cdot)=U(\cdot)U^{\dagger}$ and $\mathcal{F}(\cdot)=V(\cdot)V^{\dagger}$ are unitary channels. On each device independently, we randomly select a state $|{\psi}\rangle$ from a state $4$-design ensemble, apply $U^{\otimes s}$ (or $V^{\otimes s}$) to $|{\psi}\rangle^{\otimes s}$, and perform the symmetric collective measurement $\mathcal{M}_{s}$. Repeating this procedure $T$ times on each device yields two independent sets of classical snapshots, $\{X_{t}\}$ and $\{Y_{t}\}$. The inner product is then estimated via
$\tilde{\gamma}=\frac{1}{T^{2}d^{2}}\sum_{i,j=1}^{T}\operatorname{Tr}[X_{i}Y_{j}].$ (10)
The key difference from our algorithms is that no shared randomness is used; SPAM settings are chosen independently on each device. The following proposition, proven in Appendix D, establishes the query complexity.
Proposition 6.
For unitary channels $\mathcal{E}$ and $\mathcal{F}$, the expectation and variance of estimator $\tilde{\gamma}$ defined in Eq. (10) are given by
$\mathbb{E}\tilde{\gamma}=\frac{1}{d^{2}}\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}],\quad\mathbb{V}(\tilde{\gamma})=\mathcal{O}\left(\frac{d^{2}}{T^{2}s^{2}}+\frac{1}{T}\right).$ (11)
Consequently, the query complexity is $\mathcal{O}(\max\{d/\varepsilon,1/\varepsilon^{2}\})$. For $s=1$, the result extends to general channels.
This result reveals that independent classical shadow requires $\mathcal{O}(\max\{d/\varepsilon,1/\varepsilon^{2}\})$ queries for DSEC, regardless of the number of queries per snapshot. Notably, the lower bound in Theorem 2 also applies to independent classical shadow based algorithms, as they fall within the same learning access. In contrast, our algorithms with shared randomness achieve the optimal query complexity of $\mathcal{O}(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$, demonstrating a quadratic improvement in the dimension-dependent term. This square-root advantage directly arises from the shared randomness in SPAM settings, which enables the two devices to coordinate their measurements.
Access Model Lower Bound Algorithm Query Complexity Applicable Channel
Coherent $\Omega(\sqrt{d})$ (Theorem 3) Independent classical shadow $\mathcal{O}(d)$ (Proposition 6) Unitary
RM with shared randomness $\mathcal{O}(\sqrt{d})$ (Theorem 4) Unitary
Incoherent Independent classical shadow $\mathcal{O}(d)$ (Proposition 6) General
RM with shared randomness $\mathcal{O}(\sqrt{d})$ (Theorem 5) General
Table 1: Summary of our query complexity results for distributed similarity estimation of quantum channels (DSEC), where $d$ is the channel dimension and RM is short for randomized measurements.
Conclusion.— In this work, we characterized the optimal query complexity of distributed similarity estimation of quantum channels (DSEC). We proved that $\Theta(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$ queries are both necessary and sufficient for this task. The lower bound holds even in the strongest setting (multi-round LOCC, coherent access, and adaptive strategies), while the upper bound is achieved in the weakest setting (incoherent access, ancilla-free, and non-adaptive). For the upper bound, we proposed two randomized measurement-based algorithms that effectively utilize shared randomness between devices. Both achieve optimal query complexity, with the incoherent algorithm applicable to general channels. We also compared our algorithms with independent classical shadow, demonstrating a square-root advantage in query complexity. This advantage arises from the shared randomness employed in our algorithms, which allows the two devices to coordinate SPAM settings. The obtained results are summarized in Table 1.
Several directions remain open for future research. First, existing cross-platform verification protocols typically assume identical operations on both platforms, which may fail in practice; developing robust verification algorithms that relax this assumption is an important goal. Second, since classical shadow methods utilize information about sampled unitaries, investigating whether this information can be better exploited to reduce query complexity is an intriguing question.
Note added.
While completing this manuscript, we became aware of a related work by Ananth et al. [55], which also studied the lower bound for completing Problem 1 with coherent access. While both works identify the $\Omega(\sqrt{d})$ barrier, the proof techniques are distinct and complementary. Our approach relies on Le Cam’s two-point method and connections to optimal quantum cloning [51], whereas Ananth et al. utilize linear programming. This diversity of techniques enriches our understanding of the fundamental hardness of this problem. In addition, we establish the $\varepsilon$-dependent lower bound.
Acknowledgments.
This work was supported by the National Natural Science Foundation of China (Grant Nos. 62471126 and 12504584), the Jiangsu Frontier Technology Research and Development Plan (Grant No. BF2025066), the Fundamental Research Funds for the Central Universities (Grant No. 2242022k60001), the SEU Innovation Capability Enhancement Plan for Doctoral Students (Grant No. CXJH_SEU 24083), and the National Key R&D Program of China (Grant No. 2022YFF0712800).
Appendix A Preliminaries
In this section, we first set some notation used in this work. Then, we introduce two central tools utilized in this work: the Haar measure and approximate unitary designs. Both play fundamental roles in quantum information theory [56, 50], and they are essential for establishing the query complexities of our algorithms as well as the lower bounds for completing DSEC.
A.1 Notation
Let $\mathcal{H}$ be the Hilbert space of a $d$-dimensional system. The set of Hermitian operators on $\mathcal{H}$ is denoted by $\mathcal{B}(\mathcal{H})$, and the set of density matrices on $\mathcal{H}$ is denoted by $\mathcal{D}(\mathcal{H})$. For a quantum channel $\mathcal{E}$, we define its Choi operator as [39]
$J_{\mathcal{E}}\mathrel{\mathop{\ordinarycolon}}=(\mathcal{E}\otimes\mathcal{I})(|{\Phi}\rangle\!\langle{\Phi}|),$ (12)
where $|{\Phi}\rangle\mathrel{\mathop{\ordinarycolon}}=\sum_{i}|{ii}\rangle$ is the unnormalized maximally entangled state. The Choi operator provides the useful identity
$\operatorname{Tr}\left[O\mathcal{E}(\rho)\right]=\operatorname{Tr}\left[\left(O\otimes\rho^{T}\right)J_{\mathcal{E}}\right].$ (13)
Additionally, we can also use a set of Kraus operators to describe a quantum channel. A quantum channel $\mathcal{E}$ can be described by a set of Kraus operators $\{E_{i}\}_{i}$ [57], and we have
$\mathcal{E}(\rho)=\sum_{i}E_{i}\rho E_{i}^{\dagger}.$ (14)
For two quantum channels $\mathcal{E}$ and $\mathcal{F}$, with Kraus operators $\{E_{i}\}_{i}$ and $\{F_{j}\}_{j}$, the inner product of their Choi operators can be represented as
$\operatorname{Tr}\left[J_{\mathcal{E}}J_{\mathcal{F}}\right]=\sum_{i,j}\operatorname{Tr}^{2}\left[E_{i}^{\dagger}F_{j}\right].$ (15)
For any operator $A\in\mathcal{B}(\mathcal{H})$, we use the notation $|A\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ to represent the corresponding vectorized operator. For example,
$||{\psi}\rangle\!\langle{\phi}|\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\mathrel{\mathop{\ordinarycolon}}=|{\psi}\rangle\otimes|{\phi^{*}}\rangle,\quad|ABC^{\dagger}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=A\otimes C^{*}|B\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}.$ (16)
The inner product between two vectorized operators is defined as $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}A|B\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}=\operatorname{Tr}[A^{\dagger}B]$. For operators $A$ and $B$, we write $A\succeq B$ if ABA-B is positive semidefinite.
A.2 Haar Measure and Permutation Operators
Let $\mu_{H}$ be the Haar measure over the unitary group. The Haar random channel is defined as
$\mathcal{E}_{H}^{(k)}(A)\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mu_{H}}U^{\otimes k}AU^{\dagger\otimes k},\quad A\in\mathcal{B}(\mathcal{H}^{\otimes k}).$ (17)
To describe the Haar random channel, we introduce permutation operators. Let $\mathcal{S}_{k}$ be the symmetric group on $k$ elements. For $\pi\in\mathcal{S}_{k}$, the corresponding permutation operator $P_{\pi}$ acts on product states as
$P_{\pi}|{\psi_{1}}\rangle\otimes\cdots\otimes|{\psi_{k}}\rangle=|{\psi_{\pi^{-1}(1)}}\rangle\otimes\cdots\otimes|{\psi_{\pi^{-1}(k)}}\rangle,\quad\forall\;\psi_{1},\cdots\psi_{k}\in\mathcal{D}(\mathcal{H}).$ (18)
With the above definition, we have the following lemma to describe the Haar random channel.
Lemma 7 (Weingarten Calculus [56, 50]).
Let $\mathcal{H}$ be a Hilbert space with dimension $d$, $\mathcal{S}_{k}$ be the symmetric group, and $A\in\mathcal{B}(\mathcal{H}^{\otimes k})$. Then, we have
$\mathcal{E}_{H}^{(k)}(A)=\sum_{\sigma,\tau\in\mathcal{S}_{k}}{\rm Wg}_{\sigma,\tau}(d)\operatorname{Tr}[AP^{\dagger}_{\sigma}]P_{\tau},\quad A\in\mathcal{B}(\mathcal{H}^{\otimes k}).$ (19)
Here, ${\rm Wg}_{\sigma,\tau}(d)$ are the elements of the $k!\times k!$ Weingarten matrix, which is defined as the inverse of the Gram matrix, $G_{\sigma,\tau}(d)=\operatorname{Tr}(P_{\sigma}P_{\tau}^{\dagger})$, given by the inner products of the permutation operators. Consequently, the Choi operator of Haar random channel can be represented as
$J_{H}^{(k)}=\sum_{\sigma,\tau\in\mathcal{S}_{k}}{\rm Wg}_{\sigma,\tau}(d)P_{\tau}\otimes P_{\sigma}.$ (20)
For the special cases $k=1$ and $k=2$, the Haar twirl admits simple closed forms.
Lemma 8 (Special Cases of Weingarten Calculus, $k=1$ and $k=2$ cases [56]).
Let $\mathcal{H}$ be a Hilbert space with dimension $d$, $A\in\mathcal{B}(\mathcal{H})$, and $B\in\mathcal{B}(\mathcal{H}^{\otimes 2})$, we have
$\mathcal{E}_{H}^{(1)}(A)=\frac{\operatorname{Tr}[A]}{d}I,\quad\mathcal{E}_{H}^{(2)}(B)=\frac{d\operatorname{Tr}[B]-\operatorname{Tr}[\mathbb{F}B]}{d(d^{2}-1)}I\otimes I+\frac{d\operatorname{Tr}[\mathbb{F}B]-\operatorname{Tr}[B]}{d(d^{2}-1)}\mathbb{F},$ (21)
where $\mathbb{F}$ is the SWAP operator.
For quantum states in $\mathcal{H}$, we define the Haar measure on quantum states as [56]
$\mathbb{E}_{|{\psi}\rangle\sim\mu_{H}}|{\psi}\rangle\!\langle{\psi}|^{\otimes k}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mu_{H}}U^{\otimes k}|{\phi}\rangle\!\langle{\phi}|^{\otimes k}U^{\dagger\otimes k}=\frac{\Pi_{\rm sym}^{(d,k)}}{\kappa_{k}},\quad|{\phi}\rangle\!\langle{\phi}|\in\mathcal{D}(\mathcal{H}),$ (22)
where $\kappa_{k}\mathrel{\mathop{\ordinarycolon}}=\binom{d+k-1}{k}$ and $\Pi_{\rm sym}^{(d,k)}$ is the orthogonal projector onto the symmetric subspace of $\mathcal{H}^{\otimes k}$, defined as
$\Pi_{\rm sym}^{(d,k)}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{k!}\sum_{\pi\in\mathcal{S}_{k}}P_{\pi}.$ (23)
Additionally, a pure state ensemble $\mathcal{A}$ is said to form a state $k$-design if
$\mathbb{E}_{|{\psi}\rangle\sim\mathcal{A}}|{\psi}\rangle\!\langle{\psi}|^{\otimes k}=\mathbb{E}_{|{\psi}\rangle\sim\mu_{H}}|{\psi}\rangle\!\langle{\psi}|^{\otimes k}=\frac{\Pi_{\rm sym}^{(d,k)}}{\kappa_{k}}.$ (24)
Similarly, a unitary ensemble $\mathcal{A}$ is said to form a unitary $k$-design if and only if
$\mathcal{E}_{\mathcal{A}}^{(k)}(A)\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mathcal{A}}U^{\otimes k}AU^{\dagger\otimes k}=\mathcal{E}_{H}^{(k)}(A),\quad\forall\;A\in\mathcal{B}(\mathcal{H}^{\otimes k}).$ (25)
Based on the properties of state designs, we have the following results for the symmetric collective measurement $\mathcal{M}_{s}$.
Lemma 9 (Lemmas 13 and 14 in [52]).
Let $\mathcal{H}$ be a Hilbert space with dimension $d$. For the measurement $\mathcal{M}_{s}$ on the pure state $|{\psi}\rangle\!\langle{\psi}|^{\otimes s}$, the expectation of the measurement result is
$\mathbb{E}|{\phi}\rangle\!\langle{\phi}|=\frac{I+s|{\psi}\rangle\!\langle{\psi}|}{d+s}.$ (26)
Additionally, we have
$\mathbb{E}|{\phi}\rangle\!\langle{\phi}|^{\otimes 2}=\frac{2}{(d+s)(d+s+1)}\left[(I+s|{\psi}\rangle\!\langle{\psi}|)^{\otimes 2}-\frac{s(s+1)}{2}|{\psi}\rangle\!\langle{\psi}|^{\otimes 2}\right]\Pi_{\rm sym}^{(d,2)}.$ (27)
A.3 Approximate Unitary Designs
A unitary ensemble $\mathcal{A}$ is an $\varepsilon$-approximate unitary $k$-design if [50]
$(1-\varepsilon)\mathcal{E}_{\mathcal{A}}^{(k)}\preceq\mathcal{E}_{H}^{(k)}\preceq(1+\varepsilon)\mathcal{E}_{\mathcal{A}}^{(k)},$ (28)
where $\varepsilon$ is called relative error. Here $\mathcal{E}_{1}\preceq\mathcal{E}_{2}$ denotes that $\mathcal{E}_{2}-\mathcal{E}_{1}$ is a completely-positive map. A standard approximate $k$-design channel [50] and the corresponding Choi operator are defined as
$\mathcal{E}_{a}^{(k)}(A)\mathrel{\mathop{\ordinarycolon}}=\frac{1}{d^{k}}\sum_{\sigma\in\mathcal{S}_{k}}\operatorname{Tr}[AP_{\sigma}^{\dagger}]P_{\sigma},\quad J_{a}^{(k)}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{d^{k}}\sum_{\pi\in\mathcal{S}_{k}}P_{\pi}\otimes P_{\pi}.$ (29)
For $k^{2}\leq d$, the corresponding relative error satisfies
$\varepsilon=\frac{k^{2}/2d}{1-k^{2}/2d}.$ (30)
We also have the following inequality for $J_{a}^{(k)}$ and $J_{H}^{(k)}$.
Lemma 10 (Lemma 2 of [41]).
Suppose $\frac{d^{k}}{d(d+1)\cdots(d+T-1)}>\frac{1}{2}$, then, we have
$J^{(k)}_{H}\succeq\frac{1}{d(d+1)\cdots(d+k-1)}\sum_{\sigma\in\mathcal{S}_{k}}P_{\sigma}\otimes P_{\sigma}=\frac{d^{k}}{d(d+1)\cdots(d+k-1)}J_{a}^{(k)}.$ (31)
Appendix B Lower Bound
In this section, we establish lower bounds for completing DSEC with both incoherent and coherent accesses. Here, we consider two unitary channels $\mathcal{U}=U(\cdot)U^{\dagger}$ and $\mathcal{V}=V(\cdot)V^{\dagger}$ acting on $d$-dimensional subsystems of a Hilbert space $\mathcal{H}\simeq\mathcal{H}_{\rm main}\otimes\mathcal{H}_{\rm aux}$, where $\mathcal{H}_{\rm main}$ is the main system with dimension $d$ and $\mathcal{H}_{\rm aux}$ is an auxiliary system with dimension $d^{\prime}$. Additionally, as shown in [44, Remark 4.19], we only need to consider rank-$1$ POVMs on two quantum systems.
Our approach is to reduce the estimation task to a two-hypothesis distinguishing problem. In particular, to derive a lower bound for completing DSEC, we consider the following distinguishing problem, which is a restatement of Problem 1 in the main text:
Problem 1 (Distinguishing Problem).
We want to distinguish the following two cases:
  1. 1.
    Two quantum devices perform the same unitary $U$, which is a Haar random unitary;
  2. 2.
    Two quantum devices independently perform two unitaries $U$ and $V$, which are two independent Haar random unitaries.
Define the Choi operators of $U$ and $V$ as $J_{U}$ and $J_{V}$, respectively. If an estimation algorithm can accurately estimate $\operatorname{Tr}[J_{U}J_{V}]/d^{2}$, then it can be leveraged to solve the distinguishing problem above. Consequently, any lower bound for this distinguishing problem immediately yields a lower bound for DSEC. To bound the success probability of solving the distinguishing problem, we invoke Le Cam’s two-point method as follows.
Lemma 11 (Le Cam’s Two-Point Method [46]).
Let $p^{U,U}(\ell)$ and $p^{U,V}(\ell)$ be the probabilities of obtaining measurement outcome $\ell$ under Cases 1 and 2 of Problem 1, respectively. Then, the probability that the learning algorithm correctly solves the distinguishing problem in Problem 1 is upper bounded by total variation distance (TVD), defined as
$\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{2}\sum_{\ell}\left|\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)\right|.$ (32)
In the remainder of this section, we analyze the TVD between $p^{U,U}(\ell)$ and $p^{U,V}(\ell)$, respectively, thereby establishing the corresponding lower bounds.
B.1 Incoherent Access
Here, we prove the lower bound for completing DSEC with incoherent access. First, we introduce the tree representation of channel learning with incoherent access, which is a powerful tool in proving lower bounds for learning tasks [44, 41, 47, 48, 49].
B.1.1 Proof of Theorem 1
Definition 12 (Tree Representation for Learning Quantum Channels [44]).
Consider two channels $\mathcal{U}$ and $\mathcal{V}$ acting on $\mathcal{H}_{\rm main}$. A learning algorithm with incoherent access can be represented as a rooted tree $\mathcal{T}$ of depth $T$ such that each node encodes all measurement outcomes the algorithm has received thus far. The tree has the following properties:
  • Each node $u$ has an associated probability $p^{U,V}(u)$.
  • The root of the tree $r$ has an associated probability $p^{U,V}(r)=1$.
  • At each non-leaf node $u$, we prepare a state $|{\phi_{u,1}}\rangle\otimes|{\phi_{u,2}}\rangle$ on $\mathcal{H}^{\otimes 2}$, apply channels $\mathcal{U}$ and $\mathcal{V}$ onto two $d$-dimensional subsystems, and measure a rank-$1$ POVM $\{w_{v}^{u}d^{2}d^{\prime 2}|{\psi^{u}_{v,1}}\rangle\!\langle{\psi^{u}_{v,1}}|\otimes|{\psi^{u}_{v,2}}\rangle\!\langle{\psi^{u}_{v,2}}|\}_{v}$ (which can depend on $u$) on the entire system to obtain a classical outcome $v$. Each child node $v$ of the node $u$ corresponds to a particular POVM outcome $v$ and is connected by the edge $e_{u,v}$. We denote the set of child nodes of node $u$ as ${\rm child}(u)$.
  • If $v$ is a child node of $u$, then
    $p^{U,V}(v)=p^{U,V}(u)w_{v}^{u}d^{2}d^{\prime 2}\operatorname{Tr}\left[\left(\bigotimes_{i=1}^{2}|{\psi^{u}_{v,i}}\rangle\!\langle{\psi^{u}_{v,i}}|\right)(\mathcal{U}\otimes\mathcal{I}_{\rm aux}\otimes\mathcal{V}\otimes\mathcal{I}_{\rm aux})\left(\bigotimes_{i=1}^{2}|{\phi_{u,i}}\rangle\!\langle{\phi_{u,i}}|\right)\right].$ (33)
  • Each root-to-leaf path is of length $T$. For a leaf corresponding to node $\ell$, $p^{U,V}(\ell)$ is the probability that the classical memory is in state $\ell$ after the learning procedure.
Based on the above tree representation and distinguishing problem, we are ready to prove the lower bound for completing DSEC with incoherent access, i.e., Theorem 1. As shown in Lemma 11, our goal is to obtain an upper bound for $\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}$. By the triangle inequality, we have the following upper bound for each leaf $\ell$,
$\left|\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)\right|$ $\leq\left|\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)-p^{\mathcal{D}}(\ell)\right|+\left|p^{\mathcal{D}}(\ell)-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)\right|,$ (34)
$\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}$ $\leq\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-p^{\mathcal{D}}\right\rVert_{{\rm TV}}+\left\lVert p^{\mathcal{D}}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}},$ (35)
where $\mathcal{D}(A)\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}[A]I/d$ is the completely depolarizing channel and $p^{\mathcal{D}}(\ell)$ is the probability that both two quantum channels are $\mathcal{D}$. Using Lemma 13, we can find upper bounds for the two terms in Eq. (35) and obtain
$\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}$ $\leq\frac{6T^{2}}{d},$ (36)
which implies that we require $T=\Omega(\sqrt{d})$. Therefore, we complete the proof of Theorem 1.
B.1.2 Technical Lemmas
In the following, we prove the following key lemma.
Lemma 13.
For the two terms defined in Eq. (35), we have the following upper bounds,
$\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-p^{\mathcal{D}}\right\rVert_{{\rm TV}}$ $\leq\frac{4T^{2}}{d},\quad\left\lVert p^{\mathcal{D}}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}\leq\frac{2T^{2}}{d}.$ (37)
Proof of Lemma 13.
Before the proof, we introduce some useful notation. Each root-to-leaf path in the learning tree $\mathcal{T}$ is uniquely specified by a sequence of vertices $v_{0},v_{1},\cdots,v_{T}$, where $v_{0}=r$ is the root and $v_{T}=\ell$ is a leaf. Since $\ell$ uniquely determines the shortest path back to the root, specifying the leaf $\ell$ is equivalent to specifying the entire path $v_{0}=r,v_{1},\cdots,v_{T-1},v_{T}=\ell$. Therefore, the probabilities of reaching leaf $\ell$ can be expressed as
$p^{\mathcal{D}}(\ell)$ $=\prod_{t=1}^{T}w_{t}d^{2}d^{\prime 2}\langle{\psi_{t,1}}|\left(\mathcal{D}\otimes\mathcal{I}_{\rm aux}\right)(|{\phi_{t,1}}\rangle\!\langle{\phi_{t,1}}|)|{\psi_{t,1}}\rangle\langle{\psi_{t,2}}|\left(\mathcal{D}\otimes\mathcal{I}_{\rm aux}\right)(|{\phi_{t,2}}\rangle\!\langle{\phi_{t,2}}|)|{\psi_{t,2}}\rangle,$ (38)
$p^{U,V}(\ell)$ $=\prod_{t=1}^{T}w_{t}d^{2}d^{\prime 2}\langle{\psi_{t,1}}|\left(\mathcal{U}\otimes\mathcal{I}_{\rm aux}\right)(|{\phi_{t,1}}\rangle\!\langle{\phi_{t,1}}|)|{\psi_{t,1}}\rangle\langle{\psi_{t,2}}|\left(\mathcal{V}\otimes\mathcal{I}_{\rm aux}\right)(|{\phi_{t,2}}\rangle\!\langle{\phi_{t,2}}|)|{\psi_{t,2}}\rangle,$ (39)
similar to $p^{U,U}$. As shown in [41], we can decompose the $|{\phi_{t,i}}\rangle$ and $|{\psi_{t,i}}\rangle$ as
$|{\phi_{t,i}}\rangle=\sum_{j=0}^{d^{\prime}-1}|{\phi_{t,i,j}}\rangle\otimes|{j}\rangle,\quad|{\psi_{t,i}}\rangle=\sum_{j=0}^{d^{\prime}-1}|{\psi_{t,i,j}}\rangle\otimes|{j}\rangle,\quad i=1,2,$ (40)
where the first tensor factor corresponds to the main system $\mathcal{H}_{\rm main}$ and the second to the auxiliary system $\mathcal{H}_{\rm aux}$. Note that the vectors $|{\phi_{t,i,j}}\rangle$ and $|{\psi_{t,i,j}}\rangle$ are not required to be normalized. For convenience, let $\bm{j}\mathrel{\mathop{\ordinarycolon}}=(\bm{j}_{1},\cdots,\bm{j}_{T})$ range over all sequences of length $T$, with each $j_{t}\in\{0,\ldots,d^{\prime}-1\}$. Define $W_{\ell}\mathrel{\mathop{\ordinarycolon}}=\prod w_{t}$, and
$|{\Phi_{\ell,i,\bm{j}}}\rangle\mathrel{\mathop{\ordinarycolon}}=\bigotimes_{t=1}^{T}|{\phi_{t,i,\bm{j}_{t}}}\rangle,\quad|{\Psi_{\ell,i,\bm{j}}}\rangle\mathrel{\mathop{\ordinarycolon}}=\bigotimes_{t=1}^{T}|{\psi_{t,i,\bm{j}_{t}}}\rangle,\quad i=1,2,$ (41)
we have
$\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)$ $=\mathbb{E}_{U,V\sim\mu_{H}}\sum_{\bm{j},\bm{k},\bm{x},\bm{y}}\prod_{t=1}^{T}w_{t}d^{2}d^{\prime 2}\langle{\psi_{t,1,\bm{j}_{t}}}|\mathcal{U}(|{\phi_{t,1,\bm{j}_{t}}}\rangle\!\langle{\phi_{t,1,\bm{k}_{t}}}|)|{\psi_{t,1,\bm{k}_{t}}}\rangle\langle{\psi_{t,2,\bm{x}_{t}}}|\mathcal{V}(|{\phi_{t,2,\bm{x}_{t}}}\rangle\!\langle{\phi_{t,2,\bm{y}_{t}}}|)|{\psi_{t,2,\bm{y}_{t}}}\rangle$ (42)
$=(dd^{\prime})^{2T}W_{\ell}\sum_{\bm{j},\bm{k},\bm{x},\bm{y}}\mathbb{E}_{U,V\sim\mu_{H}}\langle{\Psi_{t,1,\bm{j}}}|U^{\otimes T}|{\Phi_{t,1,\bm{j}}}\rangle\!\langle{\Phi_{t,1,\bm{k}}}|U^{\dagger\otimes T}|{\Psi_{t,1,\bm{k}}}\rangle\langle{\Psi_{t,2,\bm{x}}}|V^{\otimes T}|{\Phi_{t,2,\bm{x}}}\rangle\!\langle{\Phi_{t,2,\bm{y}}}|V^{\dagger\otimes T}|{\Psi_{t,2,\bm{y}}}\rangle$
$=(dd^{\prime})^{2T}W_{\ell}\prod_{i=1}^{2}\left(\sum_{\bm{j},\bm{k}}\langle{\Psi_{\ell,i,\bm{j}}\otimes\Phi_{\ell,i,\bm{j}}^{*}}|J_{H}^{(T)}|{\Psi_{\ell,i,\bm{k}}\otimes\Phi_{\ell,i,\bm{k}}^{*}}\rangle\right),$ (43)
where $J_{H}^{(T)}$ is the Choi operator of Haar random channel defined in Eq. (20). Similarly, we have
$\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)$ $=(dd^{\prime})^{2T}W_{\ell}\sum_{\bm{j},\bm{k}}\langle{\Psi_{\ell,\bm{j}}\otimes\Phi_{\ell,\bm{j}}^{*}}|J_{H}^{(2T)}|{\Psi_{\ell,\bm{k}}\otimes\Phi_{\ell,\bm{k}}^{*}}\rangle,$ (44)
$p^{\mathcal{D}}(\ell)$ $=d^{\prime 2T}W_{\ell}\sum_{\bm{j},\bm{k}}\langle\Psi_{\ell,\bm{j}}|\Psi_{\ell,\bm{k}}\rangle\langle\Phi_{\ell,\bm{k}}|\Phi_{\ell,\bm{j}}\rangle,$ (45)
where
$|{\Phi_{\ell,\bm{j}}}\rangle\mathrel{\mathop{\ordinarycolon}}=|{\Phi_{\ell,1,\bm{j}^{\prime}}}\rangle\otimes|{\Phi_{\ell,2,\bm{j}^{\prime\prime}}}\rangle,\quad|{\Psi_{\ell,\bm{k}}}\rangle\mathrel{\mathop{\ordinarycolon}}=|{\Psi_{\ell,1,\bm{k}^{\prime}}}\rangle\otimes|{\Psi_{\ell,2,\bm{k}^{\prime\prime}}}\rangle,$ (46)
$\bm{j}^{\prime}=(\bm{j}_{1},\cdots,\bm{j}_{T})$, $\bm{j}^{\prime\prime}=(\bm{j}_{T+1},\cdots,\bm{j}_{2T})$, similar to $\bm{k}^{\prime}$ and $\bm{k}^{\prime\prime}$. Now, we are ready to prove the upper bound for two terms defined in Eq. (35) as follows.
  1. 1.
    For the first term, we can also write it in the following form.
    $\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-p^{\mathcal{D}}\right\rVert_{{\rm TV}}$ $=\frac{1}{2}\sum_{\ell}\left|\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)-p^{\mathcal{D}}(\ell)\right|$ (47)
    $=\sum_{\ell\mathrel{\mathop{\ordinarycolon}}p^{\mathcal{D}}(\ell)\geq\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)}p^{\mathcal{D}}(\ell)\left[1-\frac{\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)}{p^{\mathcal{D}}(\ell)}\right].$ (48)
    Therefore, we can focus on the lower bound for $\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)/p^{\mathcal{D}}(\ell)$ and have
    $\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)$ $=(dd^{\prime})^{2T}W_{\ell}\sum_{\bm{j},\bm{k}}\langle{\Psi_{\ell,\bm{j}}\otimes\Phi_{\ell,\bm{j}}^{*}}|J_{H}^{(2T)}|{\Psi_{\ell,\bm{k}}\otimes\Phi_{\ell,\bm{k}}^{*}}\rangle$ (49)
    $\geq\frac{(dd^{\prime})^{2T}W_{\ell}}{d(d+1)\cdots(d+2T-1)}\sum_{\bm{j},\bm{k}}\langle{\Psi_{\ell,\bm{j}}\otimes\Phi_{\ell,\bm{j}}^{*}}|\left(\sum_{\sigma\in\mathcal{S}_{2T}}P_{\sigma}\otimes P_{\sigma}\right)|{\Psi_{\ell,\bm{k}}\otimes\Phi_{\ell,\bm{k}}^{*}}\rangle$ Lemma 10
    $=\frac{(dd^{\prime})^{2T}W_{\ell}}{d(d+1)\cdots(d+2T-1)}\sum_{\bm{j},\bm{k}}\langle{\Psi_{\ell,\bm{j}}}|\left(\sum_{\sigma\in\mathcal{S}_{2T}}P_{\sigma}|{\Psi_{\ell,\bm{k}}}\rangle\!\langle{\Phi_{\ell,\bm{k}}}|P_{\sigma}^{\dagger}\right)|{\Phi_{\ell,\bm{j}}}\rangle.$ (50)
    Define $\mathcal{P}_{\pi}\mathrel{\mathop{\ordinarycolon}}|X_{1}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\otimes|X_{2}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\otimes\cdots\otimes|X_{k}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\mapsto|X_{\pi^{-1}(1)}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\otimes|X_{\pi^{-1}(2)}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}\otimes\cdots\otimes|X_{\pi^{-1}(k)}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ for $\pi\in\mathcal{S}_{k}$ [41], then, we have
    $\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)$ $\geq\frac{(dd^{\prime})^{2T}W_{\ell}}{d(d+1)\cdots(d+2T-1)}\sum_{\bm{j},\bm{k}}\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}|{\Psi_{\ell,\bm{j}}}\rangle\!\langle{\Phi_{\ell,\bm{j}}}||\left(\sum_{\sigma\in\mathcal{S}_{2T}}\mathcal{P}_{\sigma}\right)||{\Psi_{\ell,\bm{k}}}\rangle\!\langle{\Phi_{\ell,\bm{k}}}|\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ (51)
    $\geq\frac{(dd^{\prime})^{2T}W_{\ell}}{d(d+1)\cdots(d+2T-1)}\sum_{\bm{j},\bm{k}}\langle\Psi_{\ell,\bm{j}}|\Psi_{\ell,\bm{k}}\rangle\langle\Phi_{\ell,\bm{k}}|\Phi_{\ell,\bm{j}}\rangle.$ Lemma 14
    Thus, for each leaf $\ell$, we have
    $\frac{\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)}{p^{\mathcal{D}}(\ell)}\geq\frac{d^{2T}}{d(d+1)\cdots(d+2T-1)}=\prod_{t=1}^{2T}\left(1+\frac{t-1}{d}\right)^{-1}\geq\prod_{t=1}^{2T}\left(1-\frac{t-1}{d}\right)\geq\left(1-\frac{2T}{d}\right)^{2T}\geq 1-\frac{4T^{2}}{d}.$ (52)
    Therefore, we have
    $\left\lVert\mathbb{E}_{U\sim\mu_{H}}p^{U,U}-p^{\mathcal{D}}\right\rVert_{{\rm TV}}$ $=\sum_{\ell\mathrel{\mathop{\ordinarycolon}}p^{\mathcal{D}}(\ell)\geq\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)}p^{\mathcal{D}}(\ell)\left[1-\frac{\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)}{p^{\mathcal{D}}(\ell)}\right]\leq\frac{4T^{2}}{d}.$ (53)
  2. 2.
    Likewise, we have
    $\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)$ $=(dd^{\prime})^{2T}W_{\ell}\prod_{i=1}^{2}\left(\sum_{\bm{j},\bm{k}}\langle{\Psi_{\ell,i,\bm{j}}\otimes\Phi_{\ell,i,\bm{j}}^{*}}|J_{H}^{(T)}|{\Psi_{\ell,i,\bm{k}}\otimes\Phi_{\ell,i,\bm{k}}^{*}}\rangle\right)$ (54)
    $\geq\left(\frac{(dd^{\prime})^{T}}{d(d+1)\cdots(d+T-1)}\right)^{2}W_{\ell}\prod_{i=1}^{2}\left(\sum_{\bm{j},\bm{k}}\langle\Psi_{\ell,i,\bm{j}}|\Psi_{\ell,i,\bm{k}}\rangle\langle\Phi_{\ell,i,\bm{k}}|\Phi_{\ell,i,\bm{j}}\rangle\right)$ (55)
    $=\left(\frac{(dd^{\prime})^{T}}{d(d+1)\cdots(d+T-1)}\right)^{2}W_{\ell}\sum_{\bm{j},\bm{k}}\langle\Psi_{\ell,\bm{j}}|\Psi_{\ell,\bm{k}}\rangle\langle\Phi_{\ell,\bm{k}}|\Phi_{\ell,\bm{j}}\rangle$ (56)
    Thus, for each leaf $\ell$, we have
    $\frac{\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)}{p^{\mathcal{D}}(\ell)}$ $\geq\left(\frac{d^{T}}{d(d+1)\cdots(d+T-1)}\right)^{2}\geq\left(1-\frac{T^{2}}{d}\right)^{2}.$ (57)
    Therefore, we have
    $\left\lVert p^{\mathcal{D}}-\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}\right\rVert_{{\rm TV}}\leq 1-\left(1-\frac{T^{2}}{d}\right)^{2}\leq\frac{2T^{2}}{d}.$ (58)
Lemma 14 (Lemma 5.12 in [44] and Lemma 3 in [41]).
For any product vector $|{X}\rangle=\bigotimes_{t=1}^{k}|{x_{t}}\rangle$, we have
$\langle{X}|\sum_{\sigma\in\mathcal{S}_{k}}P_{\sigma}|{X}\rangle\geq\langle X|X\rangle.$ (59)
B.2 Coherent Access
We now consider the lower bound for completing DSEC with coherent access.
B.2.1 Proof of Theorem 2
We first establish some notations. Recall that any learning algorithms with coherent access can be described as follows. Consider two unknown unitary channels $\mathcal{U}$ and $\mathcal{V}$ acting on two separate quantum systems. On these devices, we input the initial state $|{0}\rangle\otimes|{0}\rangle$ on $\mathcal{H}^{\otimes 2}$ and interleave the channel queries with a sequence of $T$ adaptive data-processing channels $\mathcal{C}_{t}=\mathcal{C}_{t,1}\otimes\mathcal{C}_{t,2}$, $t=1,\cdots,T$. Assuming the auxiliary system is large enough, each data-processing channel can be represented as a unitary $W_{t}=W_{t,1}\otimes W_{t,2}$ [57]. After $T$ queries, the resulting states on the two devices can then be expressed as
$|{\psi_{1}}\rangle$ $=\prod_{t=1}^{T}\left[\left(U\otimes I_{\rm aux}\right)W_{t,1}\right]|{0}\rangle,\quad|{\psi_{2}}\rangle=\prod_{t=1}^{T}\left[\left(V\otimes I_{\rm aux}\right)W_{t,2}\right]|{0}\rangle.$ (60)
As shown in [50], they can also be written as
$|{\psi_{1}}\rangle=\left(I_{\mathcal{H}}\otimes\langle{\Phi}|\right)\left(I_{\mathcal{H}}\otimes U^{\otimes T}\otimes I_{\rm main}^{\otimes T}\right)|{\Psi_{I,1}}\rangle,\quad|{\psi_{2}}\rangle=\left(I_{\mathcal{H}}\otimes\langle{\Phi}|\right)\left(I_{\mathcal{H}}\otimes V^{\otimes T}\otimes I_{\rm main}^{\otimes T}\right)|{\Psi_{I,2}}\rangle,$ (61)
where $|{\Phi}\rangle=\sum_{i=0}^{d^{T}-1}|{ii}\rangle$ is the maximally entangled state on system $\mathcal{H}_{\rm main}^{\otimes T}\otimes\mathcal{H}_{\rm main}^{\otimes T}$ and
$|{\Psi_{I,i}}\rangle=\sum_{\begin{subarray}{c}x_{1},\cdots,x_{T}\in\{0,\cdots,d-1\}\\ y_{1},\cdots,y_{T}\in\{0,\cdots,d-1\}\end{subarray}}\prod_{t=1}^{T}\overbrace{\left[\left(|{y_{t}}\rangle\!\langle{x_{t}}|\otimes I_{\rm aux}\right)W_{t,i}\right]|{0}\rangle}^{\text{$\mathcal{H}$}}\otimes\overbrace{|{x_{1}\cdots x_{T}}\rangle\otimes|{y_{1}\cdots y_{T}}\rangle}^{\text{$\mathcal{H}_{\rm main}^{\otimes T}\otimes\mathcal{H}_{\rm main}^{\otimes T}$}}.$ (62)
Lastly, we can perform an adaptive POVM $\{E_{\ell}\mathrel{\mathop{\ordinarycolon}}=E_{\ell,1}\otimes E_{\ell,2}\}_{\ell}$. Based on the above definition, we can define the following two probabilities in the following form:
$p^{U,V}(\ell)$ $\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}\left[\left(E_{\ell,1}\otimes J_{U}^{(T)}\right)\Psi_{I,1}\right]\operatorname{Tr}\left[\left(E_{\ell,2}\otimes J_{V}^{(T)}\right)\Psi_{I,2}\right],$ (63)
$p^{U,U}(\ell)$ $\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}\left[\left(E_{\ell,1}\otimes J_{U}^{(T)}\right)\Psi_{I,1}\right]\operatorname{Tr}\left[\left(E_{\ell,2}\otimes J_{U}^{(T)}\right)\Psi_{I,2}\right].$ (64)
where $J_{U}^{(T)}$ is the Choi operator of unitary $U^{\otimes T}$, as defined in Eq. (12), and similarly for $J_{V}^{(T)}$. With the definition of Haar random channel, we have
$\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)=\operatorname{Tr}\left[\left(E_{\ell,1}\otimes J_{H}^{(T)}\right)\Psi_{I,1}\right]\operatorname{Tr}\left[\left(E_{\ell,2}\otimes J_{H}^{(T)}\right)\Psi_{I,2}\right],$ (65)
where $J_{H}^{(T)}$ is defined in Eq. (20). As we do in the proof of lower bound for incoherent access, we define an intermediate probability
$p^{a}(\ell)\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}\left[\left(E_{\ell,1}\otimes J_{a}^{(T)}\right)\Psi_{I,1}\right]\operatorname{Tr}\left[\left(E_{\ell,2}\otimes J_{a}^{(T)}\right)\Psi_{I,2}\right],$ (66)
where $J_{a}^{(T)}$ is the Choi operator of $\mathcal{E}_{a}^{(T)}$, defined in Eq. (29). Then, with Lemma 15, we have
$\left\lVert\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)-\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)\right\rVert_{\rm TV}$ $\leq\left\lVert\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)-p^{a}(\ell)\right\rVert_{\rm TV}+\left\lVert p^{a}(\ell)-\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)\right\rVert_{\rm TV}$ (67)
$\leq\frac{T^{2}/2d}{1-T^{2}/2d}+\frac{4T^{2}}{d},$ (68)
which hints that we require $T=\Omega(\sqrt{d})$. Therefore, we complete the proof of Theorem 2.
B.2.2 Technical Lemmas
In the following, we prove key lemmas used in the above proof.
Lemma 15.
For two terms defined in Eq. (67), we have the following upper bound,
$\left\lVert\mathbb{E}p^{U,V\sim\mu_{H}}(\ell)-p^{a}(\ell)\right\rVert_{\rm TV}$ $\leq\frac{T^{2}/2d}{1-T^{2}/2d},\quad\left\lVert p^{a}(\ell)-\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)\right\rVert_{\rm TV}\leq\frac{4T^{2}}{d}.$ (69)
Proof of Lemma 15.
First, with the definition of $J_{a}^{(T)}$ (Eq. (29)), we have
$\quad\left|\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V}(\ell)-p^{a}(\ell)\right|$ (70)
$\leq\operatorname{Tr}\left[\left(E_{\ell_{1}}\otimes J_{H}^{(T)}\right){\Psi_{I,1}}\right]\left|\operatorname{Tr}\left[\left(E_{\ell_{2}}\otimes\left(J_{H}^{(a)}-J_{a}^{(T)}\right)\right){\Psi_{I,2}}\right]\right|+\left|\operatorname{Tr}\left[\left(E_{\ell_{1}}\otimes\left(J_{H}^{(T)}-J_{a}^{(T)}\right)\right){\Psi_{I,1}}\right]\right|\operatorname{Tr}\left[\left(E_{\ell_{2}}\otimes J_{a}^{(T)}\right){\Psi_{I,2}}\right]$ (71)
$\leq\frac{T^{2}/2d}{1-T^{2}/2d}\left\{\operatorname{Tr}\left[\left(E_{\ell_{1}}\otimes J_{H}^{(T)}\right){\Psi_{I,1}}\right]\operatorname{Tr}\left[\left(E_{\ell_{2}}\otimes J_{a}^{(T)}\right){\Psi_{I,2}}\right]+\operatorname{Tr}\left[\left(E_{\ell_{1}}\otimes J_{a}^{(T)}\right){\Psi_{I,1}}\right]\operatorname{Tr}\left[\left(E_{\ell_{2}}\otimes J_{H}^{(T)}\right){\Psi_{I,2}}\right]\right\}.$ (72)
Thus, we have
$\sum_{\ell}\left|\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V\sim\mu_{H}}(\ell)-p^{a}(\ell)\right|$ $\leq 2\frac{T^{2}/2d}{1-T^{2}/2d}\quad\Rightarrow\quad\left\lVert\mathbb{E}_{U,V\sim\mu_{H}}p^{U,V\sim\mu_{H}}(\ell)-p^{a}(\ell)\right\rVert_{\rm TV}\leq\frac{T^{2}/2d}{1-T^{2}/2d}.$ (73)
For the second term, we have
$\left\lVert p^{a}(\ell)-\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)\right\rVert_{\rm TV}$ $=\sum_{\ell\mathrel{\mathop{\ordinarycolon}}p^{a}(\ell)\geq\mathbb{E}p^{U,U}(\ell)}p^{a}(\ell)\left[1-\frac{\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)}{p^{a}(\ell)}\right]$ (74)
$=\sum_{\ell\mathrel{\mathop{\ordinarycolon}}p^{a}(\ell)\geq\mathbb{E}p^{U,U}(\ell)}\operatorname{Tr}\left[M_{\ell_{1}}J_{a}^{(T)}\right]\operatorname{Tr}\left[M_{\ell_{2}}J_{a}^{(T)}\right]\left[1-\frac{\operatorname{Tr}\left[M_{\ell_{2}}\Phi_{\ell}^{(2T)}\right]}{\operatorname{Tr}\left[M_{\ell_{2}}J_{a}^{(T)}\right]}\right].$ (75)
Here we define a “measure-and-prepare” map
$\Phi_{\ell}^{(2T)}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mu_{H}}\frac{\operatorname{Tr}\left[M_{\ell_{1}}J_{U}^{(T)}\right]}{\operatorname{Tr}\left[M_{\ell_{1}}J_{a}^{(T)}\right]}J_{U}^{(T)},\quad M_{\ell_{i}}\mathrel{\mathop{\ordinarycolon}}=\Pi_{{\rm sym}}^{\prime(d,T)}\operatorname{Tr}_{1}\left[\left(E_{\ell_{i}}\otimes I\right)\Psi_{I,i}\right]\Pi_{{\rm sym}}^{\prime(d,T)},\quad i=1,2,$ (76)
where $\Pi_{{\rm sym}}^{\prime(d,T)}\mathrel{\mathop{\ordinarycolon}}=\sum_{\pi\in\mathcal{S}_{T}}P_{\pi}\otimes P_{\pi}/T!$ is the projector onto the symmetric subspace of $\mathcal{H}_{\rm main}^{\otimes T}\otimes\mathcal{H}_{\rm main}^{\otimes T}$ [50]. Thus, with the definition, $M_{\ell_{i}}$ is in the symmetric subspace of $\mathcal{H}_{\rm main}^{\otimes T}\otimes\mathcal{H}_{\rm main}^{\otimes T}$ and we have [51, 50]
$J_{a}^{(T)}=\frac{T!}{d^{T}}\Pi_{{\rm sym}}^{\prime(d,T)}\quad\Rightarrow\quad\frac{\operatorname{Tr}\left[M_{\ell_{2}}\Phi_{\ell}^{(2T)}\right]}{\operatorname{Tr}\left[M_{\ell_{2}}J_{a}^{(T)}\right]}=\frac{d^{2T}}{(T!)^{2}\operatorname{Tr}\left[M_{\ell_{1}}\right]\operatorname{Tr}\left[M_{\ell_{2}}\right]}\mathbb{E}_{U\sim\mu_{H}}\operatorname{Tr}\left[M_{\ell_{1}}J_{U}^{(T)}\right]\operatorname{Tr}\left[M_{\ell_{2}}J_{U}^{(T)}\right].$ (77)
With Lemma 16, we have
$\frac{\operatorname{Tr}\left[M_{\ell_{2}}\Phi_{\ell}^{(2T)}\right]}{\operatorname{Tr}\left[M_{\ell_{2}}J_{a}^{(T)}\right]}\geq\frac{d^{2T}}{d(d+1)\cdots(d+2T-1)}\geq 1-\frac{4T^{2}}{d}.$ (78)
Therefore, we have
$\left\lVert p^{a}(\ell)-\mathbb{E}_{U\sim\mu_{H}}p^{U,U}(\ell)\right\rVert_{\rm TV}\leq\frac{4T^{2}}{d}.$ (79)
Lemma 16.
For quantum states $\rho,\sigma$ in the symmetric subspace of $\mathcal{H}^{\otimes T}\otimes\mathcal{H}^{\otimes T}$, we have
$\mathbb{E}_{U\sim\mu_{H}}\operatorname{Tr}\left[\rho\cdot J_{U}^{(T)}\right]\operatorname{Tr}\left[\sigma\cdot J_{U}^{(T)}\right]\geq\frac{(T!)^{2}}{d(d+1)\cdots(d+2T-1)}$ (80)
Figure 2: The visualization of key proof steps. (a) The visualization of permutation map $\mathcal{R}$ (defined in Eq. (81)) when $T=3$. (b) The visualization of $\mathcal{R}^{\dagger}$ when $T=3$. (c) The visualization of Eq. (84) when $T=2$. (d) The visualization of Eq. (86) when $T=2$.
Proof of Lemma 16.
For convenience, we also provide the visualization of the key proof steps in Fig. 2. Now, we begin our proof. First, we define the following permutation map,
$\mathcal{R}\mathrel{\mathop{\ordinarycolon}}\mathcal{D}\left(\mathcal{H}^{\otimes T}\otimes\mathcal{H}^{\otimes T}\right)\to\mathcal{D}\left((\mathcal{H}^{\otimes 2})^{\otimes T}\right),\quad|{\alpha}\rangle^{\otimes T}\otimes|{\beta}\rangle^{\otimes T}\mapsto\left(|{\alpha}\rangle\otimes|{\beta}\rangle\right)^{\otimes T},$ (81)
which can map a state in the symmetric subspace of $\mathcal{H}^{\otimes T}\otimes\mathcal{H}^{\otimes T}$ to the symmetric subspace of $(\mathcal{H}^{\otimes 2})^{\otimes T}$. This map $\mathcal{R}$ is visualized in Fig 2(a). With this definition, we have
$\mathcal{R}\left(\Pi_{\rm sym}^{\prime(d,T)}\right)=\Pi_{\rm sym}^{(d^{2},T)}.$ (82)
Then, we can observe that the space of density matrices on the symmetric subspace of $(\mathcal{H}^{\otimes 2})^{\otimes T}$ is spanned by $\alpha^{\otimes T}$, where $\alpha\in\mathcal{D}(\mathcal{H}^{\otimes 2})$ [58, 51]. Here we write $\alpha\mathrel{\mathop{\ordinarycolon}}=|{\alpha}\rangle\!\langle{\alpha}|$ for pure state $|{\alpha}\rangle$. Thus, to compute the lower bound, it suffices to calculate the following function
$f(\alpha,\beta)$ $\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mu_{H}}\operatorname{Tr}\left[\beta^{\otimes T}\mathcal{R}\left(J_{U}^{(T)}\right)\right]\operatorname{Tr}\left[\alpha^{\otimes T}\mathcal{R}\left(J_{U}^{(T)}\right)\right]=\mathbb{E}_{U\sim\mu_{H}}\operatorname{Tr}\left[\left(\mathcal{R}^{\dagger}(\alpha^{\otimes T})\otimes\mathcal{R}^{\dagger}(\beta^{\otimes T})\right)\left(J_{U}^{(T)}\otimes J_{U}^{(T)}\right)\right],$ (83)
where $\mathcal{R}^{\dagger}$ is the inverse map of $\mathcal{R}$ and is visualized in Fig. 2(b). With the permutation operations shown in Fig. 2(c), we can write $f(\alpha,\beta)$ as
$f(\alpha,\beta)$ $=\operatorname{Tr}\left[\mathbb{F}_{2,3}\left(\mathcal{R}^{\dagger}(\alpha^{\otimes T})\otimes\mathcal{R}^{\dagger}(\beta^{\otimes T})\right)\mathbb{F}_{2,3}\left(\sum_{\pi,\sigma\in\mathcal{S}_{2T}}{\rm Wg}_{\pi,\sigma}P_{\pi}\otimes P_{\sigma}\right)\right],$ (84)
where $\mathbb{F}_{2,3}$ is acting on the second and third $\mathcal{H}^{\otimes T}$. With Lemma 10, we have
$f(\alpha,\beta)$ $\geq\frac{1}{d(d+1)\cdots(d+2T-1)}\operatorname{Tr}\left[\mathbb{F}_{2,3}\left(\mathcal{R}^{\dagger}(\alpha^{\otimes T})\otimes\mathcal{R}^{\dagger}(\beta^{\otimes T})\right)\mathbb{F}_{2,3}\left(\sum_{\pi\in\mathcal{S}_{2T}}P_{\pi}\otimes P_{\pi}\right)\right]$ (85)
$=\frac{(2T!)}{d(d+1)\cdots(d+2T-1)}\operatorname{Tr}\left[\Pi_{\rm sym}^{(d^{2},2T)}\left(\alpha^{\otimes T}\otimes\beta^{\otimes T}\right)\right]$ (86)
$=\frac{(2T!)}{d(d+1)\cdots(d+2T-1)}\sum_{s=0}^{T}\frac{\binom{T}{s}^{2}}{\binom{2T}{T}}x^{s}\geq\frac{(T!)^{2}}{d(d+1)\cdots(d+2T-1)},$ (87)
where $x\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}[\alpha\beta]$ and Eq. (86) is visualized in Fig. 2(d). The term on the last line is the probability that a random $\pi\in\mathcal{S}_{2T}$ satisfies $|\pi(\{1,\cdots,T\})\cap\{1,\cdots,T\}|=s$ [51]. The corresponding physical meaning is that when mapping the first $T$ $\alpha$ terms to $2T$ positions, the overlap term will be $x$, and the other terms are all $1$. This is equivalent to the probability that when $T$ balls are drawn without replacement from a bucket of $T$ white balls and $T$ black balls, the resulting sample contains TsT-s white balls and $s$ black balls. ∎
B.3 Lower bound for DSEC
Furthermore, we consider the lower bound for DSEC. The proof is similar to [15]. We first define a distinguishing problem for estimation, then analyze the lower bound for solving this problem with the result in Appendix B.2.
B.3.1 Distinguishing Problem for Estimation
Suppose that two quantum devices perform two unknown channels $\mathcal{E}$ and $\mathcal{F}$, and there is an estimation algorithm $\mathbb{A}$ that completes DSEC up to error $\varepsilon/100$ with probability at least $0.99$ with $T$ queries on each device. We try to solve the following distinguishing problem with this estimation algorithm $\mathbb{A}$.
Problem 2 (Distinguishing Problem for Estimation).
In Hilbert space $\mathcal{H}_{\rm main}$ with dimension $d$, we want to distinguish the following two cases with $T$ queries on each device:
  1. 1.
    Two quantum devices implement two operators $\sqrt{1-\varepsilon}e^{i\theta_{1}}I+\sqrt{\varepsilon}U$ and $\sqrt{1-\varepsilon}e^{i\theta_{2}}I+\sqrt{\varepsilon}U$, respectively, where $U$ is a Haar random unitary in $\mathcal{H}_{\rm main}$ and $\theta_{1},\theta_{2}$ are independent random phases in $0,2\pi]$.
  2. 2.
    Two quantum devices independently implement two operators $\sqrt{1-\varepsilon}e^{i\theta_{1}}I+\sqrt{\varepsilon}U$ and $\sqrt{1-\varepsilon}e^{i\theta_{2}}I+\sqrt{\varepsilon}V$, respectively, where $U,V$ are two independent Haar random unitaries in $\mathcal{H}_{\rm main}$ and $\theta_{1},\theta_{2}$ are independent random phases in $0,2\pi]$.
When $d$ is large enough, case 1 gives
$f_{1}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U\sim\mu_{H}}\frac{\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}]}{d^{2}}\approx\left|(1-\varepsilon)e^{i(\theta_{1}-\theta_{2})}+\varepsilon\right|^{2}=1-2\varepsilon+2\varepsilon^{2}+2\varepsilon(1-\varepsilon)\cos(\theta_{1}-\theta_{2}),$ (88)
and case 2 gives
$f_{2}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U,V\sim\mu_{H}}\frac{\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}]}{d^{2}}\approx(1-\varepsilon)^{2},$ (89)
with the fact that
$\mathbb{E}_{U,V\sim\mu_{H}}\operatorname{Tr}^{2}[U^{\dagger}V]$ $=\mathbb{E}_{U\sim\mu_{H}}\operatorname{Tr}[U\otimes U^{*}]=1.$ (90)
In case 1, $f_{1}$ is distributed roughly as $1-2\varepsilon+2\varepsilon\cos(\gamma)$ for a uniform random phase $\gamma\in[0,2\pi]$; in case 2, $f_{2}$ is concentrated around $(1-\varepsilon)^{2}$. Therefore, if we have an estimation algorithm that can complete DSEC, then we can solve the above distinguishing problem with the following procedure:
  1. 1.
    Obtain the estimation result $f$ with the estimation algorithm $\mathbb{A}$.
  2. 2.
    Output case 2 if $f\in[(1-\varepsilon)^{2}-\varepsilon/50,(1-\varepsilon)^{2}+\varepsilon/50]$, and output case 1 otherwise.
The performance of the estimation algorithm $\mathbb{A}$ guarantees that if the result $f$ is in case $i(i=1,2)$, then $|f-f_{i}|\leq\varepsilon/100$ holds with probability at least $0.99$. Then, if $f$ is from case 1, we have [15]
$\Pr\left\{f\in[(1-\varepsilon)^{2}-\varepsilon/50,(1-\varepsilon)^{2}+\varepsilon/50]\right\}$ $\leq\Pr\left\{f\in[(1-\varepsilon)^{2}-\varepsilon/50,(1-\varepsilon)^{2}+\varepsilon/50]\wedge|f-f_{1}|\leq\frac{\varepsilon}{100}\right\}+0.01$ (91)
$\leq\Pr\left\{f_{1}\in[(1-\varepsilon)^{2}-3\varepsilon/100,(1-\varepsilon)^{2}+3\varepsilon/100]\right\}+0.01$ (92)
$\leq\Pr_{\gamma\sim[0,2\pi]}\left\{\cos(\gamma)\in[-0.0203,0.0152]\right\}+0.01$ (93)
$\leq 0.03.$ (94)
Therefore, the above procedure returns case 1 with probability at least $0.97$ and can solve Problem 2 with success probability at least $2/3$. In the following, we analyze the lower bound for solving Problem 2.
As shown in Appendix B.2, given $T$ queries on each device, we can only focus on two Choi operators $J_{1}^{(T)}$ and $J_{2}^{(T)}$ of $(\sqrt{1-\varepsilon}e^{i\theta_{1}}I+\sqrt{\varepsilon}U)^{\otimes T}$ and $(\sqrt{1-\varepsilon}e^{i\theta_{2}}I+\sqrt{\varepsilon}V)^{\otimes T}$, respectively. With the definition of map $\mathcal{R}$ in Eq. (81) and [15, Lemma 10], we have
$\mathbb{E}_{\theta_{1}\sim[0,2\pi]}\mathcal{R}\left(J_{1}^{(T)}\right)$ $=\mathbb{E}_{\theta_{1}\sim[0,2\pi]}\left[\left(\sqrt{1-\varepsilon}e^{i\theta_{1}}I\otimes I+\sqrt{\varepsilon}U\otimes I\right)(|{\Phi_{1}}\rangle\!\langle{\Phi_{1}}|)\left(\sqrt{1-\varepsilon}e^{i\theta_{1}}I\otimes I+\sqrt{\varepsilon}U^{\dagger}\otimes I\right)\right]^{\otimes T}$ (95)
$=\sum_{t=0}^{T}(1-\varepsilon)^{T-t}\varepsilon^{t}\binom{T}{t}\mathcal{R}\left(|{\Phi_{U,t}}\rangle\!\langle{\Phi_{U,t}}|\right),$ (96)
where $|{\Phi_{1}}\rangle$ is the unnormalized maximally entangled state in $\mathcal{H}^{\otimes 2}_{\rm main}$ and
$|{\Phi_{U,t}}\rangle=\frac{1}{\sqrt{\binom{T}{t}}}\sum_{S\subseteq[T],|S|=t}\left(U_{S}\otimes I^{\otimes T}\right)|{\Phi}\rangle.$ (97)
Here $U_{S}$ means applying operator $U$ on the subsystems indexed by set $S$ and $T]\mathrel{\mathop{\ordinarycolon}}=\{1,\cdots,T\}$. Thus, $J_{1}^{(T)}$ can be written as
$\mathbb{E}_{\theta_{1}\sim[0,2\pi]}J_{1}^{(T)}=\mathbb{E}_{t\sim B(T,\varepsilon)}|{\Phi_{U,t}}\rangle\!\langle{\Phi_{U,t}}|.$ (98)
Likewise, we have
$\mathbb{E}_{\theta_{2}\sim[0,2\pi]}J_{2}^{(T)}=\mathbb{E}_{t\sim B(T,\varepsilon)}|{\Phi_{V,t}}\rangle\!\langle{\Phi_{V,t}}|.$ (99)
The corresponding output states can be written as
$|{\psi_{1}}\rangle=\mathbb{E}_{t\sim B(T,\varepsilon)}(I_{\mathcal{H}}\otimes\langle{\Phi_{U,t}}|)|{\Psi_{I,1}}\rangle,\quad|{\psi_{2}}\rangle=\mathbb{E}_{t^{\prime}\sim B(T,\varepsilon)}(I_{\mathcal{H}}\otimes\langle{\Phi_{V,t^{\prime}}}|)|{\Psi_{I,2}}\rangle,$ (100)
where $|{\Psi_{I,2}}\rangle$ is defined in Eq. (62). Based on the above definition, we can define the following distinguishing problem.
Problem 3.
In Hilbert space $\mathcal{H}_{\rm main}$ with dimension $d$, we want to distinguish the following two cases:
  1. 1.
    Two quantum devices perform $T$ queries and the corresponding Choi operators are $|{\Phi_{U,t}}\rangle\!\langle{\Phi_{U,t}}|$ and $|{\Phi_{U,t^{\prime}}}\rangle\!\langle{\Phi_{U,t^{\prime}}}|$, where $t,t^{\prime}\sim B(T,\varepsilon)$ and $U$ is a Haar random unitary in $\mathcal{H}_{\rm main}$.
  2. 2.
    Two quantum devices perform $T$ queries and the corresponding Choi operators are $|{\Phi_{U,t}}\rangle\!\langle{\Phi_{U,t}}|$ and $|{\Phi_{V,t^{\prime}}}\rangle\!\langle{\Phi_{V,t^{\prime}}}|$, where $t,t^{\prime}\sim B(T,\varepsilon)$ and $U,V$ are two independent Haar random unitaries in $\mathcal{H}_{\rm main}$.
It should be noted that Problem 2 is equivalent to Problem 3, as these two problems have the same average inputs. This equivalence indicates that a lower bound for Problem 3 also applies to Problem 2.
B.3.2 Proof of Theorem 3
Now, we can prove an $\Omega(\sqrt{d}/\varepsilon)$ lower bound for Problem 3 via a reduction to Problem 1. This reduction is achieved by simulating the input of Problem 3 with the input of Problem 1. Suppose that there is an algorithm $\mathbb{A}^{\prime}$ that can solve Problem 2 with success probability at least $0.99$. Now we are trying to solve Problem 1 with $K=100T\varepsilon+100$ queries on each device using the algorithm $\mathbb{A}^{\prime}$. The procedure is as follows:
  1. 1.
    For device 1, randomly sample $t\sim B(T,\varepsilon)$.
    1. (a)
      If $t\leq K$, we construct the queries corresponding to $|{\Phi_{U,t}}\rangle\!\langle{\Phi_{U,t}}|$ with $t$ queries of unknown operator $U$ based on Lemma 17;
    2. (b)
      Otherwise, all $T$ queries are identity channel queries.
  2. 2.
    For device 2, we perform the same procedure as device 1.
  3. 3.
    Run the algorithm $\mathbb{A}^{\prime}$ and decide accordingly.
It should be noted that we can simulate the input of Problem 3 with high probability, since $K=100T\varepsilon+100\Rightarrow t\leq K$ with high probability. Lastly, as shown in Appendix B.2, to complete Problem 1, we require
$100T\varepsilon+100=\Omega(\sqrt{d})\quad\Rightarrow\quad T=\Omega\left(\frac{\sqrt{d}}{\varepsilon}\right).$ (101)
On the other hand, in the non-distributed scenario (directly performing the SWAP test on two Choi states), we require $T=\Omega(1/\varepsilon^{2})$ queries [41]. Combining these two results, we complete the proof of Theorem 3, i.e.,
$T=\Omega\left(\max\left\{\frac{\sqrt{d}}{\varepsilon},\frac{1}{\varepsilon^{2}}\right\}\right).$ (102)
Lemma 17.
For any $t\in[T]$, there exists a procedure to construct the queries corresponding to $|{\Phi_{U,t}}\rangle\!\langle{\Phi_{U,t}}|$ with $t$ queries of unknown operator $U$.
Proof of Lemma 17.
We want to construct a procedure that performs the following mapping:
$(I_{\mathcal{H}}\otimes\langle{\Phi}|)(I_{\mathcal{H}}\otimes U^{\dagger\otimes t}\otimes I^{\otimes 2T-t})|{\Psi_{I}}\rangle\mapsto(I_{\mathcal{H}}\otimes\langle{\Phi_{U,t}}|)|{\Psi_{I,1}}\rangle,$ (103)
Therefore, if we can construct the vector $|{\Phi_{U,t}}\rangle$ from $|{\Phi}\rangle$ with $t$ queries of $U$, then we can implement the above mapping. To this end, we introduce an additional auxiliary system $R$ in the uniform superposition
$|{\phi_{\operatorname{aux}}}\rangle\mathrel{\mathop{\ordinarycolon}}=\frac{1}{\sqrt{\binom{T}{t}}}\sum_{S\subseteq[T],|S|=t}|{S}\rangle_{R}.$ (104)
We define a controlled-swap unitary $W=\sum_{S}|{S}\rangle\!\langle{S}|_{R}\otimes W_{S}\otimes W_{S}$, where $W_{S}$ acts on $\mathcal{H}_{\rm main}^{\otimes T}$ and permutes the systems such that the first $t$ subsystems are moved to the subsystems indexed by $S$. We then construct $|{\Phi_{U,t}}\rangle$ by inputting $|{\phi_{\operatorname{aux}}}\rangle\otimes(U^{\otimes t}\otimes I^{\otimes 2T-t}|{\Phi}\rangle)$ and performing $W$. ∎
Appendix C Upper Bound
In this section, we provide the detailed proofs of the upper bounds for completing DSEC.
C.1 Coherent Access: Proof of Theorem 4
Suppose two quantum devices perform two unitary channels $\mathcal{U}(\cdot)=U(\cdot)U^{\dagger}$ and $\mathcal{V}(\cdot)=V(\cdot)V^{\dagger}$, respectively. Let $J_{U}$ and $J_{V}$ be the Choi operators of channels $\mathcal{U}$ and $\mathcal{V}$. The algorithm solving DSEC with coherent access is summarized in Algorithm 1. In the following, we will prove Theorem 4, which concludes the query complexity of Algorithm 1.
Input: $T$ queries to unknown unitary channels $\mathcal{U}$ and $\mathcal{V}$ acting on a $d$-dimensional Hilbert space $\mathcal{H}$.
Output: an estimation of $\operatorname{Tr}[J_{U}J_{V}]/d^{2}$.
1
2Randomly generate pure states $|{\psi}\rangle^{\otimes T}$ on each device, where $|{\psi}\rangle$ is sampled from a state $4$-design ensemble.
3Apply the unitary channels $\mathcal{U}^{\otimes T}$ and $\mathcal{V}^{\otimes T}$ on $|{\psi}\rangle^{\otimes T}$.
4Measure $(\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes T}$ with the POVM $\mathcal{M}_{T}$ and obtain result $|{\phi_{A}}\rangle$.
5Measure $(\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes T}$ with the POVM $\mathcal{M}_{T}$ and obtain result $|{\phi_{B}}\rangle$.
Compute $\tilde{f}=|\langle\phi_{A}|\phi_{B}\rangle|^{2}$ and return
$\tilde{\chi}\mathrel{\mathop{\ordinarycolon}}=\frac{(d+1)(d+T)^{2}}{T^{2}d}\tilde{f}-\frac{(d+1)(d+2T)+T^{2}}{T^{2}d}.$ (105)
Algorithm 1 Distributed Similarity Estimation for Unitary Channels with Coherent Access
To obtain the expectation of $\tilde{\chi}$, we compute the expectation of $\tilde{f}$ first:
$\mathbb{E}\tilde{f}$ $=\mathbb{E}|\langle\phi_{A}|\phi_{B}\rangle|^{2}=\mathbb{E}\operatorname{Tr}\left[\left(|{\phi_{A}}\rangle\!\langle{\phi_{A}}|\right)\left(|{\phi_{B}}\rangle\!\langle{\phi_{B}}|\right)\right]$ (106)
$=\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(\frac{I+TU|{\psi}\rangle\!\langle{\psi}|U^{\dagger}}{d+T}\right)\left(\frac{I+TV|{\psi}\rangle\!\langle{\psi}|V^{\dagger}}{d+T}\right)\right]$ Lemma 9
$=\frac{d+2T}{(d+T)^{2}}+\frac{T^{2}}{(d+T)^{2}}\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)|{\psi}\rangle\!\langle{\psi}|^{\otimes 2}\right]$ (107)
$=\frac{d+2T}{(d+T)^{2}}+\frac{T^{2}\left(\operatorname{Tr}^{2}[U^{\dagger}V]+d\right)}{d(d+1)(d+T)^{2}}.$ Lemma 8
As we can see, the proof relies on Lemma 9, which requires that the output states remain pure. For a general quantum channel, the output states might be mixed, and the key identities in the lemma no longer hold. Therefore, the coherent algorithm is applicable only when the unknown channels are unitary channels. Lastly, we can prove $\tilde{\chi}$ is an unbiased estimator for $\operatorname{Tr}[J_{U}J_{V}]/d^{2}$,
$\mathbb{E}\tilde{\chi}=\frac{(d+1)(d+T)^{2}}{T^{2}d}\mathbb{E}\tilde{f}-\frac{(d+1)(d+2T)+T^{2}}{T^{2}d}=\frac{\operatorname{Tr}^{2}[U^{\dagger}V]}{d^{2}}=\frac{\operatorname{Tr}[J_{U}J_{V}]}{d^{2}}.$ (108)
with the fact that $\operatorname{Tr}[J_{U}J_{V}]=\operatorname{Tr}^{2}[U^{\dagger}V]$ (special case of Eq. (15)).
We now consider the variance of the estimator $\tilde{\chi}$. Let $\tilde{\chi}=X\tilde{f}-Y$, where
$X=\frac{(d+1)(d+T)^{2}}{T^{2}d},\quad Y=\frac{(d+1)(d+2T)+T^{2}}{T^{2}d}.$ (109)
Then, we have
[…]
We consider the first two terms, respectively. With [15, Proof of Lemma 5, Eqs. (165) and (166)], we have
$X^{2}\mathbb{E}\tilde{f}^{2}$ $\leq\left(\frac{d+1}{d}\right)^{2}\mathbb{E}_{\psi}\left(f_{\psi}^{2}+\frac{8f_{\psi}-2f_{\psi}^{2}}{T}+\frac{2df_{\psi}+f_{\psi}^{2}+8+2d}{T^{2}}+\frac{8d+4}{T^{3}}+\frac{2d^{2}+2d}{T^{4}}\right),$ (111)
where we define
$f_{\psi}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}\left[\left(U|{\psi}\rangle\!\langle{\psi}|U^{\dagger}\right)\left(V|{\psi}\rangle\!\langle{\psi}|V^{\dagger}\right)\right].$ (112)
With Lemma 8, we have
$\mathbb{E}_{\psi}f_{\psi}=\frac{\operatorname{Tr}^{2}[U^{\dagger}V]+d}{d(d+1)}\leq 1,$ (113)
and
$\mathbb{E}_{\psi}f_{\psi}^{2}$ $=\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)^{\otimes 2}|{\psi}\rangle\!\langle{\psi}|^{\otimes 4}\right]=\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)^{\otimes 2}\frac{\Pi_{\rm sym}^{(4)}}{\kappa_{4}}\right]$ (114)
$=\frac{1}{d(d+1)(d+2)(d+3)}\sum_{\pi\in\mathcal{S}_{4}}\operatorname{Tr}\left[\left(U^{\dagger}V\otimes V^{\dagger}U\right)^{\otimes 2}P_{\pi}\right]\leq\frac{\operatorname{Tr}^{4}[U^{\dagger}V]}{d^{4}}+\mathcal{O}\left(\frac{1}{d}\right).$ (115)
For the second term, we have
$Y^{2}=\left(\frac{d}{T^{2}}+\frac{2T+1}{T^{2}}+\frac{2}{Td}+\frac{1}{d}\right)^{2}=\mathcal{O}\left(\frac{d^{2}}{T^{4}}+\frac{1}{T^{2}}+\frac{1}{d^{2}}\right).$ (116)
Thus, the variance of $\tilde{\chi}$ is upper bounded by
$\mathbb{V}(\tilde{\chi})\leq\mathcal{O}\left(\frac{1}{T}+\frac{d}{T^{2}}+\frac{d}{T^{3}}+\frac{d^{2}}{T^{4}}\right)+\mathcal{O}\left(\frac{d^{2}}{T^{4}}+\frac{1}{T^{2}}+\frac{1}{d^{2}}\right)=\mathcal{O}\left(\frac{1}{T}+\frac{d}{T^{2}}+\frac{d^{2}}{T^{4}}\right).$ (117)
Therefore, to achieve $\varepsilon$ additive error, we require
$T\geq\mathcal{O}\left(\max\left\{\frac{1}{\varepsilon^{2}},\frac{\sqrt{d}}{\varepsilon}\right\}\right).$ (118)
C.2 Incoherent Access: Proof of Theorem 5
Suppose two quantum devices perform two channels $\mathcal{E}$ and $\mathcal{F}$, respectively. The algorithm solving DSEC with incoherent access is summarized in Algorithm 2. In the following, we will prove Theorem 5.
Input: number of SPAM settings $T$,
    number of measurements for each SPAM setting $m$,
    TmTm queries to unknown channels $\mathcal{E}$ and $\mathcal{F}$ acting on a $d$-dimensional Hilbert space $\mathcal{H}$.
Output: an estimation of $\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}]/d^{2}$.
1 for $t=1,\cdots,T$ do
2   Randomly generate pure states $|{\psi_{t}}\rangle$ from a state $4$-design ensemble on each quantum device.
3  Apply the unknown channels $\mathcal{E}$ and $\mathcal{F}$ on $|{\psi_{t}}\rangle$.
4  Sample a random unitary $Q_{t}$ from a unitary $4$-design ensemble.
5  Measure $m$ copies of $\mathcal{E}(|{\psi_{t}}\rangle\!\langle{\psi_{t}}|)$ in the basis $\{Q_{t}^{\dagger}|{a}\rangle\!\langle{a}|Q_{t}\}_{a}$ and obtain $A=\{a_{i}\}_{i=1}^{m}$.
6  Measure $m$ copies of $\mathcal{F}(|{\psi_{t}}\rangle\!\langle{\psi_{t}}|)$ in the basis $\{Q_{t}^{\dagger}|{b}\rangle\!\langle{b}|Q_{t}\}_{b}$ and obtain $B=\{b_{i}\}_{i=1}^{m}$.
7  Compute $\tilde{g}_{t}=\sum_{i,j=1}^{m}\mathbb{1}(a_{i},b_{j})/m^{2}$, where $\mathbb{1}(a,b)=1$ if $a=b$; otherwise 0.
8  Compute
$\tilde{\omega}_{t}=\frac{(d+1)^{2}}{d}\tilde{g}_{t}-\frac{d+2}{d}.$ (119)
Return ${\omega}\mathrel{\mathop{\ordinarycolon}}=\sum_{t}\tilde{\omega}_{t}/T$.
Algorithm 2 Distributed Similarity Estimation for General Channels with Incoherent Access
For each SPAM setting $\{Q,\psi\}$, we define the following two probabilities:
$p_{Q,\psi}(a)\mathrel{\mathop{\ordinarycolon}}=\langle{a}|Q\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle,\quad q_{Q,\psi}(b)\mathrel{\mathop{\ordinarycolon}}=\langle{b}|Q\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{b}\rangle,$ (120)
and their classical inner product function
$g(Q,\psi)\mathrel{\mathop{\ordinarycolon}}=\sum_{a}p_{Q,\psi}(a)q_{Q,\psi}(a).$ (121)
Then, the expectation of $\tilde{g}$ can be represented as
$\mathbb{E}_{Q,\psi,A,B}\;\tilde{g}$ $=\mathbb{E}_{Q,\psi}g(Q,\psi)$ (122)
$=\sum_{a}\langle{a}|Q\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle\langle{a}|Q\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle$ (123)
$=\mathbb{E}_{\psi}\operatorname{Tr}\left[\left(\sum_{a}\mathbb{E}_{Q}Q^{\dagger\otimes 2}|{aa}\rangle\!\langle{aa}|Q^{\otimes 2}\right)\left(\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)\otimes\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)\right)\right]$ (124)
$=\frac{1}{d+1}+\frac{1}{d+1}\mathbb{E}_{\psi}\operatorname{Tr}\left[\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)\right],$ (125)
where we have used Lemma 8 and the property of unitary $2$-design. For the last term, suppose that $\mathcal{E}(\cdot)=\sum_{i}E_{i}(\cdot)E_{i}^{\dagger}$ and $\mathcal{F}(\cdot)=\sum_{j}F_{j}(\cdot)F_{j}^{\dagger}$. Then we have
$\mathbb{E}_{\psi}\operatorname{Tr}\left[\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)\right]$ $=\sum_{i,j}\mathbb{E}_{\psi}\operatorname{Tr}\left[E_{i}|{\psi}\rangle\!\langle{\psi}|E_{i}^{\dagger}F_{j}|{\psi}\rangle\!\langle{\psi}|F_{j}^{\dagger}\right]$ (126)
$=\sum_{i,j}\operatorname{Tr}\left[\left(E_{i}^{\dagger}F_{j}\otimes F_{j}^{\dagger}E_{i}\right)\left(\mathbb{E}_{\psi}|{\psi}\rangle\!\langle{\psi}|^{\otimes 2}\right)\right]$ (127)
$=\sum_{i,j}\operatorname{Tr}\left[\left(E_{i}^{\dagger}F_{j}\otimes F_{j}^{\dagger}E_{i}\right)\frac{\Pi_{\rm sym}^{(d,2)}}{\kappa_{2}}\right]=\frac{\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}]+d}{d(d+1)}.$ (128)
Therefore, we have
$\quad\mathbb{E}_{Q,\psi,A,B}\;\tilde{g}=\mathbb{E}_{Q,\psi}g(Q,\psi)=\frac{1}{d+1}+\frac{1}{d}\frac{\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}]+d}{d(d+1)},$ (129)
$\Rightarrow$ $\quad\frac{(d+1)^{2}}{d}\mathbb{E}\tilde{g}-\frac{d+2}{d}=\frac{\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}]}{d^{2}}$ (130)
$\Rightarrow$ $\quad\mathbb{E}\tilde{\omega}_{t}=\frac{\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}]}{d^{2}}.$ (131)
Now, we consider the variance of the estimator so that we can obtain the query complexity of Algorithm 2. The following analysis is similar to [15]. With the definition of $\omega$}, the variance of $\omega$} is given by
[…]
Lemma 18.
The variance of $\tilde{g}$ with respect to SPAM setting $\{Q,\psi\}$ and measurement results $S\mathrel{\mathop{\ordinarycolon}}=\{A,B\}$ can be upper bounded by
$\mathbb{V}(\tilde{g})\leq\mathcal{O}\left(\frac{1}{m^{2}d}+\frac{1}{md^{2}}+\frac{1}{d^{3}}\right).$ (133)
Proof of Lemma 18.
It should be noted that $\tilde{g}$ is a function of the SPAM setting $\{Q,\psi\}$ and measurement results $S\mathrel{\mathop{\ordinarycolon}}=\{A,B\}$. Thus, we can write $\tilde{g}$ as $\tilde{g}(Q,\psi,S)$, and the law of total variance gives [15]
$\mathbb{V}(\tilde{g}(Q,\psi,S))=\mathbb{E}_{Q,\psi}\mathbb{V}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right]+\mathbb{V}_{Q,\psi}\mathbb{E}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right].$ (134)
We consider these two terms separately as follows.
  1. 1.
    For the first term, with Lemma 14 in [15], we have
    $\mathbb{E}_{Q,\psi}\mathbb{V}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right]$ $\leq\mathbb{E}_{Q,\psi}\left[\frac{g(Q,\psi)}{m^{2}}+\frac{1}{m}\sum_{a}\left(p_{Q,\psi}^{2}(a)q_{Q,\psi}(a)+p_{Q,\psi}(a)q_{Q,\psi}^{2}(a)\right)\right].$ (135)
    As shown in Eq. (129), we have
    $\mathbb{E}_{Q,\psi}\frac{g(Q,\psi)}{m^{2}}=\frac{1}{m^{2}(d+1)}+\frac{1}{m^{2}d}\frac{\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}]+d}{d(d+1)}=\mathcal{O}\left(\frac{1}{m^{2}d}\right).$ (136)
    Additionally, with Eq. (187) in [15], we have
    $\mathbb{E}_{Q}p_{Q,\psi}^{2}(a)q_{Q,\psi}(a)$ $=\mathbb{E}_{Q}\sum_{a}\langle{a}|Q\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle^{2}\langle{a}|Q\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle$ (137)
    $=d\mathbb{E}_{\phi}\langle{\phi}|\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}\langle{\phi}|\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle=\mathcal{O}\left(\frac{1}{d^{2}}\right),$ (138)
    where we use the property of unitary $3$-design. Thus, we have
    $\mathbb{E}_{Q,\psi}p_{Q,\psi}^{2}(a)q_{Q,\psi}(a)=\mathcal{O}\left(\frac{1}{d^{2}}\right),\quad\mathbb{E}_{Q,\psi}p_{Q,\psi}(a)q_{Q,\psi}^{2}(a)=\mathcal{O}\left(\frac{1}{d^{2}}\right).$ (139)
    Therefore, we have
    $\mathbb{E}_{Q,\psi}\mathbb{V}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right]\leq\mathcal{O}\left(\frac{1}{m^{2}d}+\frac{1}{md^{2}}\right).$ (140)
  2. 2.
    For the second term, we have
    $\mathbb{V}_{Q,\psi}\mathbb{E}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right]$ $=\mathbb{V}_{Q,\psi}\;g(Q,\psi)=\mathbb{E}_{Q,\psi}\;g^{2}(Q,\psi)-\left[\mathbb{E}_{Q,\psi}\;g(Q,\psi)\right]^{2}$ (141)
    $=\mathbb{E}_{Q,\psi}\;g^{2}(Q,\psi)-\mathbb{E}_{\psi}\left(\frac{1+\operatorname{Tr}[\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)]}{d+1}\right)^{2}.$ (142)
    We have
    $\mathbb{E}_{Q}\;g^{2}(Q,\psi)=\mathbb{E}_{Q}\;\left(\sum_{a}\langle{a}|Q\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle\langle{a}|Q\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)Q^{\dagger}|{a}\rangle\right)^{2}$ (143)
    $=$ $\;d\;\mathbb{E}_{\phi}\langle{\phi}|\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}\langle{\phi}|\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}+d(d-1)\mathbb{E}_{\phi,\phi^{\bot}}\langle{\phi\phi^{\bot}}|(\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes 2}|{\phi\phi^{\bot}}\rangle\langle{\phi\phi^{\bot}}|(\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes 2}|{\phi\phi^{\bot}}\rangle,$ (144)
    where $\phi^{\bot}$ is randomly sampled from the orthogonal space of $\phi$, i.e., $\langle\phi|\phi^{\bot}\rangle=0$. As shown in [15, Lemma 23 and Eq. (194)], we have
    $d\;\mathbb{E}_{\phi}\langle{\phi}|\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}\langle{\phi}|\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)|{\phi}\rangle^{2}=d\mathcal{O}\left(\frac{1}{d^{4}}\right)=\mathcal{O}\left(\frac{1}{d^{3}}\right),$ (145)
    $d(d-1)\mathbb{E}_{\phi,\phi^{\bot}}\langle{\phi\phi^{\bot}}|(\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes 2}|{\phi\phi^{\bot}}\rangle\langle{\phi\phi^{\bot}}|(\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes 2}|{\phi\phi^{\bot}}\rangle=\frac{(1+\operatorname{Tr}[\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)])^{2}}{d(d+1)}+\mathcal{O}\left(\frac{1}{d^{4}}\right).$ (146)
    Therefore, we have
    $\mathbb{V}_{Q,\psi}\mathbb{E}\left[\tilde{g}(Q,\psi,S|Q,\psi)\right]$ $\leq\mathcal{O}\left(\frac{1}{d^{3}}\right)+\left(\frac{1}{d(d+1)}-\frac{1}{(d+1)^{2}}\right)\mathbb{E}_{\psi}(1+\operatorname{Tr}[\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)])^{2}=\mathcal{O}\left(\frac{1}{d^{3}}\right).$ (147)
Combining the above two terms, we have
$\mathbb{V}(\tilde{g}(Q,\psi,S))\leq\mathcal{O}\left(\frac{1}{m^{2}d}+\frac{1}{md^{2}}+\frac{1}{d^{3}}\right).$ (148)
With Lemma 18 and Eq. (132), the variance of estimator $\omega$ can be bounded by
$\mathbb{V}(\tilde{\omega}_{t})\leq\mathcal{O}\left(\frac{d}{m^{2}}+\frac{1}{m}+\frac{1}{d}\right)\quad\Rightarrow\quad\mathbb{V}({\omega})\leq\mathcal{O}\left(\frac{d}{Tm^{2}}+\frac{1}{Tm}+\frac{1}{Td}\right).$ (149)
Therefore, to achieve $\varepsilon$ additive error, the required query times must satisfy
$Tm=\mathcal{O}\left(\max\left\{\frac{1}{\varepsilon^{2}},\frac{\sqrt{d}}{\varepsilon}\right\}\right).$ (150)
Appendix D DSEC with Independent Classical Shadow
In this section, we consider completing DSEC with independent classical shadow estimation for unitary channels (CSEU) [37]. In Appendix D.1, we describe the algorithm for completing DSEC and analyze the corresponding query complexity given in Proposition 6 of the main text. In Appendix D.2, we provide several lemmas used in the proof.
D.1 Algorithm & Proof of Proposition 6
We first briefly introduce CSEU. Given an unknown unitary channel $\mathcal{U}(\cdot)=U(\cdot)U^{\dagger}$, we can obtain the following classical snapshot by inputting the state $|{\psi}\rangle^{\otimes s}$ and performing the POVM $\mathcal{M}_{s}$ (defined in Eq. (4)),
$\tilde{X}={\rm Snap}(\psi,\phi,s)\mathrel{\mathop{\ordinarycolon}}=\frac{d(d+1)(d+s){\phi}\otimes{\psi}^{T}-(d+1+s)(I\otimes I)}{s},$ (151)
where $|{\psi}\rangle$ is sampled from a state $4$-design ensemble, and $|{\phi}\rangle$ is the measurement result of the POVM $\mathcal{M}_{s}$. Here $\phi\mathrel{\mathop{\ordinarycolon}}=|{\phi}\rangle\!\langle{\phi}|$ is the shorthand of pure state $|{\phi}\rangle$ and similar notation is used for $\psi$. Let $J_{U}$ be the Choi operator of unitary channel $\mathcal{U}$. As shown in [37, Eq. (10)], the expectation of $\tilde{X}$ is $J_{U}$, i.e.,
$\mathbb{E}\tilde{X}=J_{U}.$ (152)
See more details regarding CSEU in [37].
We can complete DSEC via CSEU as follows. First, we independently perform CSEU on each quantum device, and obtain two sets of classical snapshots of $\mathcal{U}(\cdot)=U(\cdot)U^{\dagger},\mathcal{V}(\cdot)=V(\cdot)V^{\dagger}$, labeled as $\{\tilde{X}\}$ and $\{\tilde{Y}\}$. Then, we can estimate the similarity $\operatorname{Tr}[J_{U}J_{V}]/d^{2}$ with these independent classical snapshots, where $J_{U}$ and $J_{V}$ are the Choi operators of $\mathcal{U}$ and $\mathcal{V}$, respectively. We summarize this algorithm in Algorithm 3. In the following, we prove that this algorithm can indeed complete DSEC and analyze its query complexity.
Input: number of SPAM settings $T$,
    the size of symmetric collective measurement $s$,
    TsTs queries to unknown unitary channels $\mathcal{U}$ and $\mathcal{V}$ acting on a $d$-dimensional Hilbert space $\mathcal{H}$.
Output: an estimation of $\operatorname{Tr}[J_{U}J_{V}]/d^{2}$.
1 for $t=1,\cdots,T$ do
2
3  Randomly generate pure states $|{\psi_{t,A}}\rangle^{\otimes s}$ and apply the unitary channels $\mathcal{U}^{\otimes s}$.
4  Measure $(\mathcal{U}(|{\psi_{t,A}}\rangle\!\langle{\psi_{t,A}}|))^{\otimes s}$ with the POVM $\mathcal{M}_{s}$ and obtain the result $|{\phi_{t,A}}\rangle$.
5  Randomly generate pure states $|{\psi_{t,B}}\rangle^{\otimes s}$ and apply the unitary channels $\mathcal{V}^{\otimes s}$.
6  Measure $(\mathcal{V}(|{\psi_{t,B}}\rangle\!\langle{\psi_{t,B}}|))^{\otimes s}$ with the POVM $\mathcal{M}_{s}$ and obtain the result $|{\phi_{t,B}}\rangle$.
7  Compute and store the classical snapshots with Eq. (151):
$\tilde{X}_{t}={\rm Snap}\left(\phi_{t,A},\psi_{t,A},s\right),\quad\tilde{Y}_{t}={\rm Snap}\left(\phi_{t,B},\psi_{t,B},s\right).$ (153)
Return
$\tilde{\gamma}=\frac{1}{T^{2}d^{2}}\sum_{i,j=1}^{T}\operatorname{Tr}\left[\tilde{X}_{i}\tilde{Y}_{j}\right].$ (154)
Algorithm 3 Distributed Similarity Estimation for Unitary Channels with CSEU
First, we prove that $\tilde{\gamma}$ is an unbiased estimator of $\operatorname{Tr}[J_{U}J_{V}]/d^{2}$:
$\mathbb{E}\tilde{\gamma}$ $=\frac{1}{d^{2}}\mathbb{E}\operatorname{Tr}\left[\tilde{X}\tilde{Y}\right]=\frac{1}{d^{2}}\operatorname{Tr}\left[J_{U}J_{V}\right].$ (155)
Then, we analyze the query complexity of this classical shadow-based algorithm. The variance of estimator $\tilde{\gamma}$ is given by
[…]
$=\frac{1}{T^{4}d^{4}}\mathbb{E}\left(\sum_{i,j,k,l=1}^{T}\operatorname{Tr}\left[\tilde{X}_{i}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}\tilde{Y}_{l}\right]\right)-\frac{\operatorname{Tr}^{4}[U^{\dagger}V]}{d^{4}}.$ (157)
After expanding the expectation in the above equation, there are four terms:
  1. 1.
    For $i=k$ and $j=l$, there are $T^{2}$ terms, and using Lemma 20, we have
    $\mathbb{E}\left(\sum_{i,j}\operatorname{Tr}^{2}\left[\tilde{X}_{i}\tilde{Y}_{j}\right]\right)=T^{2}\mathbb{E}\operatorname{Tr}^{2}\left[XY\right]=T^{2}\mathcal{O}\left(\frac{d^{4}(d+s)^{2}}{s^{2}}\right).$ (158)
  2. 2.
    For $i\neq k$ and $j\neq l$, there are $T^{2}(T-1)^{2}$ terms, and we have
    $\mathbb{E}\left(\sum_{i\neq k,j\neq l}\operatorname{Tr}\left[\tilde{X}_{i}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}\tilde{Y}_{l}\right]\right)=T^{2}(T-1)^{2}\operatorname{Tr}^{4}[U^{\dagger}V].$ (159)
  3. 3.
    For $i=k$ and $j\neq l$, there are $T^{2}(T-1)$ terms, and using Lemma 21 we have
    $\mathbb{E}\left(\sum_{i=k,j\neq l}\operatorname{Tr}\left[\tilde{X}_{i}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}\tilde{Y}_{l}\right]\right)=T^{2}(T-1)\mathbb{E}\operatorname{Tr}\left[X\tilde{Y}_{1}\right]\operatorname{Tr}\left[X\tilde{Y}_{2}\right]\leq\mathcal{O}\left(T^{3}d^{4}\right).$ (160)
  4. 4.
    For $i\neq k$ and $j=l$, there are $T^{2}(T-1)$ terms, and using Lemma 21 we have
    $\mathbb{E}\left(\sum_{i\neq k,j=l}\operatorname{Tr}\left[\tilde{X}_{i}\tilde{Y}_{j}\right]\operatorname{Tr}\left[\tilde{X}_{k}\tilde{Y}_{l}\right]\right)=T^{2}(T-1)\mathbb{E}\operatorname{Tr}\left[\tilde{X}_{1}Y\right]\operatorname{Tr}\left[\tilde{X}_{2}Y\right]\leq\mathcal{O}\left(T^{3}d^{4}\right).$ (161)
Therefore, we have
[…]
$\leq\mathcal{O}\left(\frac{(d+s)^{2}}{T^{2}s^{2}}+\frac{1}{T}\right)=\mathcal{O}\left(\frac{d^{2}}{T^{2}s^{2}}+\frac{1}{T}\right).$ (163)
Thus, to achieve $\varepsilon$ additive error, we require the number of queries to satisfy
$Ts=\mathcal{O}\left(\max\left\{\frac{d}{\varepsilon},\frac{1}{\varepsilon^{2}}\right\}\right).$ (164)
Therefore, we complete the proof of Proposition 6.
D.2 Technical Lemmas
Lemma 19 (Lemma D1 in [37]).
Suppose that $\phi$ and $\psi$ are the random input states and measurement outcomes of learning a unitary channel $\mathcal{U}$, then we have
$\mathbb{E}\phi^{\otimes 2}\otimes\psi^{\otimes 2}$ $=\frac{2}{(d+s)(d+s+1)}\sum_{i=1}^{4}\Delta_{U,i},$ (165)
where
$\Delta_{U,1}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{1}{\kappa_{2}}\Pi_{\rm sym}^{(d,2)}\otimes\Pi_{\rm sym}^{(d,2)},$ (166)
$\Delta_{U,2}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{s}{\kappa_{3}}(I\otimes U\otimes I\otimes I)\left[I_{1}\otimes\left(\Pi_{\rm sym}^{(d,3)}\right)_{2,3,4}\right](I\otimes U^{\dagger}\otimes I\otimes I)\left(\Pi_{\rm sym}^{(2)}\otimes I\otimes I\right),$ (167)
$\Delta_{U,3}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{s}{\kappa_{3}}(U\otimes I\otimes I\otimes I)\left[I_{2}\otimes\left(\Pi_{\rm sym}^{(d,3)}\right)_{1,3,4}\right](U^{\dagger}\otimes I\otimes I\otimes I)\left(\Pi_{\rm sym}^{(d,2)}\otimes I\otimes I\right),$ (168)
$\Delta_{U,4}$ $\mathrel{\mathop{\ordinarycolon}}=\frac{s(s-1)}{2\kappa_{4}}(U\otimes U\otimes I\otimes I)\Pi_{\rm sym}^{(d,4)}(U^{\dagger}\otimes U^{\dagger}\otimes I\otimes I).$ (169)
Lemma 20.
Let $X$ and $Y$ be classical snapshots defined in Eq. (151) with POVM $\mathcal{M}_{s}$ for unitary channels $\mathcal{U}$ and $\mathcal{V}$, we have
$\mathbb{E}\operatorname{Tr}^{2}\left[XY\right]\leq\mathcal{O}\left(\frac{d^{4}(d+s)^{2}}{s^{2}}\right).$ (170)
Proof.
Using Lemma 9 and the definition of $X,Y$, we have
$\mathbb{E}\operatorname{Tr}^{2}\left[XY\right]=\frac{1}{s^{2}}\operatorname{Tr}^{2}\left[\left(d(d+1)(d+s){\phi_{1}}\otimes{\psi_{1}}^{T}-(d+1+s)I\otimes I\right)\left(d(d+1)(d+s){\phi_{2}}\otimes{\psi_{2}}^{T}-(d+1+s)I\otimes I\right)\right]$ (171)
$=\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{1}\phi_{2}\otimes\psi_{2}\psi_{2}\right]-\frac{d(d+1)(d+s)(d+1+s)}{s^{2}}\left(\sum_{i=1}^{2}\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{i}\otimes\psi_{i}\right]\right)+\frac{d^{4}(d+1+s)^{2}}{s^{2}}.$ (172)
We consider the first two terms respectively as follows:
  1. 1.
    For the first term, with Lemma 19, there are 1616 terms:
    $\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{1}\phi_{2}\otimes\psi_{2}\psi_{2}\right]$ $=\mathbb{E}\operatorname{Tr}\left[\left(\phi_{1}^{\otimes 2}\otimes\psi_{1}^{\otimes 2}\right)\left(\phi_{2}^{\otimes 2}\otimes\psi_{2}^{\otimes 2}\right)\right]=\frac{4}{(d+s)^{2}(d+s+1)^{2}}\sum_{k,l=1}^{4}\operatorname{Tr}\left[\Delta_{U,k}\Delta_{V,l}\right]$ (173)
    We analyze these terms as follows:
    $\operatorname{Tr}[\Delta_{U,1}\Delta_{V,1}]$ $=\frac{\operatorname{Tr}^{2}\left[\Pi_{\rm sym}^{(d,2)}\right]}{\kappa_{2}^{2}}=1,$ (174)
    $\operatorname{Tr}[\Delta_{U,1}\Delta_{V,2}]$ $=\operatorname{Tr}[\Delta_{U,2}\Delta_{V,1}]=\operatorname{Tr}[\Delta_{U,1}\Delta_{V,3}]=\operatorname{Tr}[\Delta_{U,3}\Delta_{V,1}]=\frac{s}{\kappa_{2}\kappa_{3}}\frac{\kappa_{2}\kappa_{3}}{d}=\frac{s}{d},$ (175)
    $\operatorname{Tr}\left[\Delta_{U,1}\Delta_{V,4}\right]$ $=\operatorname{Tr}\left[\Delta_{U,4}\Delta_{V,1}\right]=\frac{s(s-1)}{d(d+1)}=\mathcal{O}\left(\frac{s^{2}}{d^{2}}\right),$ (176)
    $\operatorname{Tr}\left[\Delta_{U,2}\Delta_{V,2}\right]$ $=\operatorname{Tr}\left[\Delta_{U,3}\Delta_{V,3}\right]=\frac{s^{2}(d^{2}+2d+\operatorname{Tr}^{2}[U^{\dagger}V])}{d^{2}(d+1)^{2}}=\mathcal{O}\left(\frac{s^{2}}{d^{2}}\right),$ (177)
    $\operatorname{Tr}\left[\Delta_{U,2}\Delta_{V,3}\right]$ $=\operatorname{Tr}\left[\Delta_{U,3}\Delta_{V,2}\right]=\frac{s^{2}(d^{2}+2d+\operatorname{Tr}^{2}[U^{\dagger}V])}{d^{2}(d+1)^{2}}=\mathcal{O}\left(\frac{s^{2}}{d^{2}}\right),$ (178)
    $\operatorname{Tr}\left[\Delta_{U,2}\Delta_{V,4}\right]$ $\leq\frac{s^{2}(s-1)d^{4}}{2\kappa_{3}\kappa_{4}}=\mathcal{O}\left(\frac{s^{3}}{d^{3}}\right),\quad\operatorname{Tr}\left[\Delta_{U,4}\Delta_{V,2}\right]\leq\mathcal{O}\left(\frac{s^{3}}{d^{3}}\right),$ (179)
    $\operatorname{Tr}\left[\Delta_{U,3}\Delta_{V,4}\right]$ $\leq\mathcal{O}\left(\frac{s^{3}}{d^{3}}\right),\quad\operatorname{Tr}\left[\Delta_{U,4}\Delta_{V,3}\right]\leq\mathcal{O}\left(\frac{s^{3}}{d^{3}}\right),\quad\operatorname{Tr}\left[\Delta_{U,4}\Delta_{V,3}\right]\leq\frac{s^{2}(s-1)^{2}d^{4}}{4\kappa_{4}}=\mathcal{O}\left(\frac{s^{4}}{d^{4}}\right).$ (180)
    Therefore, for the first term, we have
    $\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{1}\phi_{2}\otimes(\psi_{2}\psi_{2})^{T}\right]$ $\leq\frac{4d^{2}(d+1)^{2}}{s^{2}(d+s+1)^{2}}\mathcal{O}\left(1+\frac{s}{d}+\frac{s^{2}}{d^{2}}+\frac{s^{3}}{d^{3}}+\frac{s^{4}}{d^{4}}\right)$ (181)
    $=\mathcal{O}\left(\frac{d^{4}}{s^{2}(d+s)^{2}}+\frac{s^{2}}{(d+s)^{2}}\right)$ (182)
  2. 2.
    For the second term, with Lemma 19, we have
    $\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{1}\otimes\psi^{T}_{1}\right]$ $=\mathbb{E}\operatorname{Tr}\left[\phi_{1}^{\otimes 2}\otimes\psi_{1}^{\otimes 2}\right]=\frac{2}{(d+s)(d+s+1)}\sum_{i=1}^{4}\operatorname{Tr}\left[\Delta_{U,i}\right]$ (183)
    $=\frac{2}{(d+s)(d+s+1)}\left[\kappa_{2}+\frac{2\cdot d(d+1)^{2}(d+2)}{12\cdot\kappa_{3}}+\frac{s(s-1)}{2}\right]=\mathcal{O}(1).$ (184)
    Likewise, $\mathbb{E}\operatorname{Tr}^{2}\left[\phi_{2}\otimes\psi^{T}_{2}\right]=\mathcal{O}(1)$.
Therefore, we have
$\mathbb{E}\operatorname{Tr}^{2}\left[XY\right]$ $\leq\mathcal{O}\left(\frac{d^{4}}{s^{2}(d+s)^{2}}+\frac{s^{2}}{(d+s)^{2}}\right)+\frac{d^{4}(d+1+s)^{2}}{s^{2}}=\mathcal{O}\left(\frac{d^{4}(d+s)^{2}}{s^{2}}\right).$ (185)
Lemma 21.
Let $X_{1}$ and $X_{2}$ be classical snapshots defined in Eq. (151) with POVM $\mathcal{M}_{s}$ for unitary channel $\mathcal{U}$, and $Y$ be a classical snapshot for unitary channel $\mathcal{V}$, we have
$\mathbb{E}\operatorname{Tr}\left[X_{1}Y\right]\operatorname{Tr}\left[X_{2}Y\right]$ $\leq\mathcal{O}\left(d^{4}\right).$ (186)
Let $Y_{1}$ and $Y_{2}$ be classical snapshots defined in Eq. (151) with POVM $\mathcal{M}_{s}$ for unitary channel $\mathcal{V}$, and $X$ be a classical snapshot for unitary channel $\mathcal{U}$, we have
$\mathbb{E}\operatorname{Tr}\left[XY_{1}\right]\operatorname{Tr}\left[XY_{2}\right]$ $\leq\mathcal{O}\left(d^{4}\right).$ (187)
Proof.
With the definition in Eq. (151), we have
$\;\mathbb{E}\operatorname{Tr}\left[X_{1}Y\right]\operatorname{Tr}\left[X_{2}Y\right]=\operatorname{Tr}\left[\left(\mathbb{E}X\right)^{\otimes 2}\mathbb{E}Y^{\otimes 2}\right]$ (188)
$=\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\operatorname{Tr}\left[J_{U}^{\otimes 2}\mathbb{E}\left(\phi_{2}\otimes\psi_{2}^{T}\right)^{\otimes 2}\right]-\frac{2d^{2}(d+1)(d+s)(d+s+1)}{s^{2}}\operatorname{Tr}\left[J_{U}\mathbb{E}\phi_{2}\otimes\psi_{2}^{T}\right]+\frac{d^{2}(d+s+1)^{2}}{s^{2}}.$ (189)
With Lemma 1 in [37],
$\operatorname{Tr}\left[J_{U}\mathbb{E}\phi_{2}\otimes\psi_{2}^{T}\right]$ $=\frac{d+s+1}{(d+1)(d+s)}+\frac{s}{d(d+1)(d+s)}\operatorname{Tr}\left[J_{U}(V\otimes V^{\dagger})\mathbb{F}\right]$ (190)
$=\frac{d+s+1}{(d+1)(d+s)}+\frac{s}{d(d+1)(d+s)}\operatorname{Tr}^{2}\left[U^{\dagger}V\right].$ (191)
Thus, we have
$\mathbb{E}\operatorname{Tr}\left[X_{1}Y\right]\operatorname{Tr}\left[X_{2}Y\right]$ $=\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\operatorname{Tr}\left[J_{U}^{\otimes 2}\mathbb{E}\left(\phi_{2}\otimes\psi_{2}^{T}\right)^{\otimes 2}\right]-\frac{2d(d+s+1)}{s}\operatorname{Tr}^{2}\left[U^{\dagger}V\right]-\frac{d^{2}(d+s+1)^{2}}{s^{2}}.$ (192)
Now, we focus on the first term. With Lemma 19, we have
$\operatorname{Tr}\left[J_{U}^{\otimes 2}\mathbb{E}\left(\phi_{2}\otimes\psi_{2}^{T}\right)^{\otimes 2}\right]$ $=\sum_{i,j,k,l=0}^{d-1}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\mathbb{E}\phi^{\otimes 2}_{2}\otimes\psi^{\otimes 2}_{2}\right]$ (193)
$=\frac{2}{(d+s)(d+s+1)}\sum_{i=1}^{4}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\Delta_{V,i}\right]\leq\mathcal{O}\left(\frac{s^{2}}{(d+s)(d+s+1)}\right),$ (194)
with the following calculation,
$\sum_{i,j,k,l}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\Delta_{V,1}\right]$ $=\frac{s}{\kappa_{2}}\sum_{i,j,k,l}\operatorname{Tr}^{2}\left[\Pi_{\rm sym}^{(d,2)}|{jl}\rangle\!\langle{ik}|\right]=\frac{sd(d+1)}{2\cdot\kappa_{2}}=s,$ (195)
$\sum_{i,j,k,l}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\Delta_{V,2}\right]$ $=\frac{2s\left(\operatorname{Tr}^{2}[U^{\dagger}V]+d\right)(d+2)}{12\cdot\kappa_{3}}=\frac{s\operatorname{Tr}^{2}[U^{\dagger}V]}{d(d+1)}+\frac{s}{d+1}\leq s,$ (196)
$\sum_{i,j,k,l}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\Delta_{V,3}\right]$ $=\frac{2s\left(\operatorname{Tr}^{2}[U^{\dagger}V]+d\right)(d+2)}{12\cdot\kappa_{3}}=\frac{s\operatorname{Tr}^{2}[U^{\dagger}V]}{d(d+1)}+\frac{s}{d+1}\leq s,$ (197)
$\sum_{i,j,k,l}\operatorname{Tr}\left[\left(U^{\otimes 2}|{ik}\rangle\!\langle{jl}|U^{\dagger\otimes 2}\otimes|{jl}\rangle\!\langle{ik}|\right)\Delta_{V,4}\right]$ $\leq\frac{s(s-1)}{2}.$ (198)
Therefore, we have
$\mathbb{E}\operatorname{Tr}\left[X_{1}Y\right]\operatorname{Tr}\left[X_{2}Y\right]$ $\leq\frac{d^{2}(d+1)^{2}(d+s)^{2}}{s^{2}}\mathcal{O}\left(\frac{s^{2}}{(d+s)(d+s+1)}\right)=\mathcal{O}\left(d^{4}\right).$ (199)
The same bound holds for $\mathbb{E}\operatorname{Tr}[XY_{1}]\operatorname{Tr}[XY_{2}]$ by symmetry. ∎